### v1

# Deep Linear Hawkes Processes

Abstract
Marked temporal point processes (MTPPs) are used to model sequences of different types of events with irregular arrival times, with broad applications ranging from healthcare and social networks to finance. We address shortcomings in existing point process models by drawing connections between modern deep state-space models (SSMs) and linear Hawkes processes (LHPs), culminating in an MTPP that we call the deep linear Hawkes process (DLHP). The DLHP modifies the linear differential equations in deep SSMs to be stochastic jump differential equations, akin to LHPs. After discretizing, the resulting recurrence can be implemented efficiently using a parallel scan. This brings parallelism and linear scaling to MTPP models. This contrasts with attention-based MTPPs, which scale quadratically, and RNN-based MTPPs, which do not parallelize across the sequence length. We show empirically that DLHPs match or outperform existing models across a broad range of metrics on eight real-world datasets. Our proposed DLHP model is the first instance of the unique architectural capabilities of SSMs being leveraged to construct a new class of MTPP models.
1 Introduction
Figure 1: Intensity estimates from trained models when conditioned on an empty sequence $\mathcal{H}_{t}=\emptyset$ for NHP (Mei & Eisner, 2017) and DLHP, our method. Shown in dotted lines are the ground truth, inhomogeneous Poisson process intensity. Our DLHP is able to accurately capture the background intensity. See Section 5.1 for more details.
Marked temporal point processes (MTPPs) are used to model irregular sequences of events in continuous-time, where each event has an associated type, often referred to as a mark. MTPPs model the joint distribution of marked event sequences. They have been successfully applied to modeling purchasing patterns in e-commerce (Türkmen et al., 2019; Vassøy et al., 2019; Yang et al., 2018), patient-specific medical events (Hua et al., 2022), disease propagation (Gajardo & Müller, 2023), and many other domains (Williams et al., 2020; Sharma et al., 2018; Wang et al., 2024).
An MTPP is fully characterized by a marked intensity process which specifies the expected instantaneous rate of occurrence of events of each mark conditioned on the event history. State-of-the-art methods use neural networks to compute hidden states that summarize the event history, which are then used to compute marked intensities across future values of time. However, many models are limited by inexpressive temporal dynamics, lack of support for long-range dependencies, and serial computation (Du et al., 2016; Mei & Eisner, 2017). Recent advances in transformer-based MTPPs have improved performance and gained parallelism, but scale quadratically with sequence lengths (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022).
Recently, deep state-space models (often abbreviated as SSMs) have emerged as a challenger to transformer-based models for discrete sequence modeling (Gu et al., 2022b; Smith et al., 2022; Gu & Dao, 2023). SSMs interleave a stack of linear state-space recurrences with position-wise non-linearities (Gu et al., 2021). This architecture has been found to be not only highly performant on a wide range of tasks (Goel et al., 2022; Deng et al., 2024), but retains linear scaling, can be parallelized across the length of a sequence, and can gracefully handle irregularly-spaced observations.
Inspired by this, we revisit a foundational point process model, the linear Hawkes process (LHP Hawkes, 1971), and draw connections between LHPs and deep SSMs. We combine the parameterization and parallelization strategy of SSMs with the functional form of LHPs to create what we call the deep linear Hawkes process (DLHP). More formally, the DLHP is a fully-recurrent neural MTPP parameterized by a stack of stochastic jump differential equations on the complex plane (serving as the recurrence) interleaved with position-wise non-linearities (to improve expressivity). This design yields an MTPP with two main advantages over existing neural MTPPs: (i) parallelism across the length of the sequence through the use of parallel scans, and (ii) highly flexible intensity functions. This is achieved not only through the expressivity of the SSM-style architecture, but also by tying the output intensity at time $t$ to the model’s continuously-evolved hidden state $\mathbf{x}_{t}$ (extending ideas from Mei & Eisner (2017) and Yang et al. (2022), see Figs. 1 and 2), and by going beyond the classical LHP form with input-dependent recurrent dynamics (akin to Mamba (Gu & Dao, 2023)).
The contributions of this paper are as follows: We introduce a new family of marked point process models, deep linear Hawkes processes—the first MTPP model that fully leverages the architectural features of deep SSMs. We demonstrate that DLHPs match or exceed the performance of existing models across eight real-world datasets, with an average per-event likelihood improvement of 38% across datasets, over the individually best-performing existing method per dataset. We also verify that DLHP scales more effectively to longer sequences, a crucial capability for a wide range of applications. We will release our models, datasets, and pipelines as part of the existing EasyTPP library (Xue et al., 2023). We conclude by discussing the relative advantages and disadvantages of the DLHP over existing methods and opportunities for extending this work.
Figure 2: Three different schematics of the deep linear Hawkes process (DLHP) and latent linear Hawkes (LLH) layer we propose. With increasing granularity: Left (a): On a high level, the DLHP can simply be viewed as a deep stack of neural network layers that transform an event sequence into an intensity function. Middle (b): On a more granular level, individual LLH layers can be viewed as discrete-time recurrences (see Eq. 16), directly defining an intensity evaluated at select times: $t^{*}$ using $\mathbf{x}_{t^{*}}$, right limits $t_{i}+$ using $\mathbf{x}_{t_{i}}$, and left values $t_{i}$ using $\mathbf{x}_{t_{i}-}$. Right (c): The same recurrences can be viewed as a set of non-linearly coupled stochastic jump differential equations in continuous-time. Events are embedded and impart impulses to the differential equation. Impulses for layer $L$ are omitted for visual clarity. Intermediate intensities are evaluated using $\vec{0}$ for $\mathbf{u}^{(1)}$ and no impulse.
2 Preliminaries
2.1 Marked Temporal Point Processes
Let $t_{1},t_{2},\dots\in\mathbb{R}_{\geq 0}$ be a strictly increasing sequence of positive random variables, each representing the time of occurrence for an event of interest. [^1] Please refer to Tables 2 and 3 in Appendix A for a list of notation and acronym definitions, respectively. For each $t_{i}$, let $k_{i}\in\mathcal{M}$ be a random variable representing accompanying side-information, commonly referred to as an event’s mark, with $\mathcal{M}$ being the mark-space. In this paper, we focus on discrete and finite mark spaces, i.e., :=$\mathcal{M}:=\{1,\dots,K\}$; however, in general $\mathcal{M}$ can be continuous or even a mixture of continuous and discrete. Together $t_{i}$ and $k_{i}$ fully define a given event. The joint distribution over a sequence of continuous event times and mark types is described as a marked temporal point process. We use $\mathcal{H}_{t}$ to represent the sequence, or history, of events up to some time $t$: $\mathcal{H}_{t}:=\{(t_{i},k_{i})\;|\;t_{i}\leq t\text{ for }i\in\mathbb{N}\}$, with $\mathcal{H}_{t-}$ defined similarly except that it does not include events that occur at time $t$.
One way of characterizing an MTPP is through a marked intensity process, which describes the instantaneous expected rate of occurrence for events of specific marks. Let $\mathbf{N}_{t}:=[N_{t}^{1},\dots,N_{t}^{K}]^{\top}\in\mathbb{Z}_{\geq 0}^{K}$ be the marked counting process which represents the number of occurrences of events of each type of mark in the time span $[0,t]$. The marked intensity process $\bm{\lambda}_{t}:=[\lambda_{t}^{1},\dots,\lambda_{t}^{K}]^{\top}\in\mathbb{R}_{\geq 0}^{K}$ characterizes an MTPP by describing how the counting process changes via:
$\lambda_{t}^{k}\mathrm{d}t:=$ $\;\mathbb{E}\left[\text{event of type }k\text{ occurs in }[t,t+\mathrm{d}t]\;|\;\mathcal{H}_{t-}\right]=\;\mathbb{E}\left[N^{k}_{t+\mathrm{d}t}-N^{k}_{t}\;|\;\mathcal{H}_{t-}\right],$ (1)
with the total intensity $\lambda_{t}:=\sum_{k=1}^{K}\lambda^{k}_{t}$ being the rate that any event occurs. Note that the intensity conditions on the left limit of the history $\mathcal{H}_{t-}$ to ensure that the intensity is modeling future events.
Parameterized forms of $\bm{\lambda}$ are often trained by optimizing the log-likelihood over observed data. The log-likelihood for a single sequence $\mathcal{H}_{T}$ is defined as (Daley & Vere-Jones, 2003, ch. 7.3):
$\mathcal{L}(\mathcal{H}_{T}):=\sum\nolimits_{i=1}^{|\mathcal{H}_{T}|}\log\lambda_{t_{i}}^{k_{i}}-\int_{0}^{T}\lambda_{s}\mathrm{d}s.$ (2)
Linear Hawkes Processes  An (unmarked) Hawkes process (Hawkes, 1971), or more generally a self-exciting process, is a temporal point process where event occurrences increase the rate at which subsequent events occur soon thereafter. Of particular interest to us are linear Hawkes processes (LHPs), which are characterized by the following intensity process:
$\lambda_{t}:=\nu+\int_{s=0}^{t-}h(t-s)\mathrm{d}N_{s}:=\nu+\sum\nolimits_{i=1}^{N_{t-}}h(t-t_{i}),$ (3)
where $\nu>0$ is the background intensity, $h:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is the excitation function (or kernel), and $N_{t}$ is the associated counting process characterized by intensity $\lambda_{t}$. $N_{t-}$ is used as the upper limit in Eq. 3 to ensure the intensity at time $t$ does not take into account an event that occurs at time $t$.
Should $h$ correspond to the exponential decay kernel, $h(z)=\alpha\exp\left(-\beta z\right)$, then the LHP intensity process is Markov (Law & Viens, 2016) and admits the following stochastic differential form:
$\mathrm{d}\lambda_{t}=\beta(\nu-\lambda_{t-})\mathrm{d}t+\alpha\mathrm{d}N_{t}\iff\lambda_{t}$ $=\nu+\int_{0}^{t-}\alpha\exp\left(-\beta(t-s)\right)\mathrm{d}N_{s}$ (4)
$=\nu+\sum\nolimits_{i=1}^{N_{t-}}\alpha\exp\left(-\beta(t-t_{i})\right).$ (5)
LHPs can be extended to the marked setting, with $K$ possible discrete marks, by replacing $\nu$ with a vector of $K$ background rates $\bm{\nu}:=[\nu_{1},\dots,\nu_{K}]^{\top}$, and the excitation effect $h(t-s)\mathrm{d}N_{s}$ with $\mathbf{h}(t-s)\mathrm{d}\mathbf{N}_{s}$. Here, $\mathbf{h}_{ij}$ of $\mathbf{h}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{K\times K}$ describes the excitation that events of type $i$ exerts on future events of type $j$. The counting process, $\mathrm{d}\mathbf{N}_{t}$, is then either a $K$-dimensional zero-vector if no event occurs at time $t$, or a one-hot vector indicating which mark is associated with the occurring event. Generalizing the exponential kernel to handle marks results in the following differential form:
$\mathrm{d}\bm{\lambda}_{t}=-\bm{\beta}(\bm{\lambda}_{t-}-\bm{\nu})\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (6)
where $\bm{\beta,\alpha}\in\mathbb{R}^{K\times K}_{\geq 0}$ are restricted to be non-negative to ensure non-negative marked intensities.
2.2 State-Space Models
Deep state-space models (SSMs) are a recent innovation in recurrent models that have found success in long-range sequence modeling tasks (Gu et al., 2022b) and language modeling tasks (Gu & Dao, 2023), while also having favorable computational properties. The backbone of deep SSMs is the linear state-space equations, which define a continuous-time dynamical system with input and output signals $\textbf{u}(t),\mathbf{y}(t)\in\mathbb{R}^{H}$, respectively, through linear differential equations:
$\frac{\mathrm{d}}{\mathrm{d}t}\textbf{x}(t)$ $=\mathbf{A}\textbf{x}(t)+\mathbf{B}\mathbf{u}(t)$ (7)
$\mathbf{y}(t)$ $=\mathbf{C}\mathbf{x}(t)+\mathbf{D}\mathbf{u}(t),$ (8)
where $\textbf{x}(t)\in\mathbb{R}^{P}$ is the (hidden) state of the system, and $\mathbf{A}\in\mathbb{R}^{P\times P},\mathbf{B}\in\mathbb{R}^{P\times H},\mathbf{C}\in\mathbb{R}^{H\times P},$ and $\mathbf{D}\in\mathbb{R}^{H\times H}$ are the parameters that control the system’s dynamics.
Deep SSMs then stack these recurrences interleaved with non-linear position-wise functions, $\sigma$. The function $\sigma$ can contain activation functions, residual connections and normalization layers, and transforms the output $\mathbf{y}$ of the previous recurrence into the input $\mathbf{u}$ of the next, i.e., $\mathbf{u}^{(l)}(t):=\sigma(\mathbf{y}^{(l-1)}(t))$ for layer $l$. This combination yields a sequence model where each recurrence is conditionally linear in time given the input, but is ultimately non-linear in depth due to the function $\sigma$.
To evaluate the SSM, we first discretize the continuous-time system at the desired times, and then evaluate as though it were a conventional discrete-time RNN architecture. Crucially, the linearity of the resulting discrete-time recurrence allows it to be evaluated using parallel scans (Blelloch, 1990; Smith et al., 2022; Gu & Dao, 2023), leading to linear work scaling (i.e., number of operations), and, importantly, sublinear scaling of the computation time with respect to sequence length given sufficient parallel compute. Note this contrasts with conventional sequential RNNs (e.g., LSTMs), which process sequences serially; and attention-based methods, which can be parallelized over a sequence, but have quadratic work scaling with respect to sequence length. This allows SSMs to fully and efficiently leverage modern massively parallel hardware while also a being highly expressive and performant model class. Importantly for our purposes, evaluating a linear recurrence with a parallel scan natively admits evaluations with varying observation intervals. We will leverage this to parsimoniously handle the variable inter-event times observed in MTPP settings.
3 Deep Linear Hawkes Processes
In this section, we introduce our deep linear Hawkes process (DLHP), a neural MTPP that draws a novel connection between LHPs and deep SSMs. Stochastic jump differential equations, akin to the LHP intensity, form the basis of the conditionally linear recurrent layer, which we refer to as a latent linear Hawkes (LLH) layer. The LLH layer can be viewed as a modified SSM recurrence, while still admitting parallel computation. Taking further inspiration from deep SSMs, the DLHP is then made up of a stack of LLH layers, each interleaved with non-linear, position-wise transformations to increase the overall expressivity of the model (see Fig. 2). In this section, we formalize this approach and outline implementation details.
3.1 Continuous-Time Latent Linear Hawkes Layer
We first start by generalizing the intensity of the linear Hawkes process, Eq. 6:
$\mathrm{d}\bm{\lambda}_{t}$ $=-\bm{\beta}(\bm{\lambda}_{t-}-\bm{\nu}_{t})\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t}=-\bm{\beta}\bm{\lambda}_{t-}\mathrm{d}t+\bm{\beta}\bm{\nu}_{t}\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (9)
whereby the background intensity $\bm{\nu}_{t}$ is allowed to vary over time. If we compare this to the recurrence in Eq. 7, we see that the intensity in the LHP, $\bm{\lambda}_{t}$ controlled by decay rates $\bm{\beta}$, is analogous the state in the linear SSM, $\mathbf{x}(t)$ controlled by state matrix $\mathbf{A}$. Additionally, the time-varying baseline intensity in the LHP, $\bm{\nu}_{t}$, is analogous to the SSM input signal, $\mathbf{u}(t)$. What is unique to the LHP is the (mark-specific) impulse $\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$. This impulse is important because it allows the model to instantaneously incorporate information from events as they occur, introducing discontinuities in the output signals of the otherwise continuously-integrated system, unlike conventional SSMs.
With this in mind, we adapt Eq. 9 such that it can replace the typical state-space recurrence in an SSM, Eq. 7. To do so, we replace the non-negative $\bm{\beta}$ with an unrestricted state matrix $\mathbf{A}\in\mathbb{R}^{P\times P}$. Next, given an input signal $\mathbf{u}_{t}\in\mathbb{R}^{H}$ we project it to $P$ dimensions with an input matrix $\mathbf{B}\in\mathbb{R}^{P\times H}$ to replace $\bm{\nu}_{t}$. [^2] Here, we index time $t$ via subscripts (e.g., $\mathbf{u}_{t}$) rather than an argument $\left(\mathbf{u}(t)\right)$ to emphasize that these are stochastic (jump) processes rather than deterministic functions. What was originally the intensity $\bm{\lambda}_{t}$ is now relabeled to be the state of the layer $\mathbf{x}_{t}$. Finally, we allow the impulses to be low-rank by having a shared set of mark embeddings $\bm{\alpha}\in\mathbb{R}^{R\times K}$ with rank $R$ that are brought into $P$ dimensions with a layer-specific embedding matrix $\mathbf{E}\in\mathbb{R}^{P\times R}$. For simplicity, we set $R=H$ for all our experiments. The equation for the output signal $\mathbf{y}_{t}$ is left unchanged from Eq. 8, where $\mathbf{C}\in\mathbb{R}^{H\times P}$ and $\mathbf{D}\in\mathbb{R}^{H\times H}$. All of this results in the set of equations that makes up what we call the latent linear Hawkes layer:
$\mathrm{d}\mathbf{x}_{t}$ $=-\mathbf{Ax}_{t-}\mathrm{d}t+\mathbf{ABu}_{t-}\mathrm{d}t+\mathbf{E}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (10)
$\mathbf{y}_{t}$ $=\mathbf{Cx}_{t}+\mathbf{Du}_{t},$ (11)
where the initial state $\mathbf{x}_{0}\in\mathbb{R}^{P}$ is learned. Realizations of this layer can be seen in Fig. 2.
3.2 Continuous-Time Deep Linear Hawkes Process Architecture
Inspired by deep SSMs, our MTPP is formed by stacking LLH layers, chaining the output signal $\mathbf{y}$ of one layer to the input $\mathbf{u}$ of another with non-linear transforms in between. The final layer’s output is then transformed into the intensity $\bm{\lambda}$. An illustration of the DLHP architecture is shown in Fig. 2.
Let $L$ be the number of desired LLH layers that comprise a DLHP with input and output signals $\mathbf{u}^{(l)}$ and $\mathbf{y}^{(l)}$ respectively for layers $l=1,\dots,L$. For the very first layer, the only input available to condition on are the event occurrences themselves. As such, we set $\mathbf{u}_{t}^{(1)}=\mathbf{0}$ for all $t\geq 0$.
In general, a layer’s output $\mathbf{y}^{(l)}:=\text{LLH}^{(l)}(\mathbf{u}^{(l)},\mathcal{H})$ is passed into a non-linear activation function $f$ (we use $f(z):=\text{GELU}(z)$ (Hendrycks & Gimpel, 2016)), summed with the residual stream $\mathbf{u}^{(l)}$, and normalized with LayerNorm (Ba, 2016) to compute the next layer’s input. More formally,
$\mathbf{u}^{(l+1)}_{t}:=\text{LayerNorm}^{(l)}(f(\mathbf{y}^{(l)}_{t})+\mathbf{u}^{(l)}_{t})$ (12)
for $t\geq 0$ and $l=1,\dots,L$. We use the same strategies for initialization as S5 (Smith et al., 2022), based off the performant HiPPO initialization scheme (Gu et al., 2020). Due to the transformations, unlike the original LHP, we cannot guarantee the output of the final layer is positive. Therefore, similar to Mei & Eisner (2017), we apply an affine projection followed by a rectifying transformation to enforce non-negative intensity:
$\bm{\lambda}_{t}:=\mathbf{s}\odot\text{softplus}((\mathbf{W}\mathbf{u}_{t-}^{(L+1)}+\mathbf{b})\odot\mathbf{s}^{-1})$ (13)
for $t\geq 0$ and where $\mathbf{W}\in\mathbb{R}^{K\times H}$, $\mathbf{b},\log(\mathbf{s})\in\mathbb{R}^{K}$, and $\odot$ is an element-wise product. Eq. 13 implements the “Proj. & Softplus” layer in Fig. 2. The intensity at time $t$ always uses the left-limit of $\mathbf{u}^{(L+1)}$, which in turn uses the left-limit of $\mathbf{y}^{(l)}$ and $\mathbf{u}^{(l)}$ for all $l$ to ensure that it has no information of any events that may or may not have occurred at time $t$ is used.
[…]
On the Relationship With the Linear Hawkes Process Before discussing how to compute the DLHP and its variations, we briefly reflect on the relationship between the DLHP and LHP. The derivation presented above shows the steps to modify an LHP to be a deep SSM. The connection, parameterization, and equivalence we explore does not materially affect the implementation; this was intended to concretely define how our model differs from a classical model, and to retain the intuition from the simpler LHP (even if the direct interpretability of the LHP is somewhat lost). Alternatively, one could have simply attempted to convert a deep SSM into an MTPP. Likewise, the steps taken and the result would be similar; however, the relationship to other models would be markedly less clear. We hope our exposition makes it clear how the DLHP is a natural extension of a known and well-used model rather than an arbitrary modification to a recent architecture.
3.3 Discretizing & Diagonalizing the LLH Layer
Unlike the LHP intensity, the recurrence in the LLH layer does not permit an analytic solution. As such, we must discretize the continuous-time process to compute values of the layer at specified time points. If we approximate the input signal by treating it as constant over an update interval, also known as a zero-order hold (ZOH) assumption (Iserles, 2009), then we can achieve a closed-form exact update to the recurrence relation. However, unfortunately, this involves a computationally-expensive matrix exponential in the update rule. To circumvent this, we first diagonalize the system and then impose the zero-order hold restriction on it. Doing so converts the matrix exponential into an element-wise exponential operation. This is done for all LLH layers that compose the DLHP. Note that this is same general approach taken by Smith et al. (2022) for deep SSMs.
Diagonalization Let $-\mathbf{A}$ be diagonalizable with a factorization of $\mathbf{V}\bm{\Lambda}\mathbf{V}^{-1}$, where $\mathbf{V},\bm{\Lambda}\in\mathbb{C}^{P\times P}$ and $\bm{\Lambda}$ is a diagonal matrix of eigenvalues. An equivalent, diagonalized LLH is then
$\mathrm{d}\tilde{\mathbf{x}}_{t}$ $:=\bm{\Lambda}\tilde{\mathbf{x}}_{t-}\mathrm{d}t+\bm{\Lambda}\tilde{\mathbf{B}}\mathbf{u}_{t-}\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (14)
$\mathbf{y}_{t}$ $:=\tilde{\mathbf{C}}\tilde{\mathbf{x}}_{t}+\mathbf{D}\mathbf{u}_{t}$ (15)
where $\tilde{\mathbf{x}}_{t}=\mathbf{V}^{-1}\mathbf{x}_{t}$, $\tilde{\mathbf{B}}=-\mathbf{V}^{-1}\mathbf{B}$, $\tilde{\mathbf{E}}=\mathbf{V}^{-1}\mathbf{E}$, and $\tilde{\mathbf{C}}=\mathbf{CV}$. Note that in practice we directly parameterize $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$, and $\tilde{\mathbf{E}}$ to avoid having to learn and invert $\mathbf{V}$. The eigenvalues $\bm{\Lambda}$ are also directly parameterized and constrained with negative real-components for stability (Davis, 2013). While the dynamics are diagonalized, we note this does not mean that we are independently modeling the intensities of different mark types. This can be seen two ways: First, the diagonalized dynamics are equivalent to the original dynamics (see Eq. 10, given the system can be diagonalized on the complex plane). Alternatively, the marks interact through the dense input and output matrices, the position-wise non-linearity, the mark embeddings, and the final intensity rectification layer.
Discretization We then employ a ZOH discretization to create a closed-form update from the diagonalized continuous-time system. The ZOH assumption holds the input $\mathbf{u}$ constant over the integration period. This results in the following update rule that transitions from $\mathbf{x}_{t}$ to $\mathbf{x}_{t^{\prime}}$, where, by construction, no events occur in $(t,t^{\prime})$:
$\tilde{\mathbf{x}}_{t^{\prime}}$ :=$:=\begin{cases}\bar{\bm{\Lambda}}\tilde{\mathbf{x}}_{t}+(\bar{\bm{\Lambda}}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t^{\prime}-}&\text{ if no event at }t^{\prime}\\ \bar{\bm{\Lambda}}\tilde{\mathbf{x}}_{t}+(\bar{\bm{\Lambda}}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t^{\prime}-}+\tilde{\mathbf{E}}\bm{\alpha}_{k}&\text{ if event of type }k\text{ at }t^{\prime}\end{cases}$ (16)
where $\bar{\bm{\Lambda}}:=\exp(\bm{\Lambda}(t^{\prime}-t))$ (derivation in Section B.2). Please refer to Fig. 2 for an illustration.
Note that the ZOH is an exact update when $\mathbf{u}$ is constant over the window $[t,t^{\prime})$. While we choose the constant value to set $\mathbf{u}$ to be as $\mathbf{u}_{t^{\prime}-}$, it is worth noting that technically any value $\mathbf{u}_{s}$ for $s\in[t,t^{\prime})$ is valid. We explore this design decision and the impact it has on performance in more detail in Sections B.4 and D.4. It is important that $\mathbf{u}_{t^{\prime}}$ is not used as the ZOH value to avoid data leakage.
3.4 Input-Dependent Dynamics
Inspired by recent developments in modern SSMs (e.g., Mamba (Gu & Dao, 2023)), we also consider allowing the dynamics of the system to vary depending on the input and history of previous events. This can allow for more expressive intensities. For instance, dynamically adjusting the real components of $\bm{\Lambda}$ to be smaller will result in more influence from history. Alternatively, larger values will result in more quickly “forgetting” the influence of previous events for a given hidden state channel. This is formalized with the following recurrence relation:
$\mathrm{d}\tilde{\mathbf{x}}_{t}$ $:=\bm{\Lambda}_{i}\tilde{\mathbf{x}}_{t-}\mathrm{d}t+\bm{\Lambda}_{i}\tilde{\mathbf{B}}\mathbf{u}_{t-}\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (17)
for $t\in(t_{i},t_{i+1}]$ where $\bm{\Lambda}_{i}:=\text{diag}\left(\text{softplus}(\mathbf{W}^{\prime}\mathbf{u}_{t_{i}}+\mathbf{b}^{\prime})\right)\bm{\Lambda}$ with $\mathbf{W}^{\prime}\in\mathbb{R}^{P\times H}$ and $\mathbf{b}^{\prime}\in\mathbb{R}^{P}$. Note that this is still conditionally linear in time as even though $\bm{\Lambda}_{i}$ changes it is entirely input-dependent based on $\mathbf{u}$ and not dependent on previous values of $\mathbf{x}$.
3.5 Computing LLH Recurrence
Thus far, we have created the LLH layer by diagonalizing and discretizing a modified SSM. As discussed earlier, we would like to take advantage of the efficient parallel scans leveraged by many SSM-based models (Smith et al., 2022; Gu & Dao, 2023; Dao & Gu, 2024). Below we explain how we can still use the parallel scan, despite the modified recurrence.
Parallel scans admit efficient inference over linear recurrences of the form $\mathbf{z}_{i+1}=\mathbf{A}_{i}\mathbf{z}_{i}+\mathbf{b}_{i}$ (Blelloch, 1990). Although we have added an impulse to the recurrence, this is still intrinsically of this form, where $\mathbf{z}_{i}:=\mathbf{x}_{t_{i}}$, $\mathbf{A}_{i}:=\exp(\bm{\Lambda}_{i}(t_{i+1}-t_{i}))$, and $\mathbf{b}_{i}:=(\mathbf{A}_{i}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t_{i+1}-}+\tilde{\mathbf{E}}\bm{\alpha}_{k_{i+1}}$. As a result, we can leverage efficient parallel scans to compute the sequence of right-limits $\mathbf{x}_{t_{1:N}}$ in parallel across the sequence length. The corresponding left-limits $\mathbf{x}_{t_{1:N}-}$ can then be efficiently computed after by subtracting off $\tilde{\mathbf{E}}\bm{\alpha}_{k_{1:N}}$ from $\mathbf{x}_{t_{1:N}}$. In Algorithms 1, 2 and 3 we compactly detail how to use a parallel scan to compute the sequence of right limits given events; how to evolve those right limits to compute left limits; and then how to subsequently compute the log-likelihood of the sequence.
4 Related Works
Neural MTPPs Marked temporal point processes (MTPPs) are generative models that jointly model the time and type of continuous-time sequential events, typically characterized by mark-specific intensity functions (Daley & Vere-Jones, 2003). Early approaches, such as self-exciting Hawkes processes (Hawkes, 1971; Liniger, 2009), used simple parametric forms for the intensity. More recently, neural architectures such as RNNs (Du et al., 2016; Mei & Eisner, 2017), CNNs (Zhuzhel et al., 2023), and transformers (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022) have been used to more flexibly model the conditional intensity. For intensity-free MTPPs, approaches include normalizing flows (Shchur et al., 2020a; Zagatti et al., 2024), neural processes (Bae et al., 2023), and diffusion models (Zeng et al., 2023; Zhang et al., 2024); however, the most common approach is to model intensities as it requires fewer modeling restrictions.
Efficient MTPPs Due to their recurrent nature, RNN-based MTPP models incur $\mathcal{O}(N)$ complexity for sequences of length $N$ as events must be processed sequentially. Attention-based MTPP models can be applied in parallel across the sequence, but the computational work scales as $\mathcal{O}(N^{2})$. Türkmen et al. (2020) proposed modeling events as conditionally independent so long as they occurred within the same time bin of a specified size. This resulted in parallel computation within bins, but still scales overall as $\mathcal{O}(N)$. Shchur et al. (2020b) proposed an intensity-free TPP which uses triangular maps and the time-change theorem (Daley & Vere-Jones, 2003). This was extended by Zagatti et al. (2024) to handle marks, but in doing so, lost many of the benefits of the original model and scales linearly in the mark dimension—which can rapidly become untenable with $O(NK)$ work. To the best of our knowledge, our proposed model is the first that efficiently scales with sequence length and mark space, as well as the first to fully leverage SSMs and parallel scans.
SSMs for Sequential Modeling SSMs have found recent success as alternatives to RNNs, CNNs, and transformers, enjoying reduced training cost and comparable modelling power (Gu et al., 2022b). A range of variants have been developed (Gu et al., 2021; Gupta et al., 2022; Gu et al., 2022a; Smith et al., 2022), and have been applied in language modeling (Gu & Dao, 2023), speech (Goel et al., 2022), and vision (Wang et al., 2023; Zhu et al., 2024). The linear recurrence allows for parallelism, as well as accessible long contexts which would be prohibitive for transformers due to their quadratic scaling. However, SSMs have not previously been applied to MTPPs, in part due to the irregular inter-event times and the input being a stochastic counting process.
Concurrent work by Gao et al. (2024) used Mamba (Gu & Dao, 2023), a recent deep SSM architecture, in an MTPP setting, which they call the Mamda Hawkes Process (MHP). The MHP uses a mamba SSM as the encoder in an encoder-decoder architecture, also leveraging the variable interval capabilities. Crucially, however, they use a separate parametric decoder for intermediate intensities (similar to, for instance, the THP). This misses the opportunity to “fully” leverage the SSM architecture, re-using the same variable interval evaluation to evaluate the the intensity.
5 Experiments
We now evaluate our deep linear Hawkes process model. Our core objectives in using SSMs for MTPP modeling were to define an architecture that is both (a) highly performant in its forecasting ability, and (b) able to leverage efficient parallel compute methods to accelerate inference. To this end, we first present a simple exploration of the ability of different models to represent parametric intensity functions. Then we present the main experiments in this paper, where we evaluate our model against a suite of common MTPP models on a range of datasets of different sizes. We conclude by testing the runtime of our model against a variety of baselines. We find that DLHP systematically outperforms baseline methods both in terms of log-likelihood on held-out test data and runtime across a range of sequence lengths. More results and details are included in Appendices C and D.
Metrics
[…]
Models
We compare our model (DLHP) with six of the most common MTPP models: two RNN-based models (RMTPP (Du et al., 2016), NHP (Mei & Eisner, 2017)), three transformer/attention-based models (THP (Zuo et al., 2020), SAHP (Zhang et al., 2020), AttNHP (Yang et al., 2022)), and one intensity-free model (IFTPP (Shchur et al., 2020a)). In all real-world experiments, extensive grid searches were conducted for hyperparameter tuning with configurations chosen based on validation log-likelihood, with the search spaces designed to roughly control parameter counts across models. Specifics for training, hyperparameters, and architectures are given in Appendix C.
Table 1: Per event log-likelihood ($\uparrow$ is better) results on the held-out test set averaged over five seeds (standard deviations in parentheses); OOM indicates insufficient memory. We bold the best and underline the runner-up per dataset. We also report the mean rank of models across datasets as a summary metric ($\downarrow$ is better). DLHP is consistently the best or second best-performing model. Extended results and discussion are presented in Section D.1.
Model Per Event Log-Likelihood, $\mathcal{L}_{\text{Total}}$ (nats)($\uparrow$) Avg. Ranking ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP -2.136 (0.003) -7.098 (0.217) 0.346 (0.002) 1.003 (0.004) -2.480 (0.019) -1.780 (0.005) -0.472 (0.026) -8.081 (0.025) 6.3
NHP 0.129 (0.012) -6.348 (0.000) 0.514 (0.004) 1.157 (0.004) -2.241 (0.002) -0.574 (0.011) 0.060 (0.017) -3.966 (0.058) 3.0
SAHP -2.074 (0.029) -6.708 (0.029) 0.298 (0.057) 1.168 (0.028) -2.341 (0.058) -1.646 (0.083) -0.677 (0.072) -6.804 (0.126) 5.4
THP -2.096 (0.002) -6.659 (0.007) 0.372 (0.002) 0.790 (0.002) -2.338 (0.014) -1.712 (0.011) -0.577 (0.011) -7.208 (0.096) 5.5
AttNHP 0.484 (0.077) -6.499 (0.028) 0.493 (0.009) 1.259 (0.022) -2.194 (0.016) -0.592 (0.051) -0.170 (0.077) OOM 3.1
IFTPP 0.496 (0.002) -10.344 (0.016) 0.453 (0.002) 1.318 (0.017) -2.233 (0.009) -0.492 (0.017) 0.317 (0.052) -6.596 (0.240) 2.9
DLHP (Ours) 0.781 (0.011) -6.365 (0.003) 0.522 (0.004) 1.304 (0.039) -2.163 (0.009) -0.557 (0.046) 1.243 (0.083) -2.512 (0.369) 1.4
Libraries and Compute Environment
We implement our DLHP in the EasyTPP library (Xue et al., 2023) and use their implementations of the baseline models. We also use the five standard datasets that EasyTPP immediately supports (see Section C.2 for more details). We then further include three larger datasets to stress-test the MTPP models (see Section 5.2). Unless otherwise stated, all models were trained using a single NVIDIA A5000 GPU with 24GB of onboard memory.
5.1 Synthetic Experiments
We start by performing a simple investigation into the expressivity of the DLHP intensity function and the ability to capture background intensities. We train our model and baselines on 5,000 sequences over the time period $[0,7.5]$ drawn from an unmarked, inhomogeneous Poisson process with a square-wave intensity function, $\lambda_{t}:=\mathds{1}(t\in(1,2)\cup(3,4)\cup(5,6))$ (see Fig. 1). We plot the estimated intensity functions conditioned on no events occurring, i.e., $\mathcal{H}_{t}:=\emptyset\;\forall t$.
Intensity estimates are shown for NHP and DLHP specifically in Fig. 1 (and for all models in Fig. 6 in Section D.2). We can see that our model successfully captures the true, underlying background intensity process almost perfectly. This is largely attributed to the expressivity of the linear recurrences and non-linear depth of the model. Other models have various failure modes: struggling to capture the multi-modality (RMTPP, NHP, SAHP, and THP), not matching the square shape (previous four and IFTPP), or not being able to stop the pattern from repeating a fourth time (AttNHP).
We also examine the performance of our DLHP on randomly instantiated parametric Hawkes processes, finding that DLHP is able to successfully recover the ground truth intensity. Furthermore, in this setting DLHP achieves the best held-out log-likelihood scores and competitive next-event time prediction (see Section D.3 for more details). These simple experiments confirm that the DLHP is sufficiently expressive to be able to represent more complicated intensity functions while other methods break down.
5.2 Log-likelihood Results on Real-world Datasets
We empirically investigate the performance of our proposed model against baseline methods by comparing the held-out log-likelihood per event. We evaluate our model on eight real-world datasets. Five of which are taken directly from EasyTPP (Xue et al., 2023). We also include two MTPP datasets that have been widely used throughout the literature: Last.fm, which includes data on users’ music listening patterns from Celma Herrada et al. (2009), and MIMIC-II, a subset of de-identified patient hospital visits processed from (Saeed et al., 2002). Finally, we introduce a third dataset from the recently released, publicly available electronic health record (EHR) dataset EHRSHOT (Wornow et al., 2023). To construct the dataset, we first establish the most used Current Procedural Terminology (CPT-4) codes that identify medical services and procedures as events. The processed dataset comprises sequences of CPT-4 codes issued to individual patients during their care. This dataset has a maximum sequence length 10$\times$ longer than the longest in the EasyTPP datasets (and 100$\times$ that of MIMIC-II), providing a challenging testbed (in terms of scale) beyond existing datasets. Data statistics and other details including pre-processing are provided in Table 6 and Section C.3.
Figure 3: Per event log-likelihood on the held-out test data in Table 1 decomposed into time and mark components (i.e., $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{Time}}+\mathcal{L}_{\text{Mark}}$). Models are ordered by their average ranking. Model results are adjusted by subtracting the log-likelihood achieved by RMTPP for readability.
From the held-out log-likelihood results shown in Table 1, DLHP consistently achieves the best or the second-best log-likelihood across all datasets. Compared to the best baseline model per-dataset, DLHP produces a (geometric) mean likelihood ratio of 1.38 (corresponding to 38% higher likelihood on true events). We decompose this improvements in Fig. 3, where we see the improvements in log-likelihood are mainly driven by better modeling of time. Extended plots included in Section D.1. Given the clear improvement in temporal modeling, we posit that DLHPs are particularly well suited in applications that contain more complex patterns over time. All of these results for DLHP utilize input-dependent dynamics (see Section 3.4). This was found to reliably improve forecasting performance in ablation studies (see Section D.4).
We also report and discuss additional metrics in the Appendix. We report next-mark classification accuracy and RMSE of next-event arrival time, finding that DLHP matches or outperforms all baselines. We also report model calibration with respect to next event time and mark prediction. Calibration aims to grade the predictive uncertainty of the model (Bosser & Taieb, 2023), which is not captured by other metrics such as mark classification accuracy and time RMSE. On the whole, our model (as well as the baselines) tend to produce well-calibrated time and mark predictions across the datasets. We also include full tables for the likelihood decomposition in Table 7.
5.3 Speed Testing
A key motivation for DLHP is to leverage the properties of SSMs to improve scaling and accelerate inference. To test this, we measure the wallclock time for a full forward pass and log-likelihood evaluation on random input sequences with lengths ranging from eight events to over half a million events. The architectures and mark spaces are the same as in the StackOverflow experiments (see Tables 5(a) and 6). Our EasyTPP PyTorch DLHP is written in pure PyTorch, and hence is not as optimized as other methods (compared to, for instance, IFTPP, which uses a GPU-optimized implementation of the GRU). We therefore include the runtimes of a standalone JAX DLHP implementation, which allows for similar optimization through JIT compilation.
Figure 4: Median runtime, over 20 random seeds, for various models against increasing sequence lengths. We show runtimes for both conditioning on a sequence (Algorithm 1) and likelihood evaluation (Algorithm 3). We see that DLHP is faster across a wide range of sequence lengths. Variance of runtime is negligble. Crosses indicate where the THP runs out of memory or the IFTPP throws an error.
The THP scales well before reverting to superlinear scaling, and then runs out of memory. The IFTPP is very fast at shorter runtimes, but quickly reverts to linear scaling, due to its simple but highly optimized implementation and inherently sequential operation. Both DLHP implementations scale linearly at long sequence lengths, but have near-constant runtime at shorter sequences. At shorter sequence lengths, the more optimized JAX implementation is faster than the unoptimized pure PyTorch implementation. While this indicates that there are additional opportunities to accelerate the PyTorch implementation further (e.g., exploiting kernel fusion or writing a lower-level Triton implementation), these results still confirm our core hypothesis: that our DLHP can exploit parallel scans to scale to long sequences more effectively than other methods.
6 Conclusion
We present the deep linear Hawkes process (DLHP)—a novel combination of ideas from LHPs and SSMs. Our DLHP leverages the unique properties of deep SSM architectures to achieve a flexible and performant model, without additional and restrictive intensity decoding heads. We then demonstrated that our DLHP outperforms existing methods across a range of standard and new benchmark tasks over various metrics, such as log-likelihood and runtime across sequence lengths. One limitation of our method is the increased complexity of the implementation (as we require a parallel scan), compared to, for instance, the NHP (which only requires a basic for loop). Following from this, a second limitation is that we have lost most of the interpretability of the latent dynamics and parameters enjoyed by the LHP. While there is some limited interpretability from the mark-specific impulses due to each LLH layer incrementing the residual stream of the model, this is largely correlative and imprecise. Future research directions therefore include improving on and strengthening these aspects, as well as developing additional theory around the use of deep SSMs in this novel MTPP setting, and developing heuristics and best-practices for configuring hyperparameters in various modeling settings. Specifically, exploring the relative benefits of the forward and backward discretization is a unique research direction arising from the DLHP. However, we believe the robustness, predictive performance, computational efficiency, and extensibility of DLHPs make them a very competitive model out-of-the-box for a wide range of applications.
Supplementary Materials:
Deep Linear Hawkes Processes
Appendix A Acronyms and Notation
Table 2: Key notation used repeatedly across this paper.
Symbol Space Description
$t$ $\mathbb{R}_{\geq 0}$ Time
$T$ $\mathbb{R}_{\geq 0}$ Maximum time in a given sequence’s observation window
$t_{i}$ $\mathbb{R}_{\geq 0}$ $i^{\mathrm{th}}$ time
$t-$ $\mathbb{R}_{\geq 0}$ Subscript minus indicates left-limit
$t+$ $\mathbb{R}_{\geq 0}$ Subscript plus indicates right-limit
$k$ =$\mathcal{M}=\left\{1,\ldots,K\right\}$ Event mark
$\mathcal{H}$ $\mathcal{M}^{N}\times\mathbb{R}_{\geq 0}^{N}$ Event history for $N$ events
$\mathbf{N}_{t}$ $\mathbb{Z}_{\geq 0}^{K}$ Counting process for $K$ marks at time $t$
$\lambda^{k}_{t}$ $\mathbb{R}_{\geq 0}$ Intensity of $k^{\mathrm{th}}$ mark type at time $t$
$\bm{\lambda}_{t}$ $\mathbb{R}_{\geq 0}^{K}$ Vector of $K$ mark intensities at time $t$
$\lambda_{t}$ $\mathbb{R}_{\geq 0}$ Ground/total intensity (sum of mark-specific intensities)
[…]
$\nu^{\mathrm{k}}$ $\mathbb{R}_{\geq 0}$ Background intensity for the $k^{\mathrm{th}}$ mark
$\bm{\alpha}$ $\mathbb{R}^{K,K}_{\geq 0}$ (For LHP) Matrix of intensity impulses from each type of mark
$\bm{\beta}$ $\mathbb{R}^{K,K}_{\geq 0}$ (For LHP) Dynamics matrix of intensity evolution
[…]
$\mathbf{x}_{t}$ $\mathbb{R}^{P}$ LLH/SSM hidden state at time $t$
$\mathbf{x}_{0}$ $\mathbb{R}^{P}$ Learned LLH/SSM initial hidden state
[…]
$\mathbf{y}_{t}$ $\mathbb{R}^{H}$ LLH/SSM output at time $t$
$\mathbf{u}_{t}$ $\mathbb{R}^{H}$ LLH/SSM input at time $t$
$\mathbf{A}$ $\mathbb{R}^{P\times P}$ LLH/SSM transition matrix
$\mathbf{B}$ $\mathbb{R}^{P\times H}$ LLH/SSM input matrix
$\mathbf{C}$ $\mathbb{R}^{H\times P}$ LLH/SSM output matrix
$\mathbf{D}$ $\mathbb{R}^{H\times H}$ LLH/SSM passthrough matrix
$\mathbf{E}$ $\mathbb{R}^{P\times R}$ LLH mark embedding matrix ($P\times R$ in low-rank factorization)
[…]
$\bm{\alpha}$ $\mathbb{R}^{R\times K}$ (For DLHP) Mark impulses ($R\times K$ in low-rank factorization)
$\sim$ N/A Tilde (e.g., $\tilde{\mathbf{B}}$) denotes variable is in the diagonalized eigenbasis
$\Lambda$ $\mathbb{C}^{P\times P}$ Matrix of eigenvalues of $\mathbf{A}$; diagonalized dynamics matrix
$\bar{\Lambda}$ $\mathbb{C}^{P\times P}$ Discretized diagonal dynamics matrix
${(l)}$ N/A Superscript index in parenthesis indicates layer (i.e., $\mathbf{x}$ for layer $l$)
Table 3: Key acronyms used throughout this paper.
Acronym Page number Definition
CNN 6 Convolutional neural network
LHP 1 Linear Hawkes process
LLH 2 Latent linear Hawkes
MTPP 1 Marked temporal point process
RNN 1 Recurrent neural network
SSM 1 (Deep) State-space model
TPP 7 Temporal point process
ZOH 5 Zero-order hold
RMTPP 7 Recurrent marked temporal point process (Du et al., 2016)
NHP 1 Neural Hawkes process (Mei & Eisner, 2017)
SAHP 7 Self-attentive Hawkes process (Zhang et al., 2020)
THP 7 Transformer Hawkes process (Zuo et al., 2020)
AttNHP 7 Attentive neural Hawkes process (Yang et al., 2022)
IFTPP 7 Intensity-free temporal point process (Shchur et al., 2020a)
DLHP 1 Deep linear Hawkes process (ours)
Appendix B Additional Details on Methods
B.1 DLHP Algorithms
Algorithm 1 Deep Linear Hawkes Process: Get Right State Limits
1:DLHP layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, event intervals $\Delta t_{1:N}$, nonlinearity $\sigma$, shared mark embeddings $\bm{\alpha}_{1:N}$.
2:Right state limits $\mathbf{x}_{t_{1:N}}^{(1:L)}$
3:$\mathbf{u}_{t_{1:N}-}=\mathbf{0}$triangleright Left input limits
4:for $l$ in $1:L$ do
5:     $\bar{\bm{\Lambda}}^{(l)}_{1:N}=\mathrm{Discretize}\left(\bm{\Lambda}^{(l)},\Delta t_{1:N}\right)$triangleright Zero-order hold, see Eq. 22
6:     $\tilde{\mathbf{x}}^{(l)}_{t_{1:N}}=\mathrm{ParallelScan}\left(\bar{\bm{\Lambda}}^{(l)}_{1:N},(\bar{\bm{\Lambda}}^{(l)}_{1:N}-\mathbf{I})\tilde{\mathbf{B}}^{(l)}\mathbf{u}_{t_{1:N}-}+\tilde{\mathbf{E}}^{(l)}\bm{\alpha}_{1:N}\right)$triangleright Compute right $x$ limits
7:     $\tilde{\mathbf{x}}^{(l)}_{t_{1:N}-}=\tilde{\mathbf{x}}^{(l)}_{t_{1:N}}-\tilde{\mathbf{E}}^{(l)}\bm{\alpha}_{1:N}$triangleright Compute left $x$ limits
8:     $\mathbf{u}_{t_{1:N}-}=\mathrm{LayerNorm}\left(\sigma\left(\tilde{\mathbf{C}}^{(l)}\tilde{\mathbf{x}}_{t_{1:N}-}+\mathbf{D}^{(l)}\mathbf{u}_{t_{1:N}-}\right)+\mathbf{u}_{t_{1:N}-}\right)$triangleright Compute next layer’s left $u$ limits
9:end for
10:return $\mathbf{x}_{t_{1:N}}^{(1:L)}$
Algorithm 2 Deep Linear Hawkes Process: Get Intensity From Right Limit
1:DLHP layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, Previous state right limits $\mathbf{x}_{t}^{(1:L)}$, Integration period $\delta t$, nonlinearity $\sigma$, Intensity function $\mathrm{IntensityFn}$.
2:Intensity left limit $\bm{\lambda}_{t+\delta t}$
3:$\mathbf{u}_{t+\delta t-}=\mathbf{0}$triangleright Left input limit
4:for $l$ in $1:L$ do
5:     $\bar{\bm{\Lambda}}^{(l)}=\mathrm{Discretize}\left(\bm{\Lambda}^{(l)},\delta t\right)$triangleright Zero-order hold, see Eq. 22
6:     $\tilde{\mathbf{x}}^{(l)}_{t+\delta t-}=\bar{\bm{\Lambda}}^{(l)}\mathbf{x}_{t}^{(l)}+(\bar{\bm{\Lambda}}^{(l)}-\mathbf{I})\tilde{\mathbf{B}}^{(l)}\mathbf{u}_{t+\delta t-}$triangleright Evolve state
7:     $\mathbf{u}_{t+\delta t-}=\mathrm{LayerNorm}\left(\sigma\left(\tilde{\mathbf{C}}^{(l)}\tilde{\mathbf{x}}_{t+\delta t-}^{(l)}+\mathbf{D}^{(l)}\mathbf{u}_{t+\delta t-}\right)+\mathbf{u}_{t+\delta t-}\right)$triangleright Compute event left $u$ limits
8:end for
9:$\bm{\lambda}_{t+\delta t}=\mathrm{IntensityFn}(\mathbf{u}_{t+\delta t-})$triangleright Rectify intensity, see Eq. 13
10:return $\bm{\lambda}_{t+\delta t}$
Algorithm 3 Deep Linear Hawkes Process: Compute Log-Likelihood
1:DLHP layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, Event times $t_{1:N}$, mark types $k_{1:N}$, nonlinearity $\sigma$, shared mark embedding function $\mathrm{EmbedMarks}$, number of integration points per event $M$, Intensity function $\mathrm{IntensityFn}$.
2:Log-ikelihood $\mathcal{L}$
[…]
4:$t_{0}:=0$
5:$\Delta t_{1:N}=t_{1:N}-t_{0:N-1}$
6:$s_{1:N,1:M}\sim\mathcal{U}(0,\Delta t_{1:N})$triangleright Sample $M$ integration points per interval (non-inclusive)
7:$\tilde{\mathbf{x}}_{t_{1:N}}^{(1:L)}=\mathrm{GetRightStateLimits}(\Theta,\Delta t_{1:N},\sigma,\bm{\alpha}_{1:N})$triangleright Algorithm 1, $\mathcal{O}(\log N)$ parallel time
8:for $n$ in $1:N$ do $\triangleright$ This is embarrassingly parallelizable with vmap, $\mathcal{O}(1)$ parallel time
9:     $\bm{\lambda}_{t_{n}}=\mathrm{GetIntensityFromRightLimit}\left(\Theta,\tilde{\mathbf{x}}_{t_{n}}^{(1:L)},\Delta t_{n},\sigma,\mathrm{IntensityFn}\right)$triangleright Algorithm 2, $\mathcal{O}(1)$ parallel time
10:     for $m$ in $1:M$ do $\triangleright$ This is embarrassingly parallelizable with vmap, $\mathcal{O}(1)$ parallel time
11:         $\bm{\lambda}_{s_{n,m}}=\mathrm{GetIntensityFromRightLimit}\left(\Theta,\tilde{\mathbf{x}}_{t_{n}}^{(1:L)},s_{n,m},\sigma,\mathrm{IntensityFn}\right)$triangleright Algorithm 2, $\mathcal{O}(1)$ parallel time
12:     end for
13:end for
14:$\mathcal{L}=\sum_{n=1}^{N}\log\lambda_{t_{n}}^{k_{n}}+\sum_{n=1}^{N}\frac{\Delta t_{n}}{M}\sum_{m=1}^{M}\sum_{k=1}^{K}\lambda^{k}_{s_{n,m}}$triangleright Eq. 2 with Monte-Carlo approximation of integral
15:return $\mathcal{L}$
B.2 Discretization and Zero Order Hold
The linear recurrence is defined in continuous-time. This mirrors the (M)TPP setting, where event times are not on a fixed intervals. We use the zero-order hold (ZOH) discretization method, to convert the continuous-time linear recurrence into a sequence of closed-form updates, given the integration times, that can also be efficiently computed. We refer the reader to Iserles (2009) for a comprehensive introduction to the ZOH transform.
The main assumption of the ZOH discretization is that the input signal is held constant over the time period being integrated. Under this assumption, it is possible to solve for the dynamics and input matrices that yield the correct state at the end of the integration period. For the LLH dynamics in Eq. 10, when no events occur in $(t,t^{\prime})$, this becomes
$\mathbf{x}_{t^{\prime}-}=\int_{t}^{t^{\prime}}\mathbf{A}\mathbf{x}_{t}+\mathbf{AB}\mathbf{u}_{t}\mathrm{d}t=\overline{\mathbf{A}}\mathbf{x}_{t}+\overline{\mathbf{AB}}\mathbf{u}_{t}\quad\mathrm{assuming}\quad\mathrm{d}\mathbf{u}_{t}=\mathbf{0}\in[t,t^{\prime}],$ (18)
where the resulting discretized matrices are
$\overline{\mathbf{A}}=e^{\mathbf{A}\Delta t},\quad\overline{\mathbf{AB}}=\mathbf{A}^{-1}(e^{\mathbf{A}\Delta t}-\mathbf{I})\mathbf{AB},\quad\mathrm{where}\quad\Delta t=t^{\prime}-t.$ (19)
The ZOH does not affect the output or passthrough matrices $\mathbf{C}$ and $\mathbf{D}$. To compute the matrices $\overline{\mathbf{A}}$ and $\overline{\mathbf{AB}}$ however requires computing a matrix exponential and a matrix inverse. However, Smith et al. (2022) avoid this by diagonalizing the system (also avoiding a dense matrix-matrix multiplication in the parallel scan). The diagonalized dynamics and input matrices are denoted $\bm{\Lambda}$ (a diagonal matrix) and $\bm{\Lambda}\tilde{\mathbf{B}}$ respectively. In this case, Eq. 19 reduces to
$\overline{\mathbf{A}}$ $=e^{\bm{\Lambda}\Delta t},$ (20)
$\overline{\mathbf{AB}}$ $=\bm{\Lambda}^{-1}(e^{\bm{\Lambda}\Delta t}-\mathbf{I})\bm{\Lambda}\tilde{\mathbf{B}}$ (21)
[…]
where $e^{\bm{\Lambda}\Delta t}$ is trivially computable as the exponential of the leading diagonal of $\bm{\Lambda}\Delta t$. These operations are embarrassingly parallelizable across the sequence length and state dimension given the desired evaluation times.
To contextualize, suppose an event occurs at time $t$, Eq. 22 allows us to exactly (under the constant-input assumption) efficiently evaluate the linear recurrence at subsequent times $t^{\prime}$. We use this extensively in the DLHP to efficiently evaluate the recurrence (and hence the intensity) at the irregularly-spaced event times and times used to compute the integral term.
It should be noted the discretization was done to compute a left-limit $\mathbf{x}_{t^{\prime}-}$ from a previous right-limit $\mathbf{x}_{t}$. Should an event not occur at $t^{\prime}$, then the left- and right-limits agree and $\mathbf{x}_{t^{\prime}-}=\mathbf{x}_{t^{\prime}+}=\mathbf{x}_{t^{\prime}}$. If an event does occur at time $t^{\prime}$ with mark $k$, then the left-limit $\mathbf{x}_{t^{\prime}-}$ can be incremented by $\tilde{\mathbf{E}}\bm{\alpha}_{k}$ to compute $\mathbf{x}_{t^{\prime}+}=\mathbf{x}_{t^{\prime}}$. This increment from left- to right-limit is exact and leverages no discretization assumption.
B.3 Interpretation for Input-Dependent Dynamics
Consider the input-dependent recurrence for an LLH layer, as defined in Eq. 17:
$\mathrm{d}\tilde{\mathbf{x}}_{t}$ $:=\bm{\Lambda}_{i}\tilde{\mathbf{x}}_{t-}\mathrm{d}t+\bm{\Lambda}_{i}\tilde{\mathbf{B}}\mathbf{u}_{t-}\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (23)
for $t\in(t_{i},t_{i+1}]$ where $\bm{\Lambda}_{i}:=\text{diag}(\Delta_{i})\bm{\Lambda}$ with the input-dependent factor defined as $\Delta_{i}:=\text{softplus}(\mathbf{W}^{\prime}\mathbf{u}_{t_{i}}+\mathbf{b}^{\prime})\in\mathbb{R}^{P}_{>0}$. This factor can be thought of as the input-dependent relative-time scale for the dynamics. To see this, we first note that for vectors $\mathbf{p,q}\in\mathbb{R}^{d}$, the following holds true: $\text{diag}(\mathbf{p})\mathbf{q}=\mathbf{p}\odot\mathbf{q}=\mathbf{q}\odot\mathbf{p}$ where $\odot$ is the Hadamard or element-wise product. It then follows that
$\mathrm{d}\tilde{\mathbf{x}}_{t}$ $:=\bm{\Lambda}_{i}\tilde{\mathbf{x}}_{t-}\mathrm{d}t+\bm{\Lambda}_{i}\tilde{\mathbf{B}}\mathbf{u}_{t-}\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (24)
$=\bm{\Lambda}_{i}(\tilde{\mathbf{x}}_{t-}+\tilde{\mathbf{B}}\mathbf{u}_{t-})\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (25)
$=\text{diag}(\Delta_{i})\bm{\Lambda}(\tilde{\mathbf{x}}_{t-}+\tilde{\mathbf{B}}\mathbf{u}_{t-})\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (26)
$=[\bm{\Lambda}(\tilde{\mathbf{x}}_{t-}+\tilde{\mathbf{B}}\mathbf{u}_{t-})]\odot(\Delta_{i}\mathrm{d}t)+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}.$ (27)
As shown, the positive vector $\Delta_{i}$ can be thought of as changing the relative time-scale for each channel in the hidden state $\tilde{\mathbf{x}}$. Large values of $\Delta_{i}$ will act as if time is passing quickly, encouraging the state to converge to the steady-state sooner. Conversely, smaller values will make time pass more slowly causing the model to retain the influence that prior events have on future ones (for that specific channel in $\tilde{\mathbf{x}}$ at least).
B.4 Forwards and Backwards Zero Order Hold Discretization
In Section 3.3 we highlighted that the ZOH discretization is exact when $\mathbf{u}_{t}$ is held constant over the integration window. This raises a unique design question for DLHPs: what constant value should $\mathbf{u}_{t}$ take on when evolving $\mathbf{x}$ from time $t$ to $t^{\prime}$? For the first layer of the model, the input is zero by construction, so there is no choice to be made—in fact, since $\mathbf{u}$ is constant for the first layer the updates are exact. However, the input is non-zero at deeper layers, and, crucially, varies over the integration period.
We must therefore decide how to select a $\mathbf{u}$ value over the integration period. This should be a value in (or function of) $\{\mathbf{u}_{s}\;|\;s\in[t,t^{\prime})\}$. Note this is because the value at $t^{\prime}$, $\mathbf{u}_{t^{\prime}}$, cannot be incorporated as this would cause a data leakage in our model; while values prior to $t$ would discard the most recent mark. For this work, we explore two natural choices: (i) the input value at the beginning of the interval, $\mathbf{u}_{t}$, and (ii) the left-limit at the end of the interval, $\mathbf{u}_{t^{\prime}-}$. We illustrate the backwards variant in Fig. 2, where in the rightmost panel, we use the $\mathbf{u}_{t^{*}}$ values at each layer, as opposed to $\mathbf{u}_{t_{3}}$. We refer to these options as forwards and backwards ZOH, respectively. All experiments in the main paper utilize backwards ZOH.
It is not obvious a priori which one of these modes is more performant. We therefore conducted an ablation experiment in Table 13. We see that there is little difference between the two methods. We also note that models are learned through this discretization, and so this decision does not mean that a model is “incorrectly discretized” one way or the other, but instead they define subtlety different families of models. Theoretical and empirical investigation of the interpretations of this choice is an interesting area of investigation going forwards, extending the ablations we present in Table 13.
B.5 Theoretical Complexity
We include in Table 4 a brief summary of the theoretical complexity of each of the methods we consider. We break these down by the work, memory complexity and theoretical best parallel application time of the forward pass (used when conditioning on a sequence, the left-hand term of Eq. 2) and evaluating the integral term in Eq. 2 given that the forward pass has been completed (as this is either required by the method, and is nearly always evaluated in conjunction with the forward pass). We then state the limiting best-case theoretical parallelism of the two components.
The reasoning behind this is as as follows:
  • The forward pass of RMTPP, NHP and IFTPP use non-linear RNNs, and hence incur memory and work that is linear in the sequence length, and cannot be parallelized. However, they re-use the computed hidden states to compute the integral term, and hence, while they incur work and memory that scales in the sequence length and number of events, this work can be perfectly parallelized. This results in a best-case parallelism of $\mathcal{O}(L)$ (dominated by the forward pass).
  • SAHP, THP and AttNHP all use self-attention, and hence have a work and memory that scales quadratically in the sequence length, although this work can be parallelized across the sequence length, resulting in logarithmic parallel depth. SAHP and THP re-use embeddings and a parametric decoder, and hence estimating the integral scales like the RNN, and hence the limiting parallelism is still the forward pass. AttNHP is slightly different in that it re-applies the whole independently attention mechanism for each integration point. However, this work is parallelizeable and hence still reduces to a best-case depth of $\mathcal{O}(\log L)$.
  • DLHP is an RNN and hence has linear work and memory in the forward pass, but can be parallelized to a best-case depth of $\mathcal{O}(\log L)$ using the parallel scan. We then re-use the states computed in the forward pass for estimating the integral, which, as with the other RNN methods, is perfectly parallelizable, resulting in a theoretical parallel depth of $\mathcal{O}(\log L)$.
Note that these figures do not account for the number of layers required by each model, which must be evaluated in sequence.
Table 4: Comparison of methods based on memory and compute complexity. We see that our DLHP matches the best performing baseline in all categories. $L$ denotes to the sequence length, and $M$ denotes to the number of Monte Carlo grid points per-event used in evaluating Eq. 2. As IFTPP is an intensity-free method, it does not need to estimate $\int\lambda_{t}\mathrm{d}{}t$ as the other methods do.
Method Forward Pass Estimating $\int\lambda_{t}\mathrm{d}{}t$ Overall
Memory Work Theoretical Memory Work Theoretical Theoretical
Parallelism Parallelism Parallelism
RMTPP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(L)$
NHP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(L)$
SAHP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
THP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
AttNHP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(L^{2}M)$ $\mathcal{O}(L^{2}M)$ $\mathcal{O}(\log L)$ $\mathcal{O}(\log L)$
IFTPP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ N/A N/A N/A $\mathcal{O}(L)$
DLHP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
Appendix C Experimental Configurations and Datasets
C.1 Training Details & Hyperparameter Configurations
We apply a grid search for all models on all datasets for hyperparameter tuning. We use a default batch size of 256 for training. For models/datasets that require more memory (e.g., large mark space or long sequences), we reduce the batch size and keep them as consistent as possible among all the models on each dataset. We use the Adam stochastic gradient optimizer (Kingma & Ba, 2015), with a learning rate of 0.01 and a linear warm-up schedule over the first 1% iterations, followed by a cosine decay. Initial experiments showed this setting generally worked well across different models and datasets leads to convergence within 300 epochs. We also clip the gradient norm to have a max norm of 1 for training stability. We use Monte-Carlo samples to estimate the integral in log-likelihood, where we use 10 Monte-Carlo points per event during training.
On the five EasyTPP benchmark datasets and MIMIC-II that are smaller in their scales, we choose an extended grid based on the architecture reported in the EasyTPP paper. Specifically, we search over hidden states size $h$ = {16, 32, 64, 128, 256} for RMTPP, $h$ = {32, 64, 128} for NHP, and $h$ = {16, 32, 64} for IFTPP. For SAHP, THP, and AttNHP, we searched over all combinations of number of $L$ = {1, 2, 3}, hidden state size = {16, 32, 64, 128}, and number of heads = {1, 2, 4}. Finally, for DLHP, we considered combinations for number of layers = {1, 2, 3, 4}, $p$ = {16, 32, 64, 128} and $h$ = {16, 32, 64, 256}. We fixed the activation function as GeLU (Hendrycks & Gimpel, 2016) and apply post norm with layer norm (Ba, 2016). We fix the dropout as 0.1 for DLHP on the five core benchmark datasets, and add dropout = {0, 0.1} to the grid search for the other three datasets. Due to the scale of Last.fm and EHRSHOT datasets, we perform a smaller search over architectures that roughly match the parameter counts for all models at three levels: 25k, 50k, 200k, and choose the model with the best validation results. AttNHP has expensive memory requirements that tends to have smaller batch sizes than other models. We were unable to train any AttNHP on EHRSHOT. The final model architectures used are reported in Table 5(a) and Table 5(b). These configurations are also included in the supplementary code we include.
Table 5: Model architectures for the experiments presented in Table 1
(a) Model architectures for the five EasyTPP benchmark datasets.
Model Amazon Retweet Taxi Taobao StackOverflow
RMTPP $h=128$ $h=16$ $h=128$ $h=16$ $h=256$
NHP $h=128$ $h=64$ $h=128$ $h=128$ $h=64$
SAHP $h=32,l=2,\mathrm{heads}=2$ $h=32,l=3,\mathrm{heads}=4$ $h=16,l=2,\mathrm{heads}=4$ $h=32,l=1,\mathrm{heads}=1$ $h=64,l=1,\mathrm{heads}=1$
THP $h=32,l=2,\mathrm{heads}=4$ $h=16,l=3,\mathrm{heads}=4$ $h=128,l=1,\mathrm{heads}=4$ $h=64,l=1,\mathrm{heads}=1$ $h=16,l=2,\mathrm{heads}=4$
AttNHP $h=64,t=16,l=2,\mathrm{heads}=4$ $h=16,t=16,l=2,\mathrm{heads}=4$ $h=16,t=16,l=3,\mathrm{heads}=4$ $h=32,t=16,l=3,\mathrm{heads}=4$ $h=32,t=16,l=2,\mathrm{heads}=4$
IFTPP $h=64$ $h=64$ $h=32$ $h=64$ $h=64$
DLHP $h=64,p=128,l=2$ $h=128,p=128,l=2$ $h=128,p=16,l=4$ $h=32,p=16,l=4$ $h=32,p=32,l=3$
(b) Model architectures for the additional three benchmark datasets.
Model Last.fm MIMIC-II EHRSHOT
RMTPP $h=256$ $h=128$ $h=16$
NHP $h=112$ $h=128$ $h=80$
SAHP $h=136,l=2,\mathrm{heads}=4$ $h=64,l=2,\mathrm{heads}=4$ $h=8,l=2,\mathrm{heads}=4$
THP $h=48,l=2,\mathrm{heads}=4$ $h=32,l=3,\mathrm{heads}=4$ $h=32,l=2,\mathrm{heads}=4$
AttNHP $h=28,t=16,l=2,\mathrm{heads}=4$ $h=64,t=16,l=3,\mathrm{heads}=2$ OOM
IFTPP $h=48$ $h=256$ $h=16$
DLHP $h=68,p=16,l=2$ $h=256,p=64,l=2$ $h=128,p=32,l=2$
C.2 Dataset Statistics
We report the statistics of all eight datasets we used in Table 6. We used the HuggingFace version of the five EasyTPP datasets. For all datasets, we further ensure the MTPP modeling assumptions are satisfied that no more than two events occur at the same time (i.e., inter-arrival time is strictly positive), and event times do not lie on grid points that are effectively discrete-time events. Dataset descriptions and pre-processing details are provided in Section C.3.
Table 6: Statistics of the eight datasets we experiment with.
Dataset $K$ Number of Events Sequence Length Number of Sequences
Train Valid Test Min Max Mean Train Valid Test
Amazon 16 288,377 40,995 84,048 14 94 44.8 6,454 922 1,851
Retweet 3 2,176,116 215,521 218,465 50 264 108.8 20,000 2,000 2,000
Taxi 10 51,584 7,404 14,820 36 38 37.0 1,400 200 400
Taobao 17 73,483 11,472 28,455 28 64 56.7 1,300 200 500
StackOverflow 22 90,497 25,762 26,518 41 101 64.8 1,401 401 401
Last.fm 120 1,534,738 344,542 336,676 6 501 207.2 7,488 1,604 1,604
MIMIC-II 75 9,619 1,253 1,223 2 33 3.7 2600 325 325
EHRSHOT 668 759,141 165,237 170,147 5 3,955 177.0 4,329 927 927
C.3 Dataset Pre-processing
We use the default train/validation/test splits for EasyTPP benchmark datasets. For MIMIC-II, we copy Du et al. (2016) and keep the 325 test sequences in the test split, and further split the 2,935 training sequences into 2,600 for training and 325 for validation. In our pre-processed datasets, Last.fm and EHRSHOT, we randomly partition into subsets containing 70%, 15%, 15% of all sequences for training/validation/test respectively. We provide a high-level description of all the datasets we used, followed by our pre-processing procedure of Last.fm and EHRSHOT in more detail. Note that for datasets that contain concurrent events or effectively discrete times, we apply a small amount of jittering to ensure no modeling assumptions are violated in the MTPP framework.
Amazon (Ni et al., 2019) contains user product reviews where product categories are considered as marks. Retweet (Zhao et al., 2015) predicts the popularity of a retweet cascade, where the event type is decided by if the retweet comes from users with “small”, “medium”, or “large” influences, measured by number of followers (Mei & Eisner, 2017). Taxi data (Whong, 2014; Mei et al., 2019) uses data from the pickups and dropoffs of New York taxi and the marks are defined as the Cartesian product of five discrete locations and two actions (pickup/dropoff). Taobao (Xue et al., 2022) describes the viewing patterns of users on an e-commerce site, where item categories are considered as marks. StackOverflow contains the badges (defined as marks) awarded to users on a question-answering website. Finally, MIMIC-II (Saeed et al., 2002) records different diseases (used as marks) during hospital visits of patients. We add a small amount of noise to the MIMIC-II event times so that events do not lie on a fixed grid. Both StackOverflow and MIMIC-II datasets were first pre-processed by Du et al. (2016).
Last.fm Celma Herrada et al. (2009); McFee et al. (2012) records 992 users’ music listening habits that has been widely used in MTPP literature (Kumar et al., 2019; Boyd et al., 2020; Bosser & Taieb, 2023). Mark types are defined as the genres of a song, and each event is a play of a particular genre. Each sequence represents the monthly listening behavior of each user, with sequence lengths between 5 and 500. If the song is associated with multiple genres we select a random one of the genres, resulting in a total of 120 different marks.
EHRSHOT Wornow et al. (2023) is a newly proposed large dataset of longitudinal de-identified patient medical records, and has rich information such as hospital visits, procedures, and measurements. We introduce an MTPP dataset derived from EHRSHOT, where medical services and procedures are treated as marks, as identified by Current Procedural Terminology (CPT-4) codes. Each patient defines an event sequence, and we retain only CPT-4 codes with at least 100 occurrences in the dataset. For the $<1\%$ events of events where there are more than 10 codes at a single timestamp, we retain the top 10 codes with the most frequencies and discard the rest. We then add a small amount of random noise to the event time to ensure they are not overlapping. This process ensures we still satisfy the MTPP framework, and can reasonably instead compute top-10 accuracy for the next mark prediction. Other work has considered extending the MTPP framework to consider simultaneous event occurrence (Chang et al., 2024). Then we standardize each sequence to start and end with start and end of a sequence events. Note that we do not score these events. Event times are normalized to be in hours. We discard sequences that have less than 5 events and a single timestamp. This leads to the final version of our dataset to have 668 marks, and the sequence lengths range from 5 to 3955 events, reflecting patient histories that can span multiple years.
Appendix D Additional Experimental Results
D.1 Full Results on Benchmark Datasets
We provide the full log-likelihood results and corresponding plots in Table 7 and Fig. 5 respectively, where we decompose the likelihood into time and mark likelihoods. The improvement of our DLHP model is mainly driven by better modeling of time, though we also often obtain best- or second-best predictive performance on marks from the next event prediction accuracy results conditioned on true event time in Table 8. In aggregate, our model achieves a 1.38 per-event likelihood ratio between itself and the next best method across all datasets (a 38% improvement in likelihood). This is calculated by computing the mean log-likelihood ratio across all datasets and then exponentiating. Doing so is equivalent to taking the geometric mean across likelihood ratios.
We also include root mean square error (RMSE) in Table 9 as one of the commonly used metrics in MTPP literature, with the caveat that it purely describes how well the model captures the expected value of the next-event time. In other words, it is not informative when the inter-event times have a multimodal distribution – which we often encounter in practice with real data. Intensity-based methods were sampled using the thinning algorithm and averaged over 64 sampled candidate times when memory allows. For our method, we include a simple linear probe on the output hidden vector to directly estimate the next event time, which is substantially faster in wallclock runtime than sampling even when accounting for fine-tuning the linear probe parameters. For RMSE, IFTPP is the best performer overall; however, across intensity-based methods we find our model to be competitive. We posit this is due to the simplified parameterization of IFTPP lending itself well to capturing the expected time of the next event.
Table 7: Complete per event log-likelihood results on the held-out test for the eight benchmark datasets we consider, averaged over 5 random seeds. In Table 7(a) we show the full log-likelihood. We then decompose this log-likelihood into the log-likelihood of the event time in Table 7(b), and the time-conditional log-likelihood of the mark type in Table 7(c). OOM indicates out of memory; standard deviation in parentheses. We highlight the best-performing model in bold and underline the second-best. We also report the average rank of models across datasets as a summary metric (lower is better). DLHP is consistently the best or second best-performing model across all datasets.
(a) Full log-likelihood results (equal to the summation of Table 7(b) and Table 7(c)). Extended version of Table 1.
Model Per Event Log-Likelihood, $\mathcal{L}_{\text{Total}}$ (nats) ($\uparrow$) Avg. Ranking ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP -2.136 (0.003) -7.098 (0.217) 0.346 (0.002) 1.003 (0.004) -2.480 (0.019) -1.780 (0.005) -0.472 (0.026) -8.081 (0.025) 6.3
NHP 0.129 (0.012) -6.348 (0.000) 0.514 (0.004) 1.157 (0.004) -2.241 (0.002) -0.574 (0.011) 0.060 (0.017) -3.966 (0.058) 3.0
SAHP -2.074 (0.029) -6.708 (0.029) 0.298 (0.057) 1.168 (0.029) -2.341 (0.058) -1.646 (0.083) -0.677 (0.072) -6.804 (0.126) 5.4
THP -2.096 (0.002) -6.659 (0.007) 0.372 (0.002) 0.790 (0.002) -2.338 (0.014) -1.712 (0.011) -0.577 (0.011) -7.208 (0.096) 5.5
AttNHP 0.484 (0.077) -6.499 (0.028) 0.493 (0.009) 1.259 (0.022) -2.194 (0.016) -0.592 (0.051) -0.170 (0.077) OOM 3.1
IFTPP 0.496 (0.002) -10.344 (0.016) 0.453 (0.002) 1.318 (0.017) -2.233 (0.009) -0.492 (0.017) 0.317 (0.052) -6.596 (0.240) 2.9
DLHP (Ours) 0.781 (0.011) -6.365 (0.003) 0.522 (0.004) 1.304 (0.039) -2.163 (0.009) -0.557 (0.046) 1.243 (0.083) -2.512 (0.369) 1.4
(b) Per event log-likelihood of the event times (higher is better).
Model Per Event Next Event Time Log-Likelihood, $\mathcal{L}_{\text{Time}}$ (nats) ($\uparrow$) Avg. Ranking ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP 0.011 (0.001) -6.191 (0.083) 0.622 (0.002) 2.428 (0.004) -0.797 (0.005) 0.256 (0.007) -0.188 (0.016) -1.913 (0.025) 5.8
NHP 2.116 (0.009) -5.584 (0.001) 0.727 (0.003) 2.578 (0.006) -0.699 (0.002) 1.198 (0.006) 0.225 (0.016) -0.821 (0.045) 3.3
SAHP 0.115 (0.049) -5.872 (0.062) 0.645 (0.044) 2.604 (0.008) -0.703 (0.031) 0.489 (0.078) -0.244 (0.040) -1.801 (0.049) 4.6
THP -0.068 (0.002) -5.874 (0.007) 0.621 (0.002) 2.242 (0.002) -0.772 (0.006) 0.220 (0.010) -0.271 (0.004) -1.921 (0.027) 6.4
AttNHP 2.416 (0.092) -5.726 (0.027) 0.714 (0.010) 2.654 (0.007) -0.684 (0.005) 1.203 (0.015) 0.031 (0.055) OOM 3.3
IFTPP 2.483 (0.001) -9.500 (0.011) 0.735 (0.002) 2.708 (0.018) -0.662 (0.007) 1.277 (0.016) 0.555 (0.050) -2.640 (0.115) 2.9
DLHP (Ours) 2.652 (0.009) -5.598 (0.002) 0.733 (0.003) 2.719 (0.038) -0.641 (0.003) 1.257 (0.022) 1.389 (0.053) 0.382 (0.362) 1.4
(c) Per event log-likelihood of mark type conditioned on the arrival time (higher is better).
Model Per Event Next Mark Log-Likelihood, $\mathcal{L}_{\text{Mark}}$ (nats) ($\uparrow$) Avg. Ranking ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP -2.147 (0.003) -0.908 (0.141) -0.276 (0.000) -1.425 (0.002) -1.683 (0.015) -2.035 (0.004) -0.284 (0.014) -6.168 (0.025) 5.9
NHP -1.987 (0.003) -0.764 (0.000) -0.213 (0.002) -1.421 (0.004) -1.542 (0.001) -1.772 (0.006) -0.165 (0.002) -3.144 (0.016) 2.4
SAHP -2.189 (0.030) -0.836 (0.036) -0.346 (0.024) -1.436 (0.027) -1.638 (0.032) -2.136 (0.070) -0.433 (0.031) -5.003 (0.132) 6.1
THP -2.028 (0.002) -0.785 (0.001) -0.249 (0.001) -1.451 (0.000) -1.566 (0.008) -1.932 (0.006) -0.306 (0.009) -5.287 (0.107) 5.0
AttNHP -1.933 (0.024) -0.773 (0.003) -0.221 (0.002) -1.395 (0.016) -1.510 (0.013) -1.795 (0.037) -0.201 (0.025) OOM 2.4
IFTPP -1.988 (0.001) -0.844 (0.007) -0.282 (0.001) -1.391 (0.005) -1.571 (0.003) -1.769 (0.004) -0.239 (0.002) -3.956 (0.192) 3.8
DLHP (Ours) -1.871 (0.002) -0.767 (0.000) -0.211 (0.002) -1.415 (0.005) -1.521 (0.008) -1.814 (0.025) -0.145 (0.040) -2.893 (0.009) 1.9
Figure 5: Visualization of $\mathcal{L}_{\text{Total}}$ decomposed into $\mathcal{L}_{\text{Time}}$ and $\mathcal{L}_{\text{Mark}}$ for all models and all datasets relative to RMTPP, normalized by number of events, as discussed in Section 5.2. The improvement of DLHP is mainly driven by better modeling of $\mathcal{L}_{\text{Time}}$.
Table 8: Next event prediction accuracy (reported as a percentage, higher is better) conditioned on the true event time. We report top 1 accuracy for all datasets except for top 10 accuracy for EHRSHOT, due to the pre-processing procedure described in Section C.3. We bold the best result per dataset, and underline the runner-up.
Model Next Mark Accuracy (%) ($\uparrow$) Avg. Ranking ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT (Top 10)
RMTPP 30.89 53.76 91.32 60.78 46.03 52.42 92.32 36.79 6.1
NHP 39.32 61.43 92.97 61.48 47.21 56.55 94.32 76.41 1.8
SAHP 31.45 54.97 92.23 60.45 46.54 51.82 87.42 52.24 5.5
THP 34.63 60.20 91.46 59.99 46.79 53.38 91.09 51.85 5.3
AttNHP 38.88 60.90 92.74 61.32 48.38 56.39 93.65 OOM 2.6
IFTPP 35.90 54.92 91.76 60.93 45.66 56.44 93.32 66.38 4.4
DLHP (Ours) 40.76 61.36 93.27 61.19 47.86 56.28 96.55 79.20 1.9
Table 9: Comparison of RMSE of the next-event time prediction on the benchmarks we consider in Section 5.2 in the same format as Table 8. Lower RMSE values indicate better performance.
Model Next event time RMSE ($\downarrow$) Avg. Ranking ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP 0.361 16152 0.284 0.127 1.054 15.864 0.754 3445 4.8
NHP 0.342 15322 0.283 0.127 1.028 15.841 0.739 3430 3.1
SAHP 0.345 16018 0.285 0.126 1.034 15.830 0.805 3418 3.8
THP 0.335 15848 0.286 0.126 1.029 15.878 0.781 3504 4.3
AttNHP 1.214 16220 1.273 0.130 1.311 15.889 0.852 OOM 7.0
IFTPP 0.332 2568 0.280 0.126 0.975 15.550 0.713 1899 1.0
DLHP (Ours, sampling) 0.395 15223 0.282 0.127 1.021 15.844 0.764 3421 3.4
DLHP (Ours, linear probe) 0.337 14642 0.283 0.128 1.038 15.710 0.810 3350 N/A
D.2 Full Results for Synthetic Poisson Experiments
We present the full results in Fig. 6 for all models regarding the synthetic experiments discussed in Section 5.1. All models are trained until convergence using a set of 5,000 generated sequences, where we use 20 Monte Carlo points per event to estimate the integral of log-likelihood during training to accommodate the sparsity of events. We used small models so they do not overfit; model architecture and parameter counts are reported in Table 10. We plot the background intensity conditioned on empty sequences using 1,000 equidistant grid points between the start and end points. Our model is the only one that perfectly recovers the underlying ground truth intensity, while also using the fewest parameters.
Figure 6: Results for all baseline models for the synthetic Poisson experiment introduced in Section 5.1. The estimated intensity (blue lines) conditioned on an empty sequence are plotted against the ground truth (dotted black lines).
Table 10: Model architectures and parameter counts for synthetic Poisson experiments.
Model Architecture # Parameters
RMTPP $h=16$ 627
NHP $h=8$ 1010
SAHP $h=16,l=2,\mathrm{heads}=4$ 1738
THP $h=16,l=2,\mathrm{heads}=4$ 1684
AttNHP $h=8,t=2,l=2,\mathrm{heads}=2$ 1178
IFTPP $h=16$ 1899
DLHP (Ours) $h=4,p=4,l=2$ 178
D.3 Additional Synthetic Results on Multivariate Hawkes Processes
We evaluate our model and baseline models against the true model on a randomly initiated parametric Hawkes process with three possible marks. Following the notation in Section 2.1, we draw all parameters from the following distributions: $\bm{\nu}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.1,0.5]$, $\bm{\alpha}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.5,0.8]$, and $\bm{\beta}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.4,1.2]$ for $i,j\in\{1,2,3\}$.
All models are trained until convergence using a set of 50,000 generated sequences, where we use 20 Monte Carlo points per event to estimate the integral of log-likelihood during training. Model architecture and parameter counts are reported in Table 11. We plot three example sequences drawn for an additional test set for each model in Figs. 7 and 8, using 1,000 equidistant grid points for any inter-event interval. Dotted lines refer to the intensities under the true underlying parametric model; solid lines are different model estimates from trained models.
As we see in inhomogeneous Poisson processes, our model can recover the ground truth intensities with the fewest parameters. Both neural Hawkes processes and our DLHP show almost perfect recovery of parametric Hawkes processes, especially before seeing any event happening, and at event times. It is also worth noting that our model is 7-9$\times$ quicker than NHP and AttNHP regarding wall-clock runtime on a single A5000 GPU. Our results on synthetic experiments validate the model’s ability to recover the ground truth intensities.
We further evaluate all models quantitatively using 1,000 test sequences generated from the same multivariate Hawkes process and evaluated both log-likelihood and RMSE for the immediate next event. We see our method competitive again on both metrics.
Table 11: Model architectures and parameter counts for parametric Hawkes processes experiments.
Model Architecture # Parameters
RMTPP $h=16$ 697
NHP $h=8$ 1046
SAHP $h=16,l=2,\mathrm{heads}=4$ 1902
THP $h=16,l=2,\mathrm{heads}=4$ 1756
AttNHP $h=8,t=2,l=2,\mathrm{heads}=2$ 1230
IFTPP $h=16$ 1965
DLHP (Ours) $h=8,p=4,l=2$ 358
Figure 7: Our proposed DLHP model trained with 50k training sequences drawn from a randomly instantiated multivariate Hawkes process. Three example test sequences are plotted for each model.
Figure 8: Baseline models trained with 50k training sequences drawn from a randomly instantiated multivariate Hawkes process. Three example test sequences are plotted for each model.
Table 12: Performance comparison of models on the synthetic Hawkes process experiment presented above. Higher per event log-likelihood indicates better performance, whereas lower root mean squared error (RMSE) indicates better performance.
Model Total Log-likelihood $\mathcal{L}_{\mathrm{Total}}$ ($\uparrow$) Next-Event Time RMSE ($\downarrow$)
RMTPP -0.550 0.570
NHP -0.530 0.565
SAHP -0.537 0.569
THP -0.543 0.570
AttNHP -0.533 0.655
IFTPP -0.534 0.540
DLHP (Ours) -0.527 0.566
D.4 Ablation for Different DLHP Variants
We perform an ablation study of different model variants that we proposed on all datasets and summarize the results in Table 13. We train EHRSHOT using 10% of its training data because larger dataset scale requires more training time (but use the original validation and test sets for model selection and reporting results). Forward and backward discretization are very close in performance, with backwards discretization having a slight edge. Models that are input-dependent achieve better performance on most datasets, although on certain datasets input dependence appears to harm performance. It is an interesting direction for future work to explore theoretically and empirically when each of these variants is best. We select backward discretization with input dependence for the results in the main paper.
Table 13: Ablation for different model variants log-likelihood (LL). ID stands for input-dependent, see Section 3.4. Backward and Forward respectively refer to using $\mathbf{u}_{t_{i-1}}$ and $\mathbf{u}_{t_{i}-}$ (i.e., the previous right limit or current left limit), see Section B.4.
Dataset Model variant LL Arrival time LL Mark LL conditioned on time
Amazon Forward 0.705 2.617 -1.912
Forward + ID 0.748 2.634 -1.886
Backward 0.740 2.640 -1.899
Backward + ID 0.765 2.638 -1.873
Retweet Forward -6.405 -5.625 -0.780
Forward + ID -6.370 -5.602 -0.767
Backward -6.398 -5.618 -0.780
Backward + ID -6.367 -5.600 -0.767
Taxi Forward 0.473 0.697 -0.224
Forward + ID 0.525 0.733 -0.208
Backward 0.477 0.705 -0.228
Backward + ID 0.528 0.738 -0.209
Taobao Forward 1.207 2.643 -1.435
Forward + ID 1.332 2.742 -1.410
Backward 1.215 2.648 -1.432
Backward + ID 1.332 2.742 -1.410
StackOverflow Forward -2.249 -0.676 -1.572
Forward + ID -2.174 -0.644 -1.530
Backward -2.225 -0.679 -1.547
Backward + ID -2.165 -0.636 -1.529
Last.fm Forward -0.463 1.309 -1.772
Forward + ID -0.477 1.302 -1.779
Backward -0.474 1.303 -1.777
Backward + ID -0.496 1.294 -1.790
MIMIC-II Forward 0.555 0.847 -0.292
Forward + ID 1.319 1.405 -0.086
Backward 0.322 0.601 -0.279
Backward + ID 1.231 1.345 -0.114
EHRSHOT (10%) Forward -3.885 0.105 -3.990
Forward + ID -3.848 -0.021 -3.827
Backward -4.571 -0.432 -4.139
Backward + ID -4.684 -0.641 -4.043
D.5 Model Calibration
To further probe the models, we evaluate the calibration metrics of MTPPs that are proposed in literature (Bosser & Taieb, 2023), which has a different focus than log-likelihood-based evaluation. On a high level, calibration describes how well the uncertainty in the model is reflected in the observed data. However, a model can achieve perfect calibration by predicting the marginal distribution, so better calibration does not necessarily transform into better predictive performance. We therefore present these metrics as a secondary metric (secondary to log-likelihood per Daley & Vere-Jones (2003)) for investigating the performance of different models. We provide summarized statistics for both probabilistic calibration error (PCE) for time calibration and expected calibration error (ECE) for mark calibration in Table 14, and visualize the calibration curves in Figs. 9 and 10. From our results, all MTPP models are well-calibrated on most of the datasets, especially on mark predictions.
Table 14: Calibration results for the models and datasets tests.
(a) Probabilistic calibration error (PCE) for time calibration in percentage.
Model Probabilistic Calibration Error (PCE) ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP 13.67 7.71 3.51 9.97 1.78 11.54 4.23 13.07
NHP 8.16 0.14 1.27 6.66 1.42 4.60 5.87 7.42
SAHP 10.79 6.23 1.75 3.11 1.38 10.44 2.00 20.02
THP 12.38 5.60 3.33 16.32 1.92 10.97 1.05 13.73
AttNHP 6.50 1.08 0.53 2.85 1.46 1.36 4.84 OOM
IFTPP 1.54 23.35 0.47 0.61 0.68 0.30 1.89 15.14
DLHP (Ours) 5.81 0.91 0.15 1.91 0.82 0.73 10.35 12.16
(b) Expected calibration error (ECE) for mark calibration in percentage.
Model Expected Calibration Error (ECE) ($\downarrow$)
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP 6.42 2.06 2.24 1.52 2.09 2.49 2.87 8.22
NHP 8.02 0.41 0.77 5.28 1.15 2.73 1.92 4.21
SAHP 8.93 8.66 6.96 2.87 1.48 7.79 5.58 4.80
THP 2.14 1.27 1.65 6.51 0.94 4.24 2.42 9.95
AttNHP 2.35 0.33 0.57 2.56 1.31 0.51 1.97 OOM
IFTPP 0.60 2.93 0.87 1.48 1.41 0.59 1.66 1.92
DLHP (Ours) 1.30 0.45 0.52 1.48 1.74 0.31 1.31 2.47
Figure 9: Reliability diagram for predicted inter-arrival time for each model on all datasets. Diagonal dashed lines refer to perfect calibration.
Figure 10: Reliability diagram for mark prediction of all models and all datasets. The $x$-axis specifies the confidence of model estimates grouped into 20 bins, and the $y$-axis of the bar plot is the model accuracy within that bin. The diagonal lines represent perfect calibration. The solid curves depict the distribution of confidences, and do not share the $y$-axis. The grey dashed lines indicate the overall prediction accuracy of the model for the next event conditioned on true event time.
Finally, in Figs. 11 and 12 we plot the log-likelihood of time and mark respectively, versus their corresponding calibration results, to provide an overall view of the performances of different models. Our DLHP model consistently achieves higher log-likelihood while maintaining good calibration on both time and mark components on most datasets.
Figure 11: Log-likelihood of time vs. PCE for all models grouped by datasets. Higher log-likelihood and lower PCE are better (i.e., top left corner).
Figure 12: Log-likelihood of mark vs. ECE for all models grouped by datasets. Higher log-likelihood and lower ECE are better (i.e., top left corner).

### v2

@@
-Deep Linear Hawkes Processes
+Deep Continuous-Time State-Space Models for Marked Event Sequences
 Abstract
-Marked temporal point processes (MTPPs) are used to model sequences of different types of events with irregular arrival times, with broad applications ranging from healthcare and social networks to finance.
-We address shortcomings in existing point process models by drawing connections between modern deep state-space models (SSMs) and linear Hawkes processes (LHPs), culminating in an MTPP that we call the deep linear Hawkes process (DLHP).
-The DLHP modifies the linear differential equations in deep SSMs to be stochastic jump differential equations, akin to LHPs.
-After discretizing, the resulting recurrence can be implemented efficiently using a parallel scan.
-This brings parallelism and linear scaling to MTPP models.
-This contrasts with attention-based MTPPs, which scale quadratically, and RNN-based MTPPs, which do not parallelize across the sequence length.
-We show empirically that DLHPs match or outperform existing models across a broad range of metrics on eight real-world datasets.
-Our proposed DLHP model is the first instance of the unique architectural capabilities of SSMs being leveraged to construct a new class of MTPP models.
+Marked temporal point processes (MTPPs) model sequences of events occurring at irregular time intervals, with wide-ranging applications in fields such as healthcare, finance and social networks.
+We propose the state-space point process (S2P2) model, a novel and performant model that leverages techniques derived for modern deep state-space models (SSMs) to overcome limitations of existing MTPP models, while simultaneously imbuing strong inductive biases for continuous-time event sequences that other discrete sequence models (i.e., RNNs, transformers) do not capture.
+Inspired by the classical linear Hawkes processes, we propose an architecture that interleaves stochastic jump differential equations with nonlinearities to create a highly expressive intensity-based MTPP model, without the need for restrictive parametric assumptions for the intensity.
+Our approach enables efficient training and inference with a parallel scan, bringing linear complexity and sublinear scaling while retaining expressivity to MTPPs.
+Empirically, S2P2 achieves state-of-the-art predictive likelihoods across eight real-world datasets, delivering an average improvement of 33% over the best existing approaches.
 1 Introduction
-Figure 1: Intensity estimates from trained models when conditioned on an empty sequence $\mathcal{H}_{t}=\emptyset$ for NHP (Mei & Eisner, 2017) and DLHP, our method.
-Shown in dotted lines are the ground truth, inhomogeneous Poisson process intensity.
-Our DLHP is able to accurately capture the background intensity.
-See Section 5.1 for more details.
-Marked temporal point processes (MTPPs) are used to model irregular sequences of events in continuous-time, where each event has an associated type, often referred to as a mark.
-MTPPs model the joint distribution of marked event sequences.
-They have been successfully applied to modeling purchasing patterns in e-commerce (Türkmen et al., 2019; Vassøy et al., 2019; Yang et al., 2018), patient-specific medical events (Hua et al., 2022), disease propagation (Gajardo & Müller, 2023), and many other domains (Williams et al., 2020; Sharma et al., 2018; Wang et al., 2024).
-An MTPP is fully characterized by a marked intensity process which specifies the expected instantaneous rate of occurrence of events of each mark conditioned on the event history.
-State-of-the-art methods use neural networks to compute hidden states that summarize the event history, which are then used to compute marked intensities across future values of time.
-However, many models are limited by inexpressive temporal dynamics, lack of support for long-range dependencies, and serial computation (Du et al., 2016; Mei & Eisner, 2017).
-Recent advances in transformer-based MTPPs have improved performance and gained parallelism, but scale quadratically with sequence lengths (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022).
-Recently, deep state-space models (often abbreviated as SSMs) have emerged as a challenger to transformer-based models for discrete sequence modeling (Gu et al., 2022b; Smith et al., 2022; Gu & Dao, 2023).
-SSMs interleave a stack of linear state-space recurrences with position-wise non-linearities (Gu et al., 2021).
-This architecture has been found to be not only highly performant on a wide range of tasks (Goel et al., 2022; Deng et al., 2024), but retains linear scaling, can be parallelized across the length of a sequence, and can gracefully handle irregularly-spaced observations.
-Inspired by this, we revisit a foundational point process model, the linear Hawkes process (LHP Hawkes, 1971), and draw connections between LHPs and deep SSMs.
-We combine the parameterization and parallelization strategy of SSMs with the functional form of LHPs to create what we call the deep linear Hawkes process (DLHP).
-More formally, the DLHP is a fully-recurrent neural MTPP parameterized by a stack of stochastic jump differential equations on the complex plane (serving as the recurrence) interleaved with position-wise non-linearities (to improve expressivity).
-This design yields an MTPP with two main advantages over existing neural MTPPs: (i) parallelism across the length of the sequence through the use of parallel scans, and (ii) highly flexible intensity functions.
-This is achieved not only through the expressivity of the SSM-style architecture, but also by tying the output intensity at time $t$ to the model’s continuously-evolved hidden state $\mathbf{x}_{t}$ (extending ideas from Mei & Eisner (2017) and Yang et al.
-(2022), see Figs.
-1 and 2), and by going beyond the classical LHP form with input-dependent recurrent dynamics (akin to Mamba (Gu & Dao, 2023)).
-The contributions of this paper are as follows: We introduce a new family of marked point process models, deep linear Hawkes processes—the first MTPP model that fully leverages the architectural features of deep SSMs.
-We demonstrate that DLHPs match or exceed the performance of existing models across eight real-world datasets, with an average per-event likelihood improvement of 38% across datasets, over the individually best-performing existing method per dataset.
-We also verify that DLHP scales more effectively to longer sequences, a crucial capability for a wide range of applications.
-We will release our models, datasets, and pipelines as part of the existing EasyTPP library (Xue et al., 2023).
-We conclude by discussing the relative advantages and disadvantages of the DLHP over existing methods and opportunities for extending this work.
-Figure 2: Three different schematics of the deep linear Hawkes process (DLHP) and latent linear Hawkes (LLH) layer we propose.
-With increasing granularity: Left (a): On a high level, the DLHP can simply be viewed as a deep stack of neural network layers that transform an event sequence into an intensity function.
-Middle (b): On a more granular level, individual LLH layers can be viewed as discrete-time recurrences (see Eq.
-16), directly defining an intensity evaluated at select times: $t^{*}$ using $\mathbf{x}_{t^{*}}$, right limits $t_{i}+$ using $\mathbf{x}_{t_{i}}$, and left values $t_{i}$ using $\mathbf{x}_{t_{i}-}$.
-Right (c): The same recurrences can be viewed as a set of non-linearly coupled stochastic jump differential equations in continuous-time.
-Events are embedded and impart impulses to the differential equation.
-Impulses for layer $L$ are omitted for visual clarity.
-Intermediate intensities are evaluated using $\vec{0}$ for $\mathbf{u}^{(1)}$ and no impulse.
+Marked temporal point processes (MTPPs) are used to model irregular sequences of events in continuous time, where each event has an associated type, often called a mark.
+MTPPs model the joint distribution of these sequences of event times and marks.
+They have been successfully applied to modeling purchasing patterns in e-commerce (Türkmen et al., 2019; Vassøy et al., 2019; Yang et al., 2018), patient-specific medical events (Hua et al., 2022), disease propagation (Gajardo and Müller, 2023), and event modeling and prediction across multiple other domains (Williams et al., 2020; Sharma et al., 2018; Wang et al., 2024).
+An MTPP can be fully characterized by a marked intensity process, specifying the instantaneous rate of occurrence of each mark conditioned on history.
+State-of-the-art neural methods compute hidden states to summarize the event history, which are then used to compute marked intensities at any point in time.
+However, many models are limited by inexpressive temporal dynamics, lack of support for long-range dependencies, and serial computation (Du et al., 2016; Mei and Eisner, 2017).
+Recent advances in transformer-based MTPPs improved performance and gained parallelism, but scale quadratically in sequence length (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022), preventing them from being used in practice for long sequences, such as modeling clinical events over a patient’s entire medical history spanning many years.
+Rank (Lower ↓ is better, x = best, x = second-best)
+Likelihood Next Event Prediction Calibration Composite
+Model Mark Time Mark Acc. Time RMSE Mark Time Rank
+RMTPP (Du et al., 2016) 6.8 6.1 6.9 5.0 5.8 6.1 6.1
+SAHP (Zhang et al., 2020) 6.9 4.9 7.1 4.3 6.8 5.4 5.5
+THP (Zuo et al., 2020) 5.5 7.0 5.6 4.3 5.0 5.4 5.5
+IFTPP (Shchur et al., 2020a) 4.1 3.1 5.0 4.0 1.8 2.6 3.6
+MHP (Gao et al., 2024) 5.4 6.4 4.3 5.5 4.8 6.0 5.3
+NHP (Mei and Eisner, 2017) 2.4 3.3 1.8 2.4 4.9 3.6 2.9
+AttNHP Yang et al. (2022) 2.4 3.3 2.9 6.6 3.4 3.7 3.7
+S2P2 (Ours) 1.9 1.4 1.9 2.3 3.0 2.8 2.1
+Figure 1: (a) A schematic of our proposed state-space point process (S2P2), a deep stack of novel latent linear Hawkes (LLH) layers interleaved with nonlinear and normalization layers creating an expressive MTPP architecture.
+(b) Summary table of results we present in Section˜6.
+We summarize ranks for six key metrics, ranking the average held-out test set performance across eight real-world datasets for five randomly seeded models; as well as a holistic composite rank, defined as the average of the ranks.
+Our S2P2 model outperforms all baselines by almost an entire rank, strongly indicating state-of-the-art performance and robustness across metrics and datasets.
+Recently, deep state-space models (often abbreviated as SSMs) have emerged as a challenger to transformer-based models for discrete sequence modeling (Gu et al., 2022a; Smith et al., 2022; Gu and Dao, 2023).
+SSMs interleave a stack of linear state-space recurrences with position-wise nonlinearities (Gu et al., 2021).
+This architecture not only achieves superior performance on a wide range of tasks (Goel et al., 2022; Deng et al., 2024), but retains linear scaling, can be parallelized across the length of a sequence, and can gracefully handle irregularly spaced observations.
+Despite being defined in continuous-time, applying SSMs to MTPPs is not straightforward.
+Tying the conditional intensity to the model’s state is the obvious choice, but SSMs expect a continuous-valued signal to integrate as input, whereas event sequences are discontinuous by nature.
+Taking inspiration from classical parametric linear Hawkes processes (LHPs) (Hawkes, 1971), we introduce a stochastic jump differential equation on the complex plane to serve as the SSM recurrence.
+We refer to this as a latent linear Hawkes (LLH) layer.
+We derive closed-form update rules for any time that allow the diagonalized continuous-time SSM system to use efficient parallel scans.
+We then introduce our state-space point process (S2P2) model, a stack of LLH layers interleaved with position-wise nonlinear functions (Fig.˜1(a)), inspired by deep SSMs.
+This design yields an MTPP that has both a highly flexible conditional intensity function and access to efficient, parallelizable computation.
+This paper is organized as follows: In Section˜2 we present the necessary preliminaries on MTPPs, LHPs, and deep SSMs.
+In Section˜3 we introduce our state-space point process architecture.
+We first build the connection between LHPs and deep SSMs, and then derive how to use this connection to make a highly expressive, parsimonious and parallelizable MTPP.
+In Section˜5 we then verify this expressivity on small-scale, targeted synthetic explorations comparing to key baselines.
+In Section˜6 we empirically evaluate our model at scale on a range of metrics across eight real-world datasets, finding that S2P2 matches or exceeds the average predictive performance of baselines, achieving either best- or second-best average performance on all six metrics.
+These results are summarized in Table 1(b).
+In Sections˜7 and 8 we conclude by discussing the relative advantages and limitations of S2P2 and promising future opportunities.
+Due to its robustness, performance, and efficiency, our S2P2 model is a powerful model out-of-the-box for a wide range of MTPP applications.
 2 Preliminaries
 2.1 Marked Temporal Point Processes
-Let $t_{1},t_{2},\dots\in\mathbb{R}_{\geq 0}$ be a strictly increasing sequence of positive random variables, each representing the time of occurrence for an event of interest.
-[^1] Please refer to Tables 2 and 3 in Appendix A for a list of notation and acronym definitions, respectively.
-For each $t_{i}$, let $k_{i}\in\mathcal{M}$ be a random variable representing accompanying side-information, commonly referred to as an event’s mark, with $\mathcal{M}$ being the mark-space.
-In this paper, we focus on discrete and finite mark spaces, i.e., :=$\mathcal{M}:=\{1,\dots,K\}$; however, in general $\mathcal{M}$ can be continuous or even a mixture of continuous and discrete.
-Together $t_{i}$ and $k_{i}$ fully define a given event.
-The joint distribution over a sequence of continuous event times and mark types is described as a marked temporal point process.
-We use $\mathcal{H}_{t}$ to represent the sequence, or history, of events up to some time $t$: $\mathcal{H}_{t}:=\{(t_{i},k_{i})\;|\;t_{i}\leq t\text{ for }i\in\mathbb{N}\}$, with $\mathcal{H}_{t-}$ defined similarly except that it does not include events that occur at time $t$.
-One way of characterizing an MTPP is through a marked intensity process, which describes the instantaneous expected rate of occurrence for events of specific marks.
-Let $\mathbf{N}_{t}:=[N_{t}^{1},\dots,N_{t}^{K}]^{\top}\in\mathbb{Z}_{\geq 0}^{K}$ be the marked counting process which represents the number of occurrences of events of each type of mark in the time span $[0,t]$.
-The marked intensity process $\bm{\lambda}_{t}:=[\lambda_{t}^{1},\dots,\lambda_{t}^{K}]^{\top}\in\mathbb{R}_{\geq 0}^{K}$ characterizes an MTPP by describing how the counting process changes via:
-$\lambda_{t}^{k}\mathrm{d}t:=$ $\;\mathbb{E}\left[\text{event of type }k\text{ occurs in }[t,t+\mathrm{d}t]\;|\;\mathcal{H}_{t-}\right]=\;\mathbb{E}\left[N^{k}_{t+\mathrm{d}t}-N^{k}_{t}\;|\;\mathcal{H}_{t-}\right],$ (1)
+We define an event sequence, or history, as a sequence of time-mark pairs, $\mathcal{H}_{t}:=\{(t_{i},k_{i})\;|\;t_{i}\leq t\text{ for }i\in\mathbb{Z}^{+}\}$, where $t_{i}\in\mathbb{R}_{\geq 0}$, $\forall i:t_{i-1}<t_{i}$, and $k_{i}\in\mathcal{M}$.
+[^1] Please refer to Tables 3 and 4 in Appendix A for a list of notation and acronym definitions, respectively.
+In this paper, we focus on discrete and finite mark spaces, i.e., $\mathcal{M}:=\{1,\dots,K\}$; however, $\mathcal{M}$ can be more general, such as countable or continuous.
+We also define $\mathcal{H}_{t-}$ similarly to $\mathcal{H}_{t}$, except that it excludes events at exactly time $t$.
+One way of characterizing an MTPP is through a marked intensity process.
+The intensity $\bm{\lambda}_{t}:=[\lambda_{t}^{1},\dots,\lambda_{t}^{K}]^{\top}\in\mathbb{R}_{\geq 0}^{K}$ characterizes an MTPP by describing the rate of occurance of events:
+$\lambda_{t}^{k}\mathrm{d}t:=\;\mathbb{E}\left[\text{event of type }k\text{ occurs in }[t,t+\mathrm{d}t)\;|\;\mathcal{H}_{t-}\right]$ (1)
 with the total intensity $\lambda_{t}:=\sum_{k=1}^{K}\lambda^{k}_{t}$ being the rate that any event occurs.
-Note that the intensity conditions on the left limit of the history $\mathcal{H}_{t-}$ to ensure that the intensity is modeling future events.
-Parameterized forms of $\bm{\lambda}$ are often trained by optimizing the log-likelihood over observed data.
-The log-likelihood for a single sequence $\mathcal{H}_{T}$ is defined as (Daley & Vere-Jones, 2003, ch.
+This intensity also defines a marked counting process $\mathbf{N}_{t}:=[N_{t}^{1},\dots,N_{t}^{K}]^{\top}\in\mathbb{Z}_{\geq 0}^{K}$, representing the number of occurrences of events of each type of mark in the time span $[0,t]$.
+Parameterized forms of $\bm{\lambda}$ are often trained by optimizing the log-likelihood over observed data (e.g., (Mei and Eisner, 2017; Zuo et al., 2020)).
+The log-likelihood for a single sequence $\mathcal{H}_{T}$ is defined as (Daley and Vere-Jones, 2003, ch.
 7.3):
-$\mathcal{L}(\mathcal{H}_{T}):=\sum\nolimits_{i=1}^{|\mathcal{H}_{T}|}\log\lambda_{t_{i}}^{k_{i}}-\int_{0}^{T}\lambda_{s}\mathrm{d}s.$ (2)
-Linear Hawkes Processes  An (unmarked) Hawkes process (Hawkes, 1971), or more generally a self-exciting process, is a temporal point process where event occurrences increase the rate at which subsequent events occur soon thereafter.
-Of particular interest to us are linear Hawkes processes (LHPs), which are characterized by the following intensity process:
-$\lambda_{t}:=\nu+\int_{s=0}^{t-}h(t-s)\mathrm{d}N_{s}:=\nu+\sum\nolimits_{i=1}^{N_{t-}}h(t-t_{i}),$ (3)
-where $\nu>0$ is the background intensity, $h:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is the excitation function (or kernel), and $N_{t}$ is the associated counting process characterized by intensity $\lambda_{t}$.
-$N_{t-}$ is used as the upper limit in Eq.
-3 to ensure the intensity at time $t$ does not take into account an event that occurs at time $t$.
-Should $h$ correspond to the exponential decay kernel, $h(z)=\alpha\exp\left(-\beta z\right)$, then the LHP intensity process is Markov (Law & Viens, 2016) and admits the following stochastic differential form:
-$\mathrm{d}\lambda_{t}=\beta(\nu-\lambda_{t-})\mathrm{d}t+\alpha\mathrm{d}N_{t}\iff\lambda_{t}$ $=\nu+\int_{0}^{t-}\alpha\exp\left(-\beta(t-s)\right)\mathrm{d}N_{s}$ (4)
-$=\nu+\sum\nolimits_{i=1}^{N_{t-}}\alpha\exp\left(-\beta(t-t_{i})\right).$ (5)
-LHPs can be extended to the marked setting, with $K$ possible discrete marks, by replacing $\nu$ with a vector of $K$ background rates $\bm{\nu}:=[\nu_{1},\dots,\nu_{K}]^{\top}$, and the excitation effect $h(t-s)\mathrm{d}N_{s}$ with $\mathbf{h}(t-s)\mathrm{d}\mathbf{N}_{s}$.
-Here, $\mathbf{h}_{ij}$ of $\mathbf{h}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{K\times K}$ describes the excitation that events of type $i$ exerts on future events of type $j$.
-The counting process, $\mathrm{d}\mathbf{N}_{t}$, is then either a $K$-dimensional zero-vector if no event occurs at time $t$, or a one-hot vector indicating which mark is associated with the occurring event.
-Generalizing the exponential kernel to handle marks results in the following differential form:
-$\mathrm{d}\bm{\lambda}_{t}=-\bm{\beta}(\bm{\lambda}_{t-}-\bm{\nu})\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (6)
-where $\bm{\beta,\alpha}\in\mathbb{R}^{K\times K}_{\geq 0}$ are restricted to be non-negative to ensure non-negative marked intensities.
-2.2 State-Space Models
-Deep state-space models (SSMs) are a recent innovation in recurrent models that have found success in long-range sequence modeling tasks (Gu et al., 2022b) and language modeling tasks (Gu & Dao, 2023), while also having favorable computational properties.
-The backbone of deep SSMs is the linear state-space equations, which define a continuous-time dynamical system with input and output signals $\textbf{u}(t),\mathbf{y}(t)\in\mathbb{R}^{H}$, respectively, through linear differential equations:
-$\frac{\mathrm{d}}{\mathrm{d}t}\textbf{x}(t)$ $=\mathbf{A}\textbf{x}(t)+\mathbf{B}\mathbf{u}(t)$ (7)
-$\mathbf{y}(t)$ $=\mathbf{C}\mathbf{x}(t)+\mathbf{D}\mathbf{u}(t),$ (8)
-where $\textbf{x}(t)\in\mathbb{R}^{P}$ is the (hidden) state of the system, and $\mathbf{A}\in\mathbb{R}^{P\times P},\mathbf{B}\in\mathbb{R}^{P\times H},\mathbf{C}\in\mathbb{R}^{H\times P},$ and $\mathbf{D}\in\mathbb{R}^{H\times H}$ are the parameters that control the system’s dynamics.
-Deep SSMs then stack these recurrences interleaved with non-linear position-wise functions, $\sigma$.
-The function $\sigma$ can contain activation functions, residual connections and normalization layers, and transforms the output $\mathbf{y}$ of the previous recurrence into the input $\mathbf{u}$ of the next, i.e., $\mathbf{u}^{(l)}(t):=\sigma(\mathbf{y}^{(l-1)}(t))$ for layer $l$.
-This combination yields a sequence model where each recurrence is conditionally linear in time given the input, but is ultimately non-linear in depth due to the function $\sigma$.
-To evaluate the SSM, we first discretize the continuous-time system at the desired times, and then evaluate as though it were a conventional discrete-time RNN architecture.
-Crucially, the linearity of the resulting discrete-time recurrence allows it to be evaluated using parallel scans (Blelloch, 1990; Smith et al., 2022; Gu & Dao, 2023), leading to linear work scaling (i.e., number of operations), and, importantly, sublinear scaling of the computation time with respect to sequence length given sufficient parallel compute.
-Note this contrasts with conventional sequential RNNs (e.g., LSTMs), which process sequences serially; and attention-based methods, which can be parallelized over a sequence, but have quadratic work scaling with respect to sequence length.
-This allows SSMs to fully and efficiently leverage modern massively parallel hardware while also a being highly expressive and performant model class.
-Importantly for our purposes, evaluating a linear recurrence with a parallel scan natively admits evaluations with varying observation intervals.
-We will leverage this to parsimoniously handle the variable inter-event times observed in MTPP settings.
-3 Deep Linear Hawkes Processes
-In this section, we introduce our deep linear Hawkes process (DLHP), a neural MTPP that draws a novel connection between LHPs and deep SSMs.
-Stochastic jump differential equations, akin to the LHP intensity, form the basis of the conditionally linear recurrent layer, which we refer to as a latent linear Hawkes (LLH) layer.
-The LLH layer can be viewed as a modified SSM recurrence, while still admitting parallel computation.
-Taking further inspiration from deep SSMs, the DLHP is then made up of a stack of LLH layers, each interleaved with non-linear, position-wise transformations to increase the overall expressivity of the model (see Fig.
-2).
-In this section, we formalize this approach and outline implementation details.
+$\mathcal{L}(\mathcal{H}_{T}):=\sum\nolimits_{i=1}^{N_{T}}\log\lambda_{t_{i}}^{k_{i}}-\int_{0}^{T}\lambda_{s}\mathrm{d}s.$ (2)
+Linear Hawkes Processes  A foundational MTPP is the linear Hawkes process (LHP).
+The LHP is a self-exciting process, where event occurrence increases the rate of occurrence of other events, with the influence decaying according to a kernel.
+The intensity function is the summation of influences, and if the kernel is the exponential function, it has the following integral and differential forms:
+$\bm{\lambda}_{t}=\bm{\nu}+\sum\nolimits_{i=1}^{N_{t-}}\exp\left(-\bm{\beta}(t-t_{i})\right)\bm{\alpha}\iff\mathrm{d}\bm{\lambda}_{t}=-\bm{\beta}(\bm{\lambda}_{t-}-\bm{\nu})\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (3)
+where, to ensure non-negative marked intensities, $\bm{\nu}\in\mathbb{R}^{K}_{\geq 0}$ and $\bm{\alpha},\exp{(-\bm{\beta})}\in\mathbb{R}^{K\times K}_{\geq 0}$ with $\exp$ being the matrix exponential.
+This form (which is the most common) of the LHP is incredibly limited, and hence is used primarily for its interpretability, as opposed to outright predictive performance.
+2.2 Deep State-Space Models
+Deep state-space models (SSMs) are a class of recurrent models that have excelled in long-range sequence and language modeling tasks, all while having favorable computational properties (Gu et al., 2022a; Smith et al., 2022; Gu and Dao, 2023).
+The backbone of deep SSMs are linear state-space models, which define a continuous-time dynamical system with inputs and outputs $\textbf{u}(t),\mathbf{y}(t)\in\mathbb{R}^{H}$ through linear differential equations:
+$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{x}(t)$ $=\mathbf{A}\mathbf{x}(t)+\mathbf{B}\mathbf{u}(t),\quad\mathbf{A}\in\mathbb{R}^{P\times P},\mathbf{B}\in\mathbb{R}^{P\times H}$ (4)
+$\mathbf{y}(t)$ $=\mathbf{C}\mathbf{x}(t)+\mathbf{D}\mathbf{u}(t),\quad\mathbf{C}\in\mathbb{R}^{H\times P},\mathbf{D}\in\mathbb{R}^{H\times H}$ (5)
+where $\mathbf{x}(t)\in\mathbb{R}^{P}$ is the (hidden) state of the system, and $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D}$ define the system’s dynamics.
+Deep SSMs interleave these recurrences with nonlinear position-wise functions $\sigma$ as $\mathbf{u}^{(l)}(t):=\sigma(\mathbf{y}^{(l-1)}(t))$ (for layer $l$).
+This yields a sequence model where each recurrence is conditionally linear in time but is nonlinear overall due to $\sigma$.
+To evaluate an SSM, we first discretize the continuous-time system at appropriate times to yield a discrete sequence of closed-form state updates (Smith et al., 2022).
+The resulting discrete-time recurrence can be evaluated using parallel scans (Blelloch, 1990), with linear work scaling and, importantly, sublinear (theoretically logarithmic) scaling of the computation time with respect to sequence length given sufficient parallel compute.
+This parallel evaluation natively allows varying observation intervals or latent dynamics.
+3 State-Space Point Process
+We seek to define an MTPP model that is (a) highly expressive and (b) can access efficient and parallelizable compute methods.
+In this section, we formally introduce our state-space point process (S2P2) model, as outlined below: Section˜3.1 extends and generalizes the continuous-time form of the LHP, creating a layer we refer to as the latent linear Hawkes (LLH) layer.
+This generalization decouples the choice of the width of the layer from the mark space, allowing us to make arbitrarily wide layers, and is the basis for the novel connection between the LHP and deep SSMs.
+Section˜3.2 discusses how we make each LLH efficient and tractable to compute with parallel scans, and Section˜3.3 extends the model with time-varying dynamics to be more expressive.
+We conclude in Section˜3.4 by describing S2P2, a highly expressive MTPP model built from the composition of multiple LLH layers and nonlinearities that can be computed with a novel parallel inference scheme.
 3.1 Continuous-Time Latent Linear Hawkes Layer
-We first start by generalizing the intensity of the linear Hawkes process, Eq. 6:
-$\mathrm{d}\bm{\lambda}_{t}$ $=-\bm{\beta}(\bm{\lambda}_{t-}-\bm{\nu}_{t})\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t}=-\bm{\beta}\bm{\lambda}_{t-}\mathrm{d}t+\bm{\beta}\bm{\nu}_{t}\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (9)
-whereby the background intensity $\bm{\nu}_{t}$ is allowed to vary over time.
-If we compare this to the recurrence in Eq.
-7, we see that the intensity in the LHP, $\bm{\lambda}_{t}$ controlled by decay rates $\bm{\beta}$, is analogous the state in the linear SSM, $\mathbf{x}(t)$ controlled by state matrix $\mathbf{A}$.
-Additionally, the time-varying baseline intensity in the LHP, $\bm{\nu}_{t}$, is analogous to the SSM input signal, $\mathbf{u}(t)$.
-What is unique to the LHP is the (mark-specific) impulse $\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$.
-This impulse is important because it allows the model to instantaneously incorporate information from events as they occur, introducing discontinuities in the output signals of the otherwise continuously-integrated system, unlike conventional SSMs.
-With this in mind, we adapt Eq.
-9 such that it can replace the typical state-space recurrence in an SSM, Eq.
-7.
-To do so, we replace the non-negative $\bm{\beta}$ with an unrestricted state matrix $\mathbf{A}\in\mathbb{R}^{P\times P}$.
-Next, given an input signal $\mathbf{u}_{t}\in\mathbb{R}^{H}$ we project it to $P$ dimensions with an input matrix $\mathbf{B}\in\mathbb{R}^{P\times H}$ to replace $\bm{\nu}_{t}$.
-[^2] Here, we index time $t$ via subscripts (e.g., $\mathbf{u}_{t}$) rather than an argument $\left(\mathbf{u}(t)\right)$ to emphasize that these are stochastic (jump) processes rather than deterministic functions.
-What was originally the intensity $\bm{\lambda}_{t}$ is now relabeled to be the state of the layer $\mathbf{x}_{t}$.
-Finally, we allow the impulses to be low-rank by having a shared set of mark embeddings $\bm{\alpha}\in\mathbb{R}^{R\times K}$ with rank $R$ that are brought into $P$ dimensions with a layer-specific embedding matrix $\mathbf{E}\in\mathbb{R}^{P\times R}$.
-For simplicity, we set $R=H$ for all our experiments.
-The equation for the output signal $\mathbf{y}_{t}$ is left unchanged from Eq.
-8, where $\mathbf{C}\in\mathbb{R}^{H\times P}$ and $\mathbf{D}\in\mathbb{R}^{H\times H}$.
-All of this results in the set of equations that makes up what we call the latent linear Hawkes layer:
-$\mathrm{d}\mathbf{x}_{t}$ $=-\mathbf{Ax}_{t-}\mathrm{d}t+\mathbf{ABu}_{t-}\mathrm{d}t+\mathbf{E}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (10)
-$\mathbf{y}_{t}$ $=\mathbf{Cx}_{t}+\mathbf{Du}_{t},$ (11)
-where the initial state $\mathbf{x}_{0}\in\mathbb{R}^{P}$ is learned. Realizations of this layer can be seen in Fig. 2.
-3.2 Continuous-Time Deep Linear Hawkes Process Architecture
-Inspired by deep SSMs, our MTPP is formed by stacking LLH layers, chaining the output signal $\mathbf{y}$ of one layer to the input $\mathbf{u}$ of another with non-linear transforms in between.
-The final layer’s output is then transformed into the intensity $\bm{\lambda}$.
-An illustration of the DLHP architecture is shown in Fig.
-2.
-Let $L$ be the number of desired LLH layers that comprise a DLHP with input and output signals $\mathbf{u}^{(l)}$ and $\mathbf{y}^{(l)}$ respectively for layers $l=1,\dots,L$.
-For the very first layer, the only input available to condition on are the event occurrences themselves.
-As such, we set $\mathbf{u}_{t}^{(1)}=\mathbf{0}$ for all $t\geq 0$.
-In general, a layer’s output $\mathbf{y}^{(l)}:=\text{LLH}^{(l)}(\mathbf{u}^{(l)},\mathcal{H})$ is passed into a non-linear activation function $f$ (we use $f(z):=\text{GELU}(z)$ (Hendrycks & Gimpel, 2016)), summed with the residual stream $\mathbf{u}^{(l)}$, and normalized with LayerNorm (Ba, 2016) to compute the next layer’s input.
-More formally,
-$\mathbf{u}^{(l+1)}_{t}:=\text{LayerNorm}^{(l)}(f(\mathbf{y}^{(l)}_{t})+\mathbf{u}^{(l)}_{t})$ (12)
-for $t\geq 0$ and $l=1,\dots,L$.
-We use the same strategies for initialization as S5 (Smith et al., 2022), based off the performant HiPPO initialization scheme (Gu et al., 2020).
-Due to the transformations, unlike the original LHP, we cannot guarantee the output of the final layer is positive.
-Therefore, similar to Mei & Eisner (2017), we apply an affine projection followed by a rectifying transformation to enforce non-negative intensity:
-$\bm{\lambda}_{t}:=\mathbf{s}\odot\text{softplus}((\mathbf{W}\mathbf{u}_{t-}^{(L+1)}+\mathbf{b})\odot\mathbf{s}^{-1})$ (13)
-for $t\geq 0$ and where $\mathbf{W}\in\mathbb{R}^{K\times H}$, $\mathbf{b},\log(\mathbf{s})\in\mathbb{R}^{K}$, and $\odot$ is an element-wise product.
-Eq.
-13 implements the “Proj.
-& Softplus” layer in Fig.
-2.
-The intensity at time $t$ always uses the left-limit of $\mathbf{u}^{(L+1)}$, which in turn uses the left-limit of $\mathbf{y}^{(l)}$ and $\mathbf{u}^{(l)}$ for all $l$ to ensure that it has no information of any events that may or may not have occurred at time $t$ is used.
+To develop a recurrent layer for an MTPP model, we start by reviewing the equations for an LHP intensity, Eq.˜3, and an SSM state, Eq.˜4.
+If we allow $\bm{\nu}$ in the LHP to vary over time, we obtain:
 […]
-On the Relationship With the Linear Hawkes Process Before discussing how to compute the DLHP and its variations, we briefly reflect on the relationship between the DLHP and LHP.
-The derivation presented above shows the steps to modify an LHP to be a deep SSM.
-The connection, parameterization, and equivalence we explore does not materially affect the implementation; this was intended to concretely define how our model differs from a classical model, and to retain the intuition from the simpler LHP (even if the direct interpretability of the LHP is somewhat lost).
-Alternatively, one could have simply attempted to convert a deep SSM into an MTPP.
-Likewise, the steps taken and the result would be similar; however, the relationship to other models would be markedly less clear.
-We hope our exposition makes it clear how the DLHP is a natural extension of a known and well-used model rather than an arbitrary modification to a recent architecture.
-3.3 Discretizing & Diagonalizing the LLH Layer
-Unlike the LHP intensity, the recurrence in the LLH layer does not permit an analytic solution.
-As such, we must discretize the continuous-time process to compute values of the layer at specified time points.
+[…]
+When presented together the parallels between them become apparent: the LHP intensity, $\color[rgb]{1,0,0}\bm{\lambda}_{t}$} controlled by decay rates $\color[rgb]{1,0.49,0}\bm{\beta}$}, is analogous to the state in the linear SSM, $\color[rgb]{1,0,0}\mathbf{x}(t)$} controlled by state matrix $\color[rgb]{1,0.49,0}\mathbf{A}$}.
+The time-varying background LHP intensity, $\color[rgb]{0.55,0.14,1}\bm{\nu}_{t}$}, is analogous to the SSM input signal after being projected into the state-space, $\color[rgb]{0.55,0.14,1}\mathbf{Bu}(t)$}.
+Compared to SSMs which allow for recurrence over a latent space $\mathbb{R}^{P}$, Eq.˜6 is limited in expressivity due to restricted dynamics $\color[rgb]{1,0.49,0}\bm{\beta}$} and fixed dimensionality tied to the number of possible marks $K$.
+Not present in Eq.˜7 is an impulse $\color[rgb]{0.55,0,0}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$}, which is crucial for allowing the recurrence to condition on abrupt event occurrences over time.
+Figure 2: Schematic of our state-space point process.
+We depict the internals of a single LLH layer of the model in their continuous time form (left) and as discrete computations (right).
+Black arrows can be concurrently computed in logarithmic time, and gray arrows in constant time.
+With this in mind, we combine Eqs.˜6 and 7 to resolve their respective deficiencies.
+This results in the following set of stochastic jump differential equations that we call the latent linear Hawkes (LLH) layer, which will serve as a continuous-time recurrence layer in our proposed model:
+LLH: $\mathrm{d}{\color[rgb]{1,0,0}\mathbf{x}_{t}}$ $=-{\color[rgb]{1,0.49,0}\mathbf{A}}{\color[rgb]{1,0,0}\mathbf{x}_{t-}}\mathrm{d}t+{\color[rgb]{1,0.49,0}\mathbf{A}}{\color[rgb]{0.55,0.14,1}\mathbf{Bu}_{t-}}\mathrm{d}t+{\color[rgb]{0.55,0,0}\mathbf{E}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}}$ $\quad\in\mathbb{R}^{P}$ (8)
+$\mathbf{y}_{t}$ $=\mathbf{C}{\color[rgb]{1,0,0}\mathbf{x}_{t}}+\mathbf{D}{\color[rgb]{0.55,0.14,1}\mathbf{u}_{t}}$ $\quad\in\mathbb{R}^{H}$ (9)
+where ${\color[rgb]{0.55,0,0}\bm{\alpha}}\in\mathbb{R}^{R\times K}$, ${\color[rgb]{0.55,0,0}\mathbf{E}}\in\mathbb{R}^{P\times R}$, and the standard SSM parameters share the same dimensionality as in Eqs.˜4 and 5.
+Note that the resulting dynamics within the LLH layer are much more expressive than the LHP that it was inspired by, due to both operating in a separate set of latent dimensions and by having a general dynamics matrix ${\color[rgb]{1,0.49,0}\mathbf{A}}\in\mathbb{R}^{P\times P}$, compared to the more restrictive $\exp(-{\color[rgb]{1,0.49,0}\bm{\beta}})\in\mathbb{R}^{K\times K}_{\geq 0}$.
+For parameter efficiency, we allow the new impulse term $\color[rgb]{0.55,0,0}\mathbf{E}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$} to be computed as a product of model-wide, low-rank mark embeddings $\color[rgb]{0.55,0,0}\bm{\alpha}$} of rank $R$ and a layer-specific projection into state-space $\color[rgb]{0.55,0,0}\mathbf{E}$}.
+For simplicity, we set $R:=H$ in practice.
+Realizations of this layer are shown in Fig.˜2.
+This LLH will serve as the main form of recurrence in a larger continuous-time model that we define in Section˜3.4.
+In deriving this layer as an extension from LHPs, we retain a strong inductive bias towards event sequences.
+Furthermore, by drawing connections between this layer and traditional state-space equations, we allow the downstream model to benefit from many of the innovations designed for deep SSMs in recent years, such as parallel computation.
+3.2 Diagonalizing, Discretizing & Computing the LLH Recurrence
+Unlike the LHP intensity, there is no analytical solution to the continuous-time LLH recurrence due to the continuously-integrated $\mathbf{u}(t)$ signal.
+We must therefore approximate the system at specific timepoints.
 If we approximate the input signal by treating it as constant over an update interval, also known as a zero-order hold (ZOH) assumption (Iserles, 2009), then we can achieve a closed-form exact update to the recurrence relation.
-However, unfortunately, this involves a computationally-expensive matrix exponential in the update rule.
-To circumvent this, we first diagonalize the system and then impose the zero-order hold restriction on it.
-Doing so converts the matrix exponential into an element-wise exponential operation.
-This is done for all LLH layers that compose the DLHP.
-Note that this is same general approach taken by Smith et al.
-(2022) for deep SSMs.
-Diagonalization Let $-\mathbf{A}$ be diagonalizable with a factorization of $\mathbf{V}\bm{\Lambda}\mathbf{V}^{-1}$, where $\mathbf{V},\bm{\Lambda}\in\mathbb{C}^{P\times P}$ and $\bm{\Lambda}$ is a diagonal matrix of eigenvalues.
+To avoid a computationally expensive matrix exponential in the update rule, we adopt the same general approach as Smith et al.
+(2022) for deep SSMs by first diagonalizing the system and then imposing the zero-order hold restriction on it.
+This converts the matrix exponential into an element-wise exponential operation for each LLH layer in S2P2.
+Diagonalization:
+Let $\mathbf{A}$ be diagonalizable with a factorization of $\mathbf{V}\bm{\Lambda}\mathbf{V}^{-1}$, where $\mathbf{V},\bm{\Lambda}\in\mathbb{C}^{P\times P}$ and $\bm{\Lambda}$ is a diagonal matrix of eigenvalues.
 An equivalent, diagonalized LLH is then
@@
 where $\tilde{\mathbf{x}}_{t}=\mathbf{V}^{-1}\mathbf{x}_{t}$, $\tilde{\mathbf{B}}=-\mathbf{V}^{-1}\mathbf{B}$, $\tilde{\mathbf{E}}=\mathbf{V}^{-1}\mathbf{E}$, and $\tilde{\mathbf{C}}=\mathbf{CV}$.
-Note that in practice we directly parameterize $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$, and $\tilde{\mathbf{E}}$ to avoid having to learn and invert $\mathbf{V}$.
-The eigenvalues $\bm{\Lambda}$ are also directly parameterized and constrained with negative real-components for stability (Davis, 2013).
-While the dynamics are diagonalized, we note this does not mean that we are independently modeling the intensities of different mark types.
-This can be seen two ways: First, the diagonalized dynamics are equivalent to the original dynamics (see Eq.
-10, given the system can be diagonalized on the complex plane).
-Alternatively, the marks interact through the dense input and output matrices, the position-wise non-linearity, the mark embeddings, and the final intensity rectification layer.
-Discretization We then employ a ZOH discretization to create a closed-form update from the diagonalized continuous-time system.
-The ZOH assumption holds the input $\mathbf{u}$ constant over the integration period.
-This results in the following update rule that transitions from $\mathbf{x}_{t}$ to $\mathbf{x}_{t^{\prime}}$, where, by construction, no events occur in $(t,t^{\prime})$:
-$\tilde{\mathbf{x}}_{t^{\prime}}$ :=$:=\begin{cases}\bar{\bm{\Lambda}}\tilde{\mathbf{x}}_{t}+(\bar{\bm{\Lambda}}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t^{\prime}-}&\text{ if no event at }t^{\prime}\\ \bar{\bm{\Lambda}}\tilde{\mathbf{x}}_{t}+(\bar{\bm{\Lambda}}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t^{\prime}-}+\tilde{\mathbf{E}}\bm{\alpha}_{k}&\text{ if event of type }k\text{ at }t^{\prime}\end{cases}$ (16)
-where $\bar{\bm{\Lambda}}:=\exp(\bm{\Lambda}(t^{\prime}-t))$ (derivation in Section B.2). Please refer to Fig. 2 for an illustration.
-Note that the ZOH is an exact update when $\mathbf{u}$ is constant over the window $[t,t^{\prime})$.
-While we choose the constant value to set $\mathbf{u}$ to be as $\mathbf{u}_{t^{\prime}-}$, it is worth noting that technically any value $\mathbf{u}_{s}$ for $s\in[t,t^{\prime})$ is valid.
-We explore this design decision and the impact it has on performance in more detail in Sections B.4 and D.4.
+We can then directly parameterize $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$, and $\tilde{\mathbf{E}}$ to avoid learning and inverting $\mathbf{V}$.
+We use the same initialization strategies as S5 (Smith et al., 2022) for $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$, and $\tilde{\mathbf{E}}$, based off the HiPPO initialization scheme (Gu et al., 2020).
+The eigenvalues $\bm{\Lambda}$ are also directly parameterized and constrained with negative real-components to enforce stability (Davis, 2013).
+While the dynamics are diagonalized, we note this does not mean that we are modeling the intensities of different mark types as independent.
+Marks interact through both the position-wise nonlinearities and the learned projections $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$ and $\tilde{\mathbf{E}}$ (since the diagonalized dynamics are equivalent to the original dynamics in Eq.˜8, given that the system can be diagonalized on the complex plane).
+Discretization:
+We then use a ZOH discretization to create a closed-form update from the diagonalized continuous-time system; please refer to Fig.˜2 for an illustration and Section˜B.2 for the full derivation.
+This results in the following update rule in the diagonalized eigenbasis, that transitions from $\mathbf{x}_{t}$ to $\mathbf{x}_{t^{\prime}}$ without a matrix exponential, where, by construction, no events occur in $(t,t^{\prime})$:
+[…]
+where $\bar{\bm{\Lambda}}:=\exp(\bm{\Lambda}(t^{\prime}-t))$.
+ZOH is an exact update when $\mathbf{u}$ is constant over the window $t,t^{\prime})$; as such, we can choose any value $\mathbf{u}_{s}$ for $s\in[t,t^{\prime})$ to hold constant as the input over the integration period.
+We opt to use $\mathbf{u}_{t^{\prime}-}$ so that the model can condition on the fact that no events have occurred between $t$ and $t^{\prime}$.
+This design decision and its impact on performance are explored in more detail in Sections˜B.4 and D.6.
 It is important that $\mathbf{u}_{t^{\prime}}$ is not used as the ZOH value to avoid data leakage.
+Computing LLH Recurrence:
+The final component is to derive how to use parallel scans to efficiently evaluate the closed-form updates for the modified LLH recurrence in parallel.
+Parallel scans admit efficient inference over linear recurrences of the form $\mathbf{z}_{i+1}=\mathbf{R}_{i}\mathbf{z}_{i}+\mathbf{b}_{i}$ (Blelloch, 1990).
+Although we have an impulse in the recurrence, Eq.˜8 is still intrinsically of this form, where $\mathbf{z}_{i}:=\mathbf{x}_{t_{i}}$, $\mathbf{R}_{i}:=\exp(\bm{\Lambda}_{i}(t_{i+1}-t_{i}))$, and $\mathbf{b}_{i}:=(\mathbf{R}_{i}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t_{i+1}-}+\tilde{\mathbf{E}}\bm{\alpha}_{k_{i+1}}$.
+As a result, we can leverage parallel scans to compute the sequence of right-limits $\mathbf{x}_{t_{1:N}}$ in parallel across the sequence length.
+The corresponding left-limits $\mathbf{x}_{t_{1:N}-}$, which will later be used to calculate event intensities, can then be directly and efficiently computed by subtracting the impulse, $\tilde{\mathbf{E}}\bm{\alpha}_{k_{1:N}}$, from $\mathbf{x}_{t_{1:N}}$.
 3.3 Input-Dependent Dynamics
-Inspired by recent developments in modern SSMs (e.g., Mamba (Gu & Dao, 2023)), we also consider allowing the dynamics of the system to vary depending on the input and history of previous events.
+Inspired by recent developments in modern SSMs (e.g., Mamba (Gu and Dao, 2023)), we also consider allowing the dynamics of the system to vary depending on the input and on the history of previous events.
 This can allow for more expressive intensities.
@@
 Alternatively, larger values will result in more quickly “forgetting” the influence of previous events for a given hidden state channel.
-This is formalized with the following recurrence relation:
-$\mathrm{d}\tilde{\mathbf{x}}_{t}$ $:=\bm{\Lambda}_{i}\tilde{\mathbf{x}}_{t-}\mathrm{d}t+\bm{\Lambda}_{i}\tilde{\mathbf{B}}\mathbf{u}_{t-}\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (17)
-for $t\in(t_{i},t_{i+1}]$ where $\bm{\Lambda}_{i}:=\text{diag}\left(\text{softplus}(\mathbf{W}^{\prime}\mathbf{u}_{t_{i}}+\mathbf{b}^{\prime})\right)\bm{\Lambda}$ with $\mathbf{W}^{\prime}\in\mathbb{R}^{P\times H}$ and $\mathbf{b}^{\prime}\in\mathbb{R}^{P}$.
-Note that this is still conditionally linear in time as even though $\bm{\Lambda}_{i}$ changes it is entirely input-dependent based on $\mathbf{u}$ and not dependent on previous values of $\mathbf{x}$.
-3.5 Computing LLH Recurrence
-Thus far, we have created the LLH layer by diagonalizing and discretizing a modified SSM.
-As discussed earlier, we would like to take advantage of the efficient parallel scans leveraged by many SSM-based models (Smith et al., 2022; Gu & Dao, 2023; Dao & Gu, 2024).
-Below we explain how we can still use the parallel scan, despite the modified recurrence.
-Parallel scans admit efficient inference over linear recurrences of the form $\mathbf{z}_{i+1}=\mathbf{A}_{i}\mathbf{z}_{i}+\mathbf{b}_{i}$ (Blelloch, 1990).
-Although we have added an impulse to the recurrence, this is still intrinsically of this form, where $\mathbf{z}_{i}:=\mathbf{x}_{t_{i}}$, $\mathbf{A}_{i}:=\exp(\bm{\Lambda}_{i}(t_{i+1}-t_{i}))$, and $\mathbf{b}_{i}:=(\mathbf{A}_{i}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t_{i+1}-}+\tilde{\mathbf{E}}\bm{\alpha}_{k_{i+1}}$.
-As a result, we can leverage efficient parallel scans to compute the sequence of right-limits $\mathbf{x}_{t_{1:N}}$ in parallel across the sequence length.
-The corresponding left-limits $\mathbf{x}_{t_{1:N}-}$ can then be efficiently computed after by subtracting off $\tilde{\mathbf{E}}\bm{\alpha}_{k_{1:N}}$ from $\mathbf{x}_{t_{1:N}}$.
-In Algorithms 1, 2 and 3 we compactly detail how to use a parallel scan to compute the sequence of right limits given events; how to evolve those right limits to compute left limits; and then how to subsequently compute the log-likelihood of the sequence.
-4 Related Works
-Neural MTPPs Marked temporal point processes (MTPPs) are generative models that jointly model the time and type of continuous-time sequential events, typically characterized by mark-specific intensity functions (Daley & Vere-Jones, 2003).
-Early approaches, such as self-exciting Hawkes processes (Hawkes, 1971; Liniger, 2009), used simple parametric forms for the intensity.
-More recently, neural architectures such as RNNs (Du et al., 2016; Mei & Eisner, 2017), CNNs (Zhuzhel et al., 2023), and transformers (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022) have been used to more flexibly model the conditional intensity.
-For intensity-free MTPPs, approaches include normalizing flows (Shchur et al., 2020a; Zagatti et al., 2024), neural processes (Bae et al., 2023), and diffusion models (Zeng et al., 2023; Zhang et al., 2024); however, the most common approach is to model intensities as it requires fewer modeling restrictions.
-Efficient MTPPs Due to their recurrent nature, RNN-based MTPP models incur $\mathcal{O}(N)$ complexity for sequences of length $N$ as events must be processed sequentially.
-Attention-based MTPP models can be applied in parallel across the sequence, but the computational work scales as $\mathcal{O}(N^{2})$.
+This is formalized with the following recurrence relation for $t\in(t_{i},t_{i+1}]$:
+$\mathrm{d}\tilde{\mathbf{x}}_{t}$ $:=\bm{\Lambda}_{i}\tilde{\mathbf{x}}_{t-}\mathrm{d}t+\bm{\Lambda}_{i}\tilde{\mathbf{B}}\mathbf{u}_{t-}\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (13)
+where $\bm{\Lambda}_{i}:=\text{diag}\left(\text{softplus}(\mathbf{W}^{\prime}\mathbf{u}_{t_{i}}+\mathbf{b}^{\prime})\right)\bm{\Lambda}$ with $\mathbf{W}^{\prime}\in\mathbb{R}^{P\times H}$ and $\mathbf{b}^{\prime}\in\mathbb{R}^{P}$.
+This is conditionally linear in time, as even though $\bm{\Lambda}_{i}$ changes, it is entirely input-dependent based on $\mathbf{u}$ and not dependent on previous values of $\mathbf{x}$, and hence we can still use parallel scans as discussed above.
+3.4 State-Space Point Process Architecture
+We are now well-positioned to present the state-space point process (S2P2), a flexible and parallel deep continuous-time model for marked intensities $\bm{\lambda}_{t}$.
+This model is shown in Figs.˜1(a) and 2.
+We have demonstrated how to define an efficient, parallelizable and scalable core MTPP layer in the LLH layer.
+While the diagonalized and discretized LLH layer is more expressive than the LHP, it remains fundamentally linear.
+To compensate, we take inspiration from deep SSMs and alternate $L$ LLH layers with position-wise nonlinearities.
+This creates a nonlinear model that can still use parallel computation over the sequence length, but is highly expressive compared to each linear recurrence.
+We have an input signal in two parts: (i) the continuously-integrated signal $\mathbf{u}_{t}$ and (ii) the discrete event impulses $\bm{\alpha}_{k}$.
+For the first layer, the only inputs available to condition on are the event impulses themselves, so we set $\mathbf{u}_{t}^{(1)}=\mathbf{0}$ for all $t\geq 0$.
+At deeper layers, we have a layer-specific impulse as well as the continuously integrated signal from the previous layer.
+In general, a layer’s output $\mathbf{y}^{(l)}:=\text{LLH}^{(l)}(\mathbf{u}^{(l)},\mathcal{H})$ is passed into a nonlinear activation function $\sigma$ (we use $\sigma(z):=\text{GELU}(z)$ (Hendrycks and Gimpel, 2016)), summed with the residual stream $\mathbf{u}^{(l)}$, and normalized with LayerNorm (Ba, 2016) for the next layer’s input.
+Formally, for $t\geq 0$ and $l=1,\dots,L$, then
+$\mathbf{u}^{(l+1)}_{t}:=\text{LayerNorm}^{(l)}(\sigma(\mathbf{y}^{(l)}_{t})+\mathbf{u}^{(l)}_{t})$. (14)
+Due to the unrestricted nature of the recurrences and nonlinearities (and unlike the original LHP), we enforce non-negative intensities by applying an affine projection followed by a rectifying transformation, similar to Mei and Eisner (2017).
+This is referred to as the “Proj.
+& Softplus” layer in Fig.˜1, and expressed as $\bm{\lambda}_{t}:=\mathbf{s}\odot\text{softplus}((\mathbf{W}\mathbf{u}_{t-}^{(L+1)}+\mathbf{b})\odot\mathbf{s}^{-1})$ for $t\geq 0$ and where $\odot$ is an element-wise product, $\mathbf{W}\in\mathbb{R}^{K\times H}$, and $\mathbf{b},\log(\mathbf{s})\in\mathbb{R}^{K}$.
+Then S2P2 can be trained by maximizing the log-likelihood of each sequence, Eq.˜2.
+Similar to other neural MTPPs, we use Monte Carlo estimation for the integral term $\int_{0}^{T}\lambda_{s}\mathrm{d}N_{s}$ (Mei and Eisner, 2017).
+Training the model requires computing intensities at both event times $t_{1:N}$ and at randomly sampled times $t\!\sim\!\mathcal{U}(0,T)$.
+To compute intermediate intensities at these sampled points, we take advantage of the continuous-time nature of S2P2 and partially evolve the latent state through the system dynamics.
+To do so, we first compute the right limits of the hidden states at event times, $\mathbf{x}_{t_{i}}^{(1:L)}$, as described above (we can do this conditioning in logarithmic depth).
+Then, Eq.˜12 is applied, with no impulse, as no events are occurring at these intermediate points, to find $\mathbf{x}_{t}^{(1:L)}$.
+From there, the evolved hidden states are used to evaluate the model across depth to compute the intermediate intensity $\bm{\lambda}_{t}$.
+This operation is both efficient and can be done in constant complexity because intermediate evaluations are conditionally independent given the right limits at events.
+Crucially, there is no separate parametric decoding head, unlike, for instance, Mamba Hawkes processes (Gao et al., 2024) or transformer Hawkes processes (Zuo et al., 2020).
+Instead, this tying of the intensity to the model’s continuously evolving hidden states, more like the neural Hawkes processes (Mei and Eisner, 2017), makes S2P2 a continuous-time model and contributes to its enhanced expressivity (see Section˜5 and Section˜7 for more discussions).
+In Algorithms˜1, 2 and 3, we explicitly detail how to use a parallel scan to compute the sequence of right limits at events, how to then evolve those to compute left limits, and then how to subsequently compute the log-likelihood of the sequence.
+4 Related Work
+Neural MTPPs: MTPPs are generative models that jointly model the time and type of continuous-time sequential events, typically characterized by mark-specific intensity functions (Daley and Vere-Jones, 2003).
+Early approaches used parametric intensity functions, such as self-exciting Hawkes processes (Hawkes, 1971; Liniger, 2009).
+More recently, neural models such as RNNs (Du et al., 2016; Mei and Eisner, 2017) and transformers (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022) were developed to enable flexible modeling of conditional intensities.
+Intensity-free MTPPs include normalizing flows (Shchur et al., 2020a; Zagatti et al., 2024), transformers (Draxler et al., 2025), neural processes (Bae et al., 2023), and diffusion models (Zeng et al., 2023; Zhang et al., 2024); however, modeling intensities is more common, requiring fewer modeling restrictions.
+Efficient MTPPs: Due to their recurrent nature, RNN-based MTPPs incur $\mathcal{O}(N)$ scaling for length-$N$ sequences as events are processed sequentially.
+Attention-based MTPP models can be applied in parallel across the sequence, but with $\mathcal{O}(N^{2})$ computational work.
 Türkmen et al.
-(2020) proposed modeling events as conditionally independent so long as they occurred within the same time bin of a specified size.
-This resulted in parallel computation within bins, but still scales overall as $\mathcal{O}(N)$.
+(2020) model events as conditionally independent if they occur within the same time bin of a specified size, resulting in parallel computation within bins, but still scaling as $\mathcal{O}(N)$ overall.
 Shchur et al.
-(2020b) proposed an intensity-free TPP which uses triangular maps and the time-change theorem (Daley & Vere-Jones, 2003).
+(2020b) proposed an intensity-free TPP using triangular maps and the time-change theorem (Daley and Vere-Jones, 2003).
 This was extended by Zagatti et al.
-(2024) to handle marks, but in doing so, lost many of the benefits of the original model and scales linearly in the mark dimension—which can rapidly become untenable with $O(NK)$ work.
-To the best of our knowledge, our proposed model is the first that efficiently scales with sequence length and mark space, as well as the first to fully leverage SSMs and parallel scans.
-SSMs for Sequential Modeling SSMs have found recent success as alternatives to RNNs, CNNs, and transformers, enjoying reduced training cost and comparable modelling power (Gu et al., 2022b).
-A range of variants have been developed (Gu et al., 2021; Gupta et al., 2022; Gu et al., 2022a; Smith et al., 2022), and have been applied in language modeling (Gu & Dao, 2023), speech (Goel et al., 2022), and vision (Wang et al., 2023; Zhu et al., 2024).
-The linear recurrence allows for parallelism, as well as accessible long contexts which would be prohibitive for transformers due to their quadratic scaling.
-However, SSMs have not previously been applied to MTPPs, in part due to the irregular inter-event times and the input being a stochastic counting process.
-Concurrent work by Gao et al.
-(2024) used Mamba (Gu & Dao, 2023), a recent deep SSM architecture, in an MTPP setting, which they call the Mamda Hawkes Process (MHP).
-The MHP uses a mamba SSM as the encoder in an encoder-decoder architecture, also leveraging the variable interval capabilities.
-Crucially, however, they use a separate parametric decoder for intermediate intensities (similar to, for instance, the THP).
-This misses the opportunity to “fully” leverage the SSM architecture, re-using the same variable interval evaluation to evaluate the the intensity.
-5 Experiments
-We now evaluate our deep linear Hawkes process model.
-Our core objectives in using SSMs for MTPP modeling were to define an architecture that is both (a) highly performant in its forecasting ability, and (b) able to leverage efficient parallel compute methods to accelerate inference.
-To this end, we first present a simple exploration of the ability of different models to represent parametric intensity functions.
-Then we present the main experiments in this paper, where we evaluate our model against a suite of common MTPP models on a range of datasets of different sizes.
-We conclude by testing the runtime of our model against a variety of baselines.
-We find that DLHP systematically outperforms baseline methods both in terms of log-likelihood on held-out test data and runtime across a range of sequence lengths.
-More results and details are included in Appendices C and D.
-Metrics
-[…]
-Models
-We compare our model (DLHP) with six of the most common MTPP models: two RNN-based models (RMTPP (Du et al., 2016), NHP (Mei & Eisner, 2017)), three transformer/attention-based models (THP (Zuo et al., 2020), SAHP (Zhang et al., 2020), AttNHP (Yang et al., 2022)), and one intensity-free model (IFTPP (Shchur et al., 2020a)).
-In all real-world experiments, extensive grid searches were conducted for hyperparameter tuning with configurations chosen based on validation log-likelihood, with the search spaces designed to roughly control parameter counts across models.
-Specifics for training, hyperparameters, and architectures are given in Appendix C.
-Table 1: Per event log-likelihood ($\uparrow$ is better) results on the held-out test set averaged over five seeds (standard deviations in parentheses); OOM indicates insufficient memory.
-We bold the best and underline the runner-up per dataset.
-We also report the mean rank of models across datasets as a summary metric ($\downarrow$ is better).
-DLHP is consistently the best or second best-performing model.
-Extended results and discussion are presented in Section D.1.
-Model Per Event Log-Likelihood, $\mathcal{L}_{\text{Total}}$ (nats)($\uparrow$) Avg. Ranking ($\downarrow$)
-Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
-RMTPP -2.136 (0.003) -7.098 (0.217) 0.346 (0.002) 1.003 (0.004) -2.480 (0.019) -1.780 (0.005) -0.472 (0.026) -8.081 (0.025) 6.3
+(2024) to handle marks but losing benefits of the original model and scaling linearly in the mark dimension—which can rapidly become untenable as $\mathcal{O}(NK)$ work.
+Our S2P2 scales as $\mathcal{O}(\log N)$ and efficiently in marks; more discussion in Section˜B.5.
+SSMs for Sequential Modeling: SSMs have found recent success as alternatives to RNNs, CNNs, and transformers, enjoying reduced training cost and comparable modelling power (Gu et al., 2022a).
+A range of variants have been developed (Gu et al., 2021; Gupta et al., 2022; Gu et al., 2022b; Smith et al., 2022) and applied in language modeling (Gu and Dao, 2023), speech (Goel et al., 2022), and vision (Wang et al., 2023; Zhu et al., 2024).
+The linear recurrence enables parallelism, as well as accessible long contexts that are prohibitive for transformers due to quadratic scaling.
+However, SSMs have not previously been used as continuous-time models for MTPPs, in part due to the intensity functions having different left and right limits, and the input being a stochastic counting process.
+SSMs for TPPs:Gao et al.
+(2024) propose using an SSM as a discrete sequence model, encoding an event sequence into a fixed set of static hidden states, then computing intensities with a separate parametric decoder (similar to transformer Hawkes processes (Zuo et al., 2020), but with the transformer replaced with Mamba (Gu and Dao, 2023)).
+This is fundamentally different to our continuous-time model, where instead of a separate parametric decoding head, we leverage the continuously evolving latent state at any time $t$ to compute predicted intensities at corresponding times.
+We empirically compare to their model MHP in Section˜6, finding that MHP (i) performs comparably to the THP, and (ii) is comprehensively outperformed by S2P2.
+5 Synthetic Experiments
+We first verify the ability of our S2P2 model to represent known intensity functions through a series of targeted experiments on synthetic data.
+Recent results from the deep SSM literature (Muca Cirone et al., 2024) show that sufficiently deep SSMs are able to approximate arbitrary continuous functions; their results apply to our S2P2 model, and hence we should be able to (verifiably) recover arbitrary and known intensity functions.
+To demonstrate this, we showcase recovering ground truth intensities in three different settings: the classical point processes of Hawkes and self-correcting processes; an inhomogeneous Poisson process with a discontinuous intensity function; and finally a marked process with long-range dependencies between marks.
+Our S2P2 recovers the correct intensity in all cases.
+We include full results for all baselines and reproducibility details for all the following examples and tasks in Appendix˜D, including intensity visualizations, where S2P2 works as intended, corrects recovers intensities, and matches or exceeds baseline methods performance.
+Figure 3: Intensity estimates from trained models when conditioned on an empty sequence $\mathcal{H}_{t}=\emptyset$ for NHP (Mei and Eisner, 2017) and our S2P2.
+Dotted lines show the ground truth intensity for an inhomogeneous Poisson process.
+S2P2 accurately captures the background intensity.
+Classical Point Processes: We first apply our model and baseline models to two classical point processes: a Hawkes process and a self-correcting process with intensity functions $\lambda_{t}=0.5+\sum_{i}0.5\exp(t_{i}-t)$ and $\lambda_{t}=\exp(t-N_{t})$, respectively.
+Examples of the recovered intensity functions are shown in Fig.˜4(a).
+We see that our S2P2 recovers the ground truth intensities nearly perfectly.
+We also explore applying our model and other baseline models to randomly generated synthetic multivariate Hawkes processes in Section˜D.3, finding that all methods perform comparably.
+Inhomogeneous Poisson Process: Our S2P2 does not have a fixed parametric form for the intensity decoder (cf.
+THP or MHP) or limited recurrent dynamics (cf.
+NHP).
+This flexibility should allow the S2P2 to capture intensities where other methods fail.
+This is showcased in Fig.˜3.
+Models are trained on sequences drawn from an inhomogeneous Poisson process with a square wave for an intensity function (except for $t>7$).
+We observe that baseline models fail in predictable ways due to their expressivity limitations (please refer to Section˜D.2 for full results), whereas our model successfully captures the true background intensity process almost perfectly.
+Long-Range Dependencies: Lastly, we study a task with known long-range dependencies, depicted in Fig.˜4(b).
+In this task, a “trigger” mark (green) is drawn from a homogeneous Poisson process ($\lambda$=0.1); each trigger is then followed by a “target” mark (orange) a predictably long time later.
+Here “long” refers to the fact that (a) there are often many trigger marks before a single target mark, and (b) the variance of the distribution over the trigger-target time is comparatively narrow compared to the mean.
+Gray marks are “distractors” drawn from a homogeneous Poison process ($\lambda$=1).
+We see that S2P2 successfully captures the long-range dependencies, whereas NHP struggles.
+In spite of the long delay between cause and effect within the data, S2P2 was able to recover 98% of the true data likelihood, whereas NHP only achieved 88%.
+This shows that the state-space point process successfully captures long-range dependencies and is flexible as an approximation family.
+(a) S2P2 successfully recovers parametric intensities from self-correcting and Hawkes processes.
+(b) Long-range dependencies between $\rightarrow||$ are accurately captured by S2P2, whereas NHP struggles to recover $\lambda_{t}$.
+Figure 4: Results for synthetic experiments in Section˜5.
+6 Real-World Experiments
+We now present the main experiments in this paper, evaluating on real-world datasets.
+[^2] Our model is fully integrated into the EasyTPP (Xue et al., 2023) library [link].
+Other code changes to reproduce our results can be found in our forked repository [link].
+Model checkpoints are available on request.
+Further results and extensive experimental details are included in Appendices˜C and D.
+Datasets:
+We compare models on eight different datasets, including five datasets available from EasyTPP (Xue et al., 2023) (Amazon, Retweet, Taxi, Taobao and StackOverflow).
+We also add two commonly used datasets in the literature (Last.fm and MIMIC-II), as well as a new medical events dataset derived from the publicly available EHRSHOT dataset (Wornow et al., 2023).
+EHRSHOT is an order of magnitude larger than any dataset in EasyTPP in both the number of marks and the maximum sequence length (see summary statistics in Table˜7), representing a challenging application.
+Metrics:
+We performed a comprehensive evaluation using three common families of metrics: (i) total per-event log-likelihood (Mei and Eisner, 2017; Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022; Gao et al., 2024) (also broken down into mark- and time-specific likelihoods); (ii) predictive accuracy summary metrics of next-mark classification and next event time RMSE (Du et al., 2016; Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022; Gao et al., 2024); and (iii) mark and event time calibration metrics (Zhang et al., 2020; Bosser and Taieb, 2023) measuring the reliability of the implied uncertainties (see Section˜D.7).
+In order to characterize overall performance across multiple diverse metrics, we also compute a summary composite rank in Table 1(b), defined as the average rank obtained for a given model across all dataset-metric combinations across five random seeds (i.e., is an average of 240 individual evaluations and five differently initialized trained models).
+Models:
+We use the EasyTPP library (Xue et al., 2023) for baseline models and S2P2.
+We perform extensive hyperparameter searches for each individual pair of baseline model and dataset, then select the configurations that maximizes validation log-likelihood per model/dataset.
+The range of grid search and other details are elaborated in Section˜C.1; results are reported on a fully held-out test set.
+All models were trained on a single 24GB NVIDIA A5000 GPU.
+Table 2: Results for S2P2 and baselines across key metrics.
+We show the mean on the held-out test set evaluated across five random training seeds (std.
+dev.
+in parentheses).
+OOM indicates insufficient memory.
+We also report the mean rank of models across datasets as a summary metric (for which lower $\downarrow$ is better).
+S2P2 is consistently the best or second best-performing model on average on each dataset and metric.
+Extended results and discussion are shown in Section˜D.1.
+(a) Per event total log-likelihood. Higher values indicate better performance.
+Model Per Event Log-Likelihood, $\mathcal{L}_{\text{Total}}$ (nats) (Higher $\uparrow$ is better, x = best, x = second-best) Average
+Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT Rank ($\downarrow$)
+RMTPP -2.136 (0.003) -7.098 (0.217) 0.346 (0.002) 1.003 (0.004) -2.480 (0.019) -1.780 (0.005) -0.472 (0.026) -8.081 (0.025) 7.1
+SAHP -2.074 (0.029) -6.708 (0.029) 0.298 (0.057) 1.168 (0.028) -2.341 (0.058) -1.646 (0.083) -0.677 (0.072) -6.804 (0.126) 5.8
+THP -2.096 (0.002) -6.659 (0.007) 0.372 (0.002) 0.790 (0.002) -2.338 (0.014) -1.712 (0.011) -0.577 (0.011) -7.208 (0.096) 6.1
+IFTPP 0.496 (0.002) -10.344 (0.016) 0.453 (0.002) 1.318 (0.017) -2.233 (0.009) -0.492 (0.017) 0.317 (0.052) -6.596 (0.240) 3.0
+MHP -2.091 (0.002) -6.564 (0.015) 0.370 (0.008) 0.636 (0.004) -2.346 (0.012) -1.676 (0.004) -0.351 (0.012) -7.206 (0.407) 5.9
 NHP 0.129 (0.012) -6.348 (0.000) 0.514 (0.004) 1.157 (0.004) -2.241 (0.002) -0.574 (0.011) 0.060 (0.017) -3.966 (0.058) 3.0
-SAHP -2.074 (0.029) -6.708 (0.029) 0.298 (0.057) 1.168 (0.028) -2.341 (0.058) -1.646 (0.083) -0.677 (0.072) -6.804 (0.126) 5.4
-THP -2.096 (0.002) -6.659 (0.007) 0.372 (0.002) 0.790 (0.002) -2.338 (0.014) -1.712 (0.011) -0.577 (0.011) -7.208 (0.096) 5.5
 AttNHP 0.484 (0.077) -6.499 (0.028) 0.493 (0.009) 1.259 (0.022) -2.194 (0.016) -0.592 (0.051) -0.170 (0.077) OOM 3.1
-IFTPP 0.496 (0.002) -10.344 (0.016) 0.453 (0.002) 1.318 (0.017) -2.233 (0.009) -0.492 (0.017) 0.317 (0.052) -6.596 (0.240) 2.9
-DLHP (Ours) 0.781 (0.011) -6.365 (0.003) 0.522 (0.004) 1.304 (0.039) -2.163 (0.009) -0.557 (0.046) 1.243 (0.083) -2.512 (0.369) 1.4
-Libraries and Compute Environment
-We implement our DLHP in the EasyTPP library (Xue et al., 2023) and use their implementations of the baseline models.
-We also use the five standard datasets that EasyTPP immediately supports (see Section C.2 for more details).
-We then further include three larger datasets to stress-test the MTPP models (see Section 5.2).
-Unless otherwise stated, all models were trained using a single NVIDIA A5000 GPU with 24GB of onboard memory.
-5.1 Synthetic Experiments
-We start by performing a simple investigation into the expressivity of the DLHP intensity function and the ability to capture background intensities.
-We train our model and baselines on 5,000 sequences over the time period $[0,7.5]$ drawn from an unmarked, inhomogeneous Poisson process with a square-wave intensity function, $\lambda_{t}:=\mathds{1}(t\in(1,2)\cup(3,4)\cup(5,6))$ (see Fig.
-1).
-We plot the estimated intensity functions conditioned on no events occurring, i.e., $\mathcal{H}_{t}:=\emptyset\;\forall t$.
-Intensity estimates are shown for NHP and DLHP specifically in Fig.
-1 (and for all models in Fig.
-6 in Section D.2).
-We can see that our model successfully captures the true, underlying background intensity process almost perfectly.
-This is largely attributed to the expressivity of the linear recurrences and non-linear depth of the model.
-Other models have various failure modes: struggling to capture the multi-modality (RMTPP, NHP, SAHP, and THP), not matching the square shape (previous four and IFTPP), or not being able to stop the pattern from repeating a fourth time (AttNHP).
-We also examine the performance of our DLHP on randomly instantiated parametric Hawkes processes, finding that DLHP is able to successfully recover the ground truth intensity.
-Furthermore, in this setting DLHP achieves the best held-out log-likelihood scores and competitive next-event time prediction (see Section D.3 for more details).
-These simple experiments confirm that the DLHP is sufficiently expressive to be able to represent more complicated intensity functions while other methods break down.
-5.2 Log-likelihood Results on Real-world Datasets
-We empirically investigate the performance of our proposed model against baseline methods by comparing the held-out log-likelihood per event.
-We evaluate our model on eight real-world datasets.
-Five of which are taken directly from EasyTPP (Xue et al., 2023).
-We also include two MTPP datasets that have been widely used throughout the literature: Last.fm, which includes data on users’ music listening patterns from Celma Herrada et al.
-(2009), and MIMIC-II, a subset of de-identified patient hospital visits processed from (Saeed et al., 2002).
-Finally, we introduce a third dataset from the recently released, publicly available electronic health record (EHR) dataset EHRSHOT (Wornow et al., 2023).
-To construct the dataset, we first establish the most used Current Procedural Terminology (CPT-4) codes that identify medical services and procedures as events.
-The processed dataset comprises sequences of CPT-4 codes issued to individual patients during their care.
-This dataset has a maximum sequence length 10$\times$ longer than the longest in the EasyTPP datasets (and 100$\times$ that of MIMIC-II), providing a challenging testbed (in terms of scale) beyond existing datasets.
-Data statistics and other details including pre-processing are provided in Table 6 and Section C.3.
-Figure 3: Per event log-likelihood on the held-out test data in Table 1 decomposed into time and mark components (i.e., $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{Time}}+\mathcal{L}_{\text{Mark}}$).
-Models are ordered by their average ranking.
-Model results are adjusted by subtracting the log-likelihood achieved by RMTPP for readability.
-From the held-out log-likelihood results shown in Table 1, DLHP consistently achieves the best or the second-best log-likelihood across all datasets.
-Compared to the best baseline model per-dataset, DLHP produces a (geometric) mean likelihood ratio of 1.38 (corresponding to 38% higher likelihood on true events).
-We decompose this improvements in Fig.
-3, where we see the improvements in log-likelihood are mainly driven by better modeling of time.
-Extended plots included in Section D.1.
-Given the clear improvement in temporal modeling, we posit that DLHPs are particularly well suited in applications that contain more complex patterns over time.
-All of these results for DLHP utilize input-dependent dynamics (see Section 3.4).
-This was found to reliably improve forecasting performance in ablation studies (see Section D.4).
-We also report and discuss additional metrics in the Appendix.
-We report next-mark classification accuracy and RMSE of next-event arrival time, finding that DLHP matches or outperforms all baselines.
-We also report model calibration with respect to next event time and mark prediction.
-Calibration aims to grade the predictive uncertainty of the model (Bosser & Taieb, 2023), which is not captured by other metrics such as mark classification accuracy and time RMSE.
-On the whole, our model (as well as the baselines) tend to produce well-calibrated time and mark predictions across the datasets.
-We also include full tables for the likelihood decomposition in Table 7.
-5.3 Speed Testing
-A key motivation for DLHP is to leverage the properties of SSMs to improve scaling and accelerate inference.
-To test this, we measure the wallclock time for a full forward pass and log-likelihood evaluation on random input sequences with lengths ranging from eight events to over half a million events.
-The architectures and mark spaces are the same as in the StackOverflow experiments (see Tables 5(a) and 6).
-Our EasyTPP PyTorch DLHP is written in pure PyTorch, and hence is not as optimized as other methods (compared to, for instance, IFTPP, which uses a GPU-optimized implementation of the GRU).
-We therefore include the runtimes of a standalone JAX DLHP implementation, which allows for similar optimization through JIT compilation.
-Figure 4: Median runtime, over 20 random seeds, for various models against increasing sequence lengths.
-We show runtimes for both conditioning on a sequence (Algorithm 1) and likelihood evaluation (Algorithm 3).
-We see that DLHP is faster across a wide range of sequence lengths.
-Variance of runtime is negligble.
-Crosses indicate where the THP runs out of memory or the IFTPP throws an error.
-The THP scales well before reverting to superlinear scaling, and then runs out of memory.
-The IFTPP is very fast at shorter runtimes, but quickly reverts to linear scaling, due to its simple but highly optimized implementation and inherently sequential operation.
-Both DLHP implementations scale linearly at long sequence lengths, but have near-constant runtime at shorter sequences.
-At shorter sequence lengths, the more optimized JAX implementation is faster than the unoptimized pure PyTorch implementation.
-While this indicates that there are additional opportunities to accelerate the PyTorch implementation further (e.g., exploiting kernel fusion or writing a lower-level Triton implementation), these results still confirm our core hypothesis: that our DLHP can exploit parallel scans to scale to long sequences more effectively than other methods.
+S2P2 (Ours) 0.781 (0.011) -6.365 (0.003) 0.522 (0.004) 1.304 (0.039) -2.163 (0.009) -0.557 (0.046) 0.919 (0.069) -2.512 (0.369) 1.4
+(b) RMSE of the next event time prediction. Lower RMSE values indicate better performance.
+Model Next Event Time RMSE (Lower $\downarrow$ is better, x = best, x = second-best) Average
+Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT Rank ($\downarrow$)
+RMTPP 0.338 (0.000) 16488 (70.5) 0.283 (0.001) 0.126 (0.000) 1.049 (0.003) 15.873 (0.000) 0.749 (0.010) 3425 (0.2) 5.0
+SAHP 0.335 (0.001) 16102 (62.4) 0.290 (0.008) 0.126 (0.000) 1.031 (0.011) 15.757 (0.007) 1.142 (0.198) 3374 (9.4) 4.3
+THP 0.332 (0.000) 16268 (18.7) 0.285 (0.001) 0.125 (0.000) 1.033 (0.005) 15.871 (0.000) 0.768 (0.005) 3414 (1.0) 4.3
+IFTPP 0.327 (0.000) 16625 (0.2) 0.362 (0.178) 0.125 (0.000) 1.340 (0.724) 16.508 (0.555) 0.767 (0.029) 3616 (17.6) 5.0
+MHP 0.329 (0.000) 16109 (36.9) 0.284 (0.003) 0.126 (0.000) 1.046 (0.030) 15.871 (0.000) 0.758 (0.065) 3418 (5.8) 4.0
+NHP 0.339 (0.000) 15911 (4.0) 0.282 (0.001) 0.126 (0.000) 1.019 (0.001) 15.733 (0.008) 0.726 (0.001) 3330 (30.9) 2.4
+AttNHP 2.656 (1.950) 16171 (284.2) 1.739 (0.422) 0.130 (0.000) 1.256 (0.030) 15.865 (0.017) 0.860 (0.022) OOM 6.6
+S2P2 (Ours) 0.327 (0.000) 15987 (13.7) 0.281 (0.000) 0.126 (0.000) 1.014 (0.001) 15.720 (0.000) 0.894 (0.054) 3368 (14.4) 2.3
+(c) Next mark classification accuracy (top-10 accuracy on EHRSHOT). Higher values indicate better performance.
+Model Next Mark Classification Accuracy (%) (Higher $\uparrow$ is better, x = best, x = second-best) Average
+Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT Rank ($\downarrow$)
+RMTPP 30.8 (0.1) 53.4 (0.6) 91.4 (0.1) 60.9 (0.1) 45.6 (0.3) 52.5 (0.1) 92.3 (0.3) 36.5 (0.2) 6.9
+SAHP 32.4 (1.0) 57.5 (2.2) 91.4 (0.7) 60.5 (0.2) 44.7 (2.0) 51.8 (0.7) 86.8 (0.9) 49.6 (3.1) 7.1
+THP 34.6 (0.1) 60.2 (0.1) 91.4 (0.0) 60.0 (0.0) 46.6 (0.2) 53.3 (0.1) 90.9 (0.2) 49.7 (1.7) 5.6
+IFTPP 35.9 (0.1) 50.4 (2.5) 91.8 (0.0) 61.0 (0.1) 45.6 (0.1) 56.4 (0.1) 93.4 (0.1) 62.7 (3.8) 4.3
+MHP 35.1 (0.1) 60.0 (0.2) 91.4 (0.3) 60.7 (0.2) 46.5 (0.1) 54.3 (0.1) 93.2 (0.3) 54.8 (4.5) 5.0
+NHP 39.4 (0.1) 61.4 (0.0) 92.9 (0.1) 61.5 (0.2) 47.1 (0.1) 56.5 (0.1) 94.3 (0.0) 76.5 (0.2) 1.8
+AttNHP 38.9 (0.9) 60.7 (0.2) 92.6 (0.1) 61.3 (0.2) 48.2 (0.2) 55.8 (0.6) 92.9 (0.6) OOM 2.9
+S2P2 (Ours) 40.7 (0.0) 61.3 (0.0) 93.1 (0.1) 61.1 (0.1) 47.5 (0.3) 55.8 (0.4) 96.0 (0.4) 79.5 (0.3) 1.9
+Main Results: We report full numerical results in Table˜2, and a summary of rankings in Table 1(b).
+Table˜2(a) shows that S2P2 consistently achieves the best or the second-best held-out log-likelihood results across all datasets, beating other methods by over a whole rank and with a (geometric) mean likelihood ratio of 1.33 (corresponding to a 33% higher likelihood of true events).
+This is calculated by computing the mean log-likelihood ratio across all datasets and then exponentiating.
+We further investigate the log-likelihood improvement of S2P2 by separating the time- and mark-specific log-likelihood, finding that S2P2 is mainly driven by better temporal modeling and while achieving gains in both time and mark modeling over existing methods (see Tables˜8 and 6 in Section˜D.1 for full results).
+Additionally, we report prediction summary metrics, next event time prediction and next mark classification, in Table˜2(b) and Table˜2(c).
+We again see that S2P2 performs well, matching or exceeding the best-performing baseline, and far outpacing other baseline methods.
+Additional Experiments: We defer several additional experiments and explorations to Appendices:
+  1. 1.
+    Calibration results: We also include full calibration results, both in raw numerical form and as reliability graphs (Bosser and Taieb, 2023).
+These show that S2P2 is well calibrated, placing highest amongst intensity-based methods, only being beaten by IFTPP (Shchur et al., 2020a).
+  2. 2.
+    Computational scaling: We define the theoretical time complexity of each baseline, where S2P2 is best with logarithmic complexity and linear work. We verify this scaling empirically.
+  3. 3.
+    Ablations: We also perform an ablation study on whether input dependent dynamics and the choice of which input to hold constant over the integration interval (introduced in Sections˜3.2 and 3.3) affect performance.
+We find input-dependent dynamics nearly always improve performance, and that performance is less sensitive to the specific form of the LLH’s input signal $\mathbf{u}_{t}$ in ZOH.
+7 Discussion
+Continuous-Time Hidden States:
+The S2P2 and NHP (Mei and Eisner, 2017) architectures are closely related, in that both model continuous-time latent states.
+For the NHP, the intensity function is parameterized by a hidden state $\mathbf{h}(t)$ evolved via a continuous-time LSTM variant.
+Similarly, S2P2 continuously evolves a set of continuous-time latent states $\mathbf{h}^{1:L}(t)$, one for each layer $\{1,\ldots,L\}$.
+Then the hidden state $\mathbf{h}^{L}(t)$ of the top layer is decoded into intensities at any $t$.
+In contrast, models such as THP, MHP, IFTPP, only have latent states $\mathbf{h}(t_{j})$ at discrete event times $t_{j\in\{1:N\}}$.
+Consequently, some parametric shapes need to be defined to “interpolate” the hidden state values between events, thus sacrificing models’ flexibility due to these parametric functions.
+Our results are grouped according to whether or not models have continuous-time hidden states.
+We empirically verify that continuously evolving hidden states enable more expressive latent dynamics, where the expressiveness in explaining the data is measured through log-likelihood; see Section˜D.1.
+S2P2 further improves scalability by incorporating SSMs that are naturally in continuous-time and by leveraging efficient parallelized scans.
+S2P2 also introduces a stronger inductive bias than most MTPP models (similar to NHP), helping to achieve superior next event predictions.
+Limitation and Future Directions:
+S2P2 forgoes the interpretability of the parameters and latent states of linear Hawkes processes (LHPs), meaning it is not appropriate when the interpretation of the underlying system is crucial.
+This tradeoff is common with neural MTPP models (e.g., NHP), because the limited expressivity of LHP may lead to poor predictive performance, as well as misleading interpretations due to underfitting.
+Therefore, an exciting opportunity for future research is exploiting the connection to recover LHP-like interpretability while retaining the enhanced predictive power of S2P2.
+For example, any recently developed techniques for interpreting deep SSMs can be extended to S2P2, such as Ali et al.
+(2025), while fully taking advantage of the continuously varying attention from continuous-time latent dynamics of S2P2 will provide richer insights compared to discrete-time attention maps (e.g., large language models).
+Beyond this, further directions include leveraging theories and best practices from deep SSMs, such as the enhanced parameterizations presented by Merrill et al.
+(2024).
+S2P2 can also naturally be extended and accommodate non-categorical marks by tailoring its architecture to the structure of the mark space, such as marks having continuous values or containing richer information.
+A concrete example is spatio-temporal point processes that have applications in seismic and weather forecasting.
+Other valuable extensions include serving S2P2 as a pre-trained backbone for downstream applications, such as EHRSHOT clinical classification tasks.
+Additionally, censoring or adversarial conditions are commonly seen in practice (Boyd et al., 2023; Chakraborty et al., 2025).
+Extending the evaluation of S2P2 and MTPP models in general remains an open direction for real-world MTPP deployments.
 8 Conclusion
-We present the deep linear Hawkes process (DLHP)—a novel combination of ideas from LHPs and SSMs.
-Our DLHP leverages the unique properties of deep SSM architectures to achieve a flexible and performant model, without additional and restrictive intensity decoding heads.
-We then demonstrated that our DLHP outperforms existing methods across a range of standard and new benchmark tasks over various metrics, such as log-likelihood and runtime across sequence lengths.
-One limitation of our method is the increased complexity of the implementation (as we require a parallel scan), compared to, for instance, the NHP (which only requires a basic for loop).
-Following from this, a second limitation is that we have lost most of the interpretability of the latent dynamics and parameters enjoyed by the LHP.
-While there is some limited interpretability from the mark-specific impulses due to each LLH layer incrementing the residual stream of the model, this is largely correlative and imprecise.
-Future research directions therefore include improving on and strengthening these aspects, as well as developing additional theory around the use of deep SSMs in this novel MTPP setting, and developing heuristics and best-practices for configuring hyperparameters in various modeling settings.
-Specifically, exploring the relative benefits of the forward and backward discretization is a unique research direction arising from the DLHP.
-However, we believe the robustness, predictive performance, computational efficiency, and extensibility of DLHPs make them a very competitive model out-of-the-box for a wide range of applications.
-Supplementary Materials:
-Deep Linear Hawkes Processes
+In this paper, we present the state-space point process (S2P2) model—a novel fusion of concepts from LHPs and SSMs.
+S2P2 uses deep stacks of stochastic jump differential equations to create an expressive and parsimonious MTPP without additional and restrictive intensity decoding heads, while simultaneously being able to leverage parameterizations and techniques borrowed from deep SSM architectures.
+We demonstrated that S2P2 outperforms existing methods across a range of standard and new benchmark tasks, and over a range of predictive metrics and efficiency evaluations.
+As an intensity-based model, S2P2 requires numerical integration for predicting the expected next event times (as with any neural intensity-based MTPP model) and is not able to model point masses in time (due to the MTPP assumption of no concurrent events).
+All other baselines suffer from at least one key additional deficiency (e.g., NHP has linear scaling, THP has quadratic work, IFTPP has a parametric decoder, etc.), all of which are resolved by S2P2, while also achieving better performance.
+We provide a PyTorch implementation via EasyTPP for out-of-the-box usage.
+We believe the robustness, state-of-the-art predictive performance across a range of metrics, computational efficiency, and extensibility of S2P2 make it a very competitive model for a wide range of MTPP applications.
+Acknowledgements
+We thank the reviewers for their invaluable feedback on improving the paper.
+This work was supported by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at the University of California, Irvine, by the National Institutes of Health under awards R01-LM013344 and R01CA297869, and by GE HealthCare.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: All of the claims are further expanded upon and justified in either the methodology section or the two experiment sections.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We discuss the limitations in the last two sections in our concluding thoughts.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: There are no theoretical results in our work.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Extensive reproducibility details are provided in Appendices˜C and D.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Please refer to the footnote on page 8.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: All details can be found in Appendix˜C.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: All numerical results for experiments are presented with standard deviations describing the spread over five different random seeds.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: As mentioned in the real-world experiments section, all models were trained on a single 24GB NVIDIA A5000 GPU.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We fully comply with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49.
+Justification: We do not believe this paper has any societal impacts, positive or negative.
+Neural MTPPs have existed for almost a decade, and our work is the most recent iteration in this well-established domain.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: All of the models were trained on publicly available datasets without direct risk for misuse.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: Credit is given in detail in Section˜C.3.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64.
+Justification: Data preprocessing scripts and instructions for acquisition for the new dataset (EHRSHOT) we propose are included in Section˜C.3 and our forked repository (see the footnote on page 8).
+Direct access to the data is not provided due to separate usage agreements and required training from the original source.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: Crowdsourcing nor human subjects were used in our work.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Human subjects were not used in our work.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: Our method nor paper used LLMs.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Supplementary Materials: Deep Continuous-Time State-Space Models for Marked Event Sequences
 Appendix A Acronyms and Notation
@@
 $t+$ $\mathbb{R}_{\geq 0}$ Subscript plus indicates right-limit
-$k$ =$\mathcal{M}=\left\{1,\ldots,K\right\}$ Event mark
+$k$ $\mathcal{M}=\left\{1,\ldots,K\right\}$ Event mark
 $\mathcal{H}$ $\mathcal{M}^{N}\times\mathbb{R}_{\geq 0}^{N}$ Event history for $N$ events
@@
 $\lambda_{t}$ $\mathbb{R}_{\geq 0}$ Ground/total intensity (sum of mark-specific intensities)
-[…]
+$\mathcal{L}(\cdot)$ $\mathbb{R}$ Log-likelihood of the argument under the model
 $\nu^{\mathrm{k}}$ $\mathbb{R}_{\geq 0}$ Background intensity for the $k^{\mathrm{th}}$ mark
@@
 $\bm{\beta}$ $\mathbb{R}^{K,K}_{\geq 0}$ (For LHP) Dynamics matrix of intensity evolution
-[…]
-[…]
+$R$ $\mathbb{Z}^{+}$ Mark embedding rank
+$P$ $\mathbb{Z}^{+}$ LLH/SSM hidden dimension
 $\mathbf{x}_{t}$ $\mathbb{R}^{P}$ LLH/SSM hidden state at time $t$
 $\mathbf{x}_{0}$ $\mathbb{R}^{P}$ Learned LLH/SSM initial hidden state
-[…]
+$H$ $\mathbb{N}$ LLH/SSM output dimension
 $\mathbf{y}_{t}$ $\mathbb{R}^{H}$ LLH/SSM output at time $t$
@@
 $\mathbf{E}$ $\mathbb{R}^{P\times R}$ LLH mark embedding matrix ($P\times R$ in low-rank factorization)
-[…]
-$\bm{\alpha}$ $\mathbb{R}^{R\times K}$ (For DLHP) Mark impulses ($R\times K$ in low-rank factorization)
+$L$ $\mathbb{Z}^{+}$ Number of linear recurrences in a S2P2 model; model “depth”
+$\bm{\alpha}$ $\mathbb{R}^{R\times K}$ (For S2P2) Mark impulses ($R\times K$ in low-rank factorization)
 $\sim$ N/A Tilde (e.g., $\tilde{\mathbf{B}}$) denotes variable is in the diagonalized eigenbasis
@@
 $\bar{\Lambda}$ $\mathbb{C}^{P\times P}$ Discretized diagonal dynamics matrix
-${(l)}$ N/A Superscript index in parenthesis indicates layer (i.e., $\mathbf{x}$ for layer $l$)
+(l){(l)} N/A Superscript index in parenthesis indicates layer (i.e., $\mathbf{x}$ for layer $l$)
 Table 4: Key acronyms used throughout this paper.
 Acronym Page number Definition
-CNN 6 Convolutional neural network
-LHP 1 Linear Hawkes process
+CNN 7 Convolutional neural network
+LHP 2 Linear Hawkes process
 LLH 2 Latent linear Hawkes
@@
 ZOH 5 Zero-order hold
-RMTPP 7 Recurrent marked temporal point process (Du et al., 2016)
-NHP 1 Neural Hawkes process (Mei & Eisner, 2017)
-SAHP 7 Self-attentive Hawkes process (Zhang et al., 2020)
-THP 7 Transformer Hawkes process (Zuo et al., 2020)
-AttNHP 7 Attentive neural Hawkes process (Yang et al., 2022)
-IFTPP 7 Intensity-free temporal point process (Shchur et al., 2020a)
-DLHP 1 Deep linear Hawkes process (ours)
+RMTPP 2 Recurrent marked temporal point process [Du et al., 2016]
+NHP 2 Neural Hawkes process [Mei and Eisner, 2017]
+SAHP 2 Self-attentive Hawkes process [Zhang et al., 2020]
+THP 2 Transformer Hawkes process [Zuo et al., 2020]
+AttNHP 2 Attentive neural Hawkes process [Yang et al., 2022]
+IFTPP 2 Intensity-free temporal point process [Shchur et al., 2020a]
+MHP 2 Mamba Hawkes process [Gao et al., 2024]
+S2P2 1 State-space point process (ours)
 Appendix B Additional Details on Methods
-B.1 DLHP Algorithms
-Algorithm 1 Deep Linear Hawkes Process: Get Right State Limits
-1:DLHP layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, event intervals $\Delta t_{1:N}$, nonlinearity $\sigma$, shared mark embeddings $\bm{\alpha}_{1:N}$.
+B.1 State-Space Point Process Algorithms
+Algorithm 1 State-Space Point Process: Get Right State Limits
+1:S2P2 layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, event intervals $\Delta t_{1:N}$, nonlinearity $\sigma$, shared mark embeddings $\bm{\alpha}_{1:N}$.
 2:Right state limits $\mathbf{x}_{t_{1:N}}^{(1:L)}$
-3:$\mathbf{u}_{t_{1:N}-}=\mathbf{0}$triangleright Left input limits
+[…]
 4:for $l$ in $1:L$ do
-5:     $\bar{\bm{\Lambda}}^{(l)}_{1:N}=\mathrm{Discretize}\left(\bm{\Lambda}^{(l)},\Delta t_{1:N}\right)$triangleright Zero-order hold, see Eq. 22
-6:     $\tilde{\mathbf{x}}^{(l)}_{t_{1:N}}=\mathrm{ParallelScan}\left(\bar{\bm{\Lambda}}^{(l)}_{1:N},(\bar{\bm{\Lambda}}^{(l)}_{1:N}-\mathbf{I})\tilde{\mathbf{B}}^{(l)}\mathbf{u}_{t_{1:N}-}+\tilde{\mathbf{E}}^{(l)}\bm{\alpha}_{1:N}\right)$triangleright Compute right $x$ limits
-7:     $\tilde{\mathbf{x}}^{(l)}_{t_{1:N}-}=\tilde{\mathbf{x}}^{(l)}_{t_{1:N}}-\tilde{\mathbf{E}}^{(l)}\bm{\alpha}_{1:N}$triangleright Compute left $x$ limits
-8:     $\mathbf{u}_{t_{1:N}-}=\mathrm{LayerNorm}\left(\sigma\left(\tilde{\mathbf{C}}^{(l)}\tilde{\mathbf{x}}_{t_{1:N}-}+\mathbf{D}^{(l)}\mathbf{u}_{t_{1:N}-}\right)+\mathbf{u}_{t_{1:N}-}\right)$triangleright Compute next layer’s left $u$ limits
+5:  $\bar{\bm{\Lambda}}^{(l)}_{1:N}=\mathrm{Discretize}\left(\bm{\Lambda}^{(l)},\Delta t_{1:N}\right)$ $\triangleright$ Zero-order hold, see Eq.˜19
+6:  $\tilde{\mathbf{x}}^{(l)}_{t_{1:N}}=\mathrm{ParallelScan}\left(\bar{\bm{\Lambda}}^{(l)}_{1:N},(\bar{\bm{\Lambda}}^{(l)}_{1:N}-\mathbf{I})\tilde{\mathbf{B}}^{(l)}\mathbf{u}_{t_{1:N}-}+\tilde{\mathbf{E}}^{(l)}\bm{\alpha}_{1:N}\right)$ $\triangleright$ Compute right $x$ limits
+7:  $\tilde{\mathbf{x}}^{(l)}_{t_{1:N}-}=\tilde{\mathbf{x}}^{(l)}_{t_{1:N}}-\tilde{\mathbf{E}}^{(l)}\bm{\alpha}_{1:N}$ $\triangleright$ Compute left $x$ limits
+8:  $\mathbf{u}_{t_{1:N}-}=\mathrm{LayerNorm}\left(\sigma\left(\tilde{\mathbf{C}}^{(l)}\tilde{\mathbf{x}}_{t_{1:N}-}+\mathbf{D}^{(l)}\mathbf{u}_{t_{1:N}-}\right)+\mathbf{u}_{t_{1:N}-}\right)$ $\triangleright$ Compute next layer’s left $u$ limits
 9:end for
 10:return $\mathbf{x}_{t_{1:N}}^{(1:L)}$
-Algorithm 2 Deep Linear Hawkes Process: Get Intensity From Right Limit
-1:DLHP layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, Previous state right limits $\mathbf{x}_{t}^{(1:L)}$, Integration period $\delta t$, nonlinearity $\sigma$, Intensity function $\mathrm{IntensityFn}$.
+Algorithm 2 State-Space Point Process: Get Intensity From Right Limit
+1:S2P2 layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, Previous state right limits $\mathbf{x}_{t}^{(1:L)}$, Integration period $\delta t$, nonlinearity $\sigma$, Intensity function $\mathrm{IntensityFn}$.
 2:Intensity left limit $\bm{\lambda}_{t+\delta t}$
-3:$\mathbf{u}_{t+\delta t-}=\mathbf{0}$triangleright Left input limit
+[…]
 4:for $l$ in $1:L$ do
-5:     $\bar{\bm{\Lambda}}^{(l)}=\mathrm{Discretize}\left(\bm{\Lambda}^{(l)},\delta t\right)$triangleright Zero-order hold, see Eq. 22
-6:     $\tilde{\mathbf{x}}^{(l)}_{t+\delta t-}=\bar{\bm{\Lambda}}^{(l)}\mathbf{x}_{t}^{(l)}+(\bar{\bm{\Lambda}}^{(l)}-\mathbf{I})\tilde{\mathbf{B}}^{(l)}\mathbf{u}_{t+\delta t-}$triangleright Evolve state
-7:     $\mathbf{u}_{t+\delta t-}=\mathrm{LayerNorm}\left(\sigma\left(\tilde{\mathbf{C}}^{(l)}\tilde{\mathbf{x}}_{t+\delta t-}^{(l)}+\mathbf{D}^{(l)}\mathbf{u}_{t+\delta t-}\right)+\mathbf{u}_{t+\delta t-}\right)$triangleright Compute event left $u$ limits
+5:  $\bar{\bm{\Lambda}}^{(l)}=\mathrm{Discretize}\left(\bm{\Lambda}^{(l)},\delta t\right)$ $\triangleright$ Zero-order hold, see Eq.˜19
+6:  $\tilde{\mathbf{x}}^{(l)}_{t+\delta t-}=\bar{\bm{\Lambda}}^{(l)}\mathbf{x}_{t}^{(l)}+(\bar{\bm{\Lambda}}^{(l)}-\mathbf{I})\tilde{\mathbf{B}}^{(l)}\mathbf{u}_{t+\delta t-}$ $\triangleright$ Evolve state
+7:  $\mathbf{u}_{t+\delta t-}=\mathrm{LayerNorm}\left(\sigma\left(\tilde{\mathbf{C}}^{(l)}\tilde{\mathbf{x}}_{t+\delta t-}^{(l)}+\mathbf{D}^{(l)}\mathbf{u}_{t+\delta t-}\right)+\mathbf{u}_{t+\delta t-}\right)$ $\triangleright$ Compute event left $u$ limits
 8:end for
-9:$\bm{\lambda}_{t+\delta t}=\mathrm{IntensityFn}(\mathbf{u}_{t+\delta t-})$triangleright Rectify intensity, see Eq. 13
+[…]
 10:return $\bm{\lambda}_{t+\delta t}$
-Algorithm 3 Deep Linear Hawkes Process: Compute Log-Likelihood
-1:DLHP layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, Event times $t_{1:N}$, mark types $k_{1:N}$, nonlinearity $\sigma$, shared mark embedding function $\mathrm{EmbedMarks}$, number of integration points per event $M$, Intensity function $\mathrm{IntensityFn}$.
-2:Log-ikelihood $\mathcal{L}$
+Algorithm 3 State-Space Point Process: Compute Log-Likelihood
+1:S2P2 layer parameters $\bm{\Theta}=\left\{\bm{\Lambda}^{(l)},\tilde{\mathbf{B}}^{(l)},\tilde{\mathbf{C}}^{(l)},\mathbf{D}^{(l)},\tilde{\mathbf{E}}^{(l)},\tilde{\mathbf{x}}_{0}^{(l)}\right\}_{l=1}^{L}$, Event times $t_{1:N}$, mark types $k_{1:N}$, nonlinearity $\sigma$, shared mark embedding function $\mathrm{EmbedMarks}$, number of integration points per event $M$, Intensity function $\mathrm{IntensityFn}$.
+2:Log-likelihood $\mathcal{L}$
 […]
-4:$t_{0}:=0$
-5:$\Delta t_{1:N}=t_{1:N}-t_{0:N-1}$
-6:$s_{1:N,1:M}\sim\mathcal{U}(0,\Delta t_{1:N})$triangleright Sample $M$ integration points per interval (non-inclusive)
-7:$\tilde{\mathbf{x}}_{t_{1:N}}^{(1:L)}=\mathrm{GetRightStateLimits}(\Theta,\Delta t_{1:N},\sigma,\bm{\alpha}_{1:N})$triangleright Algorithm 1, $\mathcal{O}(\log N)$ parallel time
+[…]
+[…]
+[…]
+[…]
 8:for $n$ in $1:N$ do $\triangleright$ This is embarrassingly parallelizable with vmap, $\mathcal{O}(1)$ parallel time
-9:     $\bm{\lambda}_{t_{n}}=\mathrm{GetIntensityFromRightLimit}\left(\Theta,\tilde{\mathbf{x}}_{t_{n}}^{(1:L)},\Delta t_{n},\sigma,\mathrm{IntensityFn}\right)$triangleright Algorithm 2, $\mathcal{O}(1)$ parallel time
+9:  $\bm{\lambda}_{t_{n}}=\mathrm{GetIntensityFromRightLimit}\left(\Theta,\tilde{\mathbf{x}}_{t_{n}}^{(1:L)},\Delta t_{n},\sigma,\mathrm{IntensityFn}\right)$ $\triangleright$ Algorithm 2, $\mathcal{O}(1)$ parallel time
 10:  for $m$ in $1:M$ do $\triangleright$ This is embarrassingly parallelizable with vmap, $\mathcal{O}(1)$ parallel time
-11:         $\bm{\lambda}_{s_{n,m}}=\mathrm{GetIntensityFromRightLimit}\left(\Theta,\tilde{\mathbf{x}}_{t_{n}}^{(1:L)},s_{n,m},\sigma,\mathrm{IntensityFn}\right)$triangleright Algorithm 2, $\mathcal{O}(1)$ parallel time
+11:   $\bm{\lambda}_{s_{n,m}}=\mathrm{GetIntensityFromRightLimit}\left(\Theta,\tilde{\mathbf{x}}_{t_{n}}^{(1:L)},s_{n,m},\sigma,\mathrm{IntensityFn}\right)$ $\triangleright$ Algorithm 2, $\mathcal{O}(1)$ parallel time
 12:  end for
 13:end for
-14:$\mathcal{L}=\sum_{n=1}^{N}\log\lambda_{t_{n}}^{k_{n}}+\sum_{n=1}^{N}\frac{\Delta t_{n}}{M}\sum_{m=1}^{M}\sum_{k=1}^{K}\lambda^{k}_{s_{n,m}}$triangleright Eq.
-2 with Monte-Carlo approximation of integral
+[…]
 15:return $\mathcal{L}$
@@
 We use the zero-order hold (ZOH) discretization method, to convert the continuous-time linear recurrence into a sequence of closed-form updates, given the integration times, that can also be efficiently computed.
-We refer the reader to Iserles (2009) for a comprehensive introduction to the ZOH transform.
+We refer the reader to Iserles [2009] for a comprehensive introduction to the ZOH transform.
 The main assumption of the ZOH discretization is that the input signal is held constant over the time period being integrated.
 Under this assumption, it is possible to solve for the dynamics and input matrices that yield the correct state at the end of the integration period.
-For the LLH dynamics in Eq.
-10, when no events occur in $(t,t^{\prime})$, this becomes
+For the LLH dynamics in Eq.˜8, when no events occur in $(t,t^{\prime})$, this becomes
 $\mathbf{x}_{t^{\prime}-}=\int_{t}^{t^{\prime}}\mathbf{A}\mathbf{x}_{t}+\mathbf{AB}\mathbf{u}_{t}\mathrm{d}t=\overline{\mathbf{A}}\mathbf{x}_{t}+\overline{\mathbf{AB}}\mathbf{u}_{t}\quad\mathrm{assuming}\quad\mathrm{d}\mathbf{u}_{t}=\mathbf{0}\in[t,t^{\prime}],$ (15)
@@
 To compute the matrices $\overline{\mathbf{A}}$ and $\overline{\mathbf{AB}}$ however requires computing a matrix exponential and a matrix inverse.
-However, Smith et al.
-(2022) avoid this by diagonalizing the system (also avoiding a dense matrix-matrix multiplication in the parallel scan).
+Smith et al.
+[2022] avoid this by diagonalizing the system (also avoiding a dense matrix-matrix multiplication in the parallel scan).
 The diagonalized dynamics and input matrices are denoted $\bm{\Lambda}$ (a diagonal matrix) and $\bm{\Lambda}\tilde{\mathbf{B}}$ respectively.
-In this case, Eq.
-19 reduces to
+In this case, Eq.˜16 reduces to
 $\overline{\mathbf{A}}$ $=e^{\bm{\Lambda}\Delta t},$ (17)
 $\overline{\mathbf{AB}}$ $=\bm{\Lambda}^{-1}(e^{\bm{\Lambda}\Delta t}-\mathbf{I})\bm{\Lambda}\tilde{\mathbf{B}}$ (18)
-[…]
+$=(e^{\bm{\Lambda}\Delta t}-\mathbf{I})\tilde{\mathbf{B}}\quad\text{(diagonal matrices commute)}$ (19)
 where $e^{\bm{\Lambda}\Delta t}$ is trivially computable as the exponential of the leading diagonal of $\bm{\Lambda}\Delta t$.
 These operations are embarrassingly parallelizable across the sequence length and state dimension given the desired evaluation times.
-To contextualize, suppose an event occurs at time $t$, Eq.
-22 allows us to exactly (under the constant-input assumption) efficiently evaluate the linear recurrence at subsequent times $t^{\prime}$.
-We use this extensively in the DLHP to efficiently evaluate the recurrence (and hence the intensity) at the irregularly-spaced event times and times used to compute the integral term.
+To contextualize, suppose an event occurs at time $t$, Eq.˜19 allows us to exactly (under the constant-input assumption) efficiently evaluate the linear recurrence at subsequent times $t^{\prime}$.
+We use this extensively in S2P2 to efficiently evaluate the recurrence (and hence the intensity) at the irregularly-spaced event times and times used to compute the integral term.
 It should be noted the discretization was done to compute a left-limit $\mathbf{x}_{t^{\prime}-}$ from a previous right-limit $\mathbf{x}_{t}$.
@@
 If an event does occur at time $t^{\prime}$ with mark $k$, then the left-limit $\mathbf{x}_{t^{\prime}-}$ can be incremented by $\tilde{\mathbf{E}}\bm{\alpha}_{k}$ to compute $\mathbf{x}_{t^{\prime}+}=\mathbf{x}_{t^{\prime}}$.
-This increment from left- to right-limit is exact and leverages no discretization assumption.
+This increment is exact and leverages no discretization assumption.
 B.3 Interpretation for Input-Dependent Dynamics
-Consider the input-dependent recurrence for an LLH layer, as defined in Eq. 17:
+Consider the input-dependent recurrence for an LLH layer, as defined in Eq.˜13:
 $\mathrm{d}\tilde{\mathbf{x}}_{t}$ $:=\bm{\Lambda}_{i}\tilde{\mathbf{x}}_{t-}\mathrm{d}t+\bm{\Lambda}_{i}\tilde{\mathbf{B}}\mathbf{u}_{t-}\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (20)
@@
 B.4 Forwards and Backwards Zero Order Hold Discretization
-In Section 3.3 we highlighted that the ZOH discretization is exact when $\mathbf{u}_{t}$ is held constant over the integration window.
-This raises a unique design question for DLHPs: what constant value should $\mathbf{u}_{t}$ take on when evolving $\mathbf{x}$ from time $t$ to $t^{\prime}$?
+In Section˜3.2 we highlighted that the ZOH discretization is exact when $\mathbf{u}_{t}$ is held constant over the integration window.
+This raises a unique design question for S2P2: what constant value should $\mathbf{u}_{t}$ take on when evolving $\mathbf{x}$ from time $t$ to $t^{\prime}$?
 For the first layer of the model, the input is zero by construction, so there is no choice to be made—in fact, since $\mathbf{u}$ is constant for the first layer the updates are exact.
@@
 This should be a value in (or function of) $\{\mathbf{u}_{s}\;|\;s\in[t,t^{\prime})\}$.
-Note this is because the value at $t^{\prime}$, $\mathbf{u}_{t^{\prime}}$, cannot be incorporated as this would cause a data leakage in our model; while values prior to $t$ would discard the most recent mark.
+Note this is because the value at $t^{\prime}$, $\mathbf{u}_{t^{\prime}}$, cannot be incorporated as this would cause a data leakage in our model; while values prior to $t$ would discard the most recent event occurrence.
 For this work, we explore two natural choices: (i) the input value at the beginning of the interval, $\mathbf{u}_{t}$, and (ii) the left-limit at the end of the interval, $\mathbf{u}_{t^{\prime}-}$.
-We illustrate the backwards variant in Fig.
-2, where in the rightmost panel, we use the $\mathbf{u}_{t^{*}}$ values at each layer, as opposed to $\mathbf{u}_{t_{3}}$.
+Note the end of the interval need not align with an event (crucial for when computing intermediate intensity values).
 We refer to these options as forwards and backwards ZOH, respectively.
+We illustrate the backwards variant in Fig.˜2, where in the leftmost panel, we use the $\mathbf{u}_{t-}$ values at each layer to calculate $\mathbf{x}_{t-}$ and $\mathbf{x}_{t}$, as opposed to $\mathbf{u}_{t_{i}}$ for $t_{i}<t<t_{i+1}$.
 All experiments in the main paper utilize backwards ZOH.
 It is not obvious a priori which one of these modes is more performant.
-We therefore conducted an ablation experiment in Table 13.
+We therefore conducted an ablation experiment in Table˜14.
 We see that there is little difference between the two methods.
 We also note that models are learned through this discretization, and so this decision does not mean that a model is “incorrectly discretized” one way or the other, but instead they define subtlety different families of models.
-Theoretical and empirical investigation of the interpretations of this choice is an interesting area of investigation going forwards, extending the ablations we present in Table 13.
+Theoretical and empirical investigation of the interpretations of this choice is an interesting area of investigation going forwards, extending the ablations we present in Table˜14.
 B.5 Theoretical Complexity
-We include in Table 4 a brief summary of the theoretical complexity of each of the methods we consider.
-We break these down by the work, memory complexity and theoretical best parallel application time of the forward pass (used when conditioning on a sequence, the left-hand term of Eq.
-2) and evaluating the integral term in Eq.
-2 given that the forward pass has been completed (as this is either required by the method, and is nearly always evaluated in conjunction with the forward pass).
+We include in Table˜5 a brief summary of the theoretical complexity of each of the methods we consider, broadly grouped by their architectures.
+We analyze complexity by the work, memory complexity and theoretical best parallel application time of the forward pass (used when conditioning on a sequence, the left-hand term of Eq.˜2) and evaluating the integral term in Eq.˜2 given that the forward pass has been completed (as this is either required by the method, and is nearly always evaluated in conjunction with the forward pass).
 We then state the limiting best-case theoretical parallelism of the two components.
-The reasoning behind this is as as follows:
+The reasoning behind the calculated values are as follows:
   • The forward pass of RMTPP, NHP and IFTPP use non-linear RNNs, and hence incur memory and work that is linear in the sequence length, and cannot be parallelized.
-However, they re-use the computed hidden states to compute the integral term, and hence, while they incur work and memory that scales in the sequence length and number of events, this work can be perfectly parallelized.
-This results in a best-case parallelism of $\mathcal{O}(L)$ (dominated by the forward pass).
+MHP uses an RNN, but that is logarithmically parallelizable.
+These models re-use the computed hidden states to compute the integral term, and hence, while they incur work and memory that scales in the sequence length and number of events, this work can be perfectly parallelized.
+This results in a best-case parallelism of $\mathcal{O}(N)$ (dominated by the forward pass; $\mathcal{O}(\log N)$ for MHP).
   • SAHP, THP and AttNHP all use self-attention, and hence have a work and memory that scales quadratically in the sequence length, although this work can be parallelized across the sequence length, resulting in logarithmic parallel depth.
@@
 AttNHP is slightly different in that it re-applies the whole independently attention mechanism for each integration point.
-However, this work is parallelizeable and hence still reduces to a best-case depth of $\mathcal{O}(\log L)$.
-  • DLHP is an RNN and hence has linear work and memory in the forward pass, but can be parallelized to a best-case depth of $\mathcal{O}(\log L)$ using the parallel scan.
-We then re-use the states computed in the forward pass for estimating the integral, which, as with the other RNN methods, is perfectly parallelizable, resulting in a theoretical parallel depth of $\mathcal{O}(\log L)$.
+However, this work is parallelizeable and hence still reduces to a best-case depth of $\mathcal{O}(\log N)$.
+  • S2P2 is an RNN and hence has linear work and memory in the forward pass, but can be parallelized to a best-case depth of $\mathcal{O}(\log N)$ using the parallel scan.
+We then re-use the states computed in the forward pass for estimating the integral, which, as with the other RNN methods, is perfectly parallelizable, resulting in a theoretical parallel depth of $\mathcal{O}(\log N)$.
 Note that these figures do not account for the number of layers required by each model, which must be evaluated in sequence.
 Table 5: Comparison of methods based on memory and compute complexity.
-We see that our DLHP matches the best performing baseline in all categories.
-$L$ denotes to the sequence length, and $M$ denotes to the number of Monte Carlo grid points per-event used in evaluating Eq.
-2.
+We see that our S2P2 matches the best performing baseline in all categories.
+NN denotes to the sequence length, and $M$ denotes to the number of Monte Carlo grid points per-event used in evaluating Eq.˜2.
 As IFTPP is an intensity-free method, it does not need to estimate $\int\lambda_{t}\mathrm{d}{}t$ as the other methods do.
@@
 Parallelism Parallelism Parallelism
-RMTPP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(L)$
-NHP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(L)$
-SAHP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
-THP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
-AttNHP $\mathcal{O}(L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(\log L)$ $\mathcal{O}(L^{2}M)$ $\mathcal{O}(L^{2}M)$ $\mathcal{O}(\log L)$ $\mathcal{O}(\log L)$
-IFTPP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(L)$ N/A N/A N/A $\mathcal{O}(L)$
-DLHP $\mathcal{O}(L)$ $\mathcal{O}(L)$ $\mathcal{O}(\log L)$ $\mathcal{O}(LM)$ $\mathcal{O}(LM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log L)$
+RMTPP $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(N)$
+NHP $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(N)$
+MHP $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(\log N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log N)$
+IFTPP $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(N)$ N/A N/A N/A $\mathcal{O}(N)$
+SAHP $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(\log N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log N)$
+THP $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(\log N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log N)$
+AttNHP $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(\log N)$ $\mathcal{O}(N^{2}M)$ $\mathcal{O}(N^{2}M)$ $\mathcal{O}(\log N)$ $\mathcal{O}(\log N)$
+S2P2 $\mathcal{O}(N)$ $\mathcal{O}(N)$ $\mathcal{O}(\log N)$ $\mathcal{O}(NM)$ $\mathcal{O}(NM)$ $\mathcal{O}(1)$ $\mathcal{O}(\log N)$
+To validate the scaling properties, we measure the wallclock time for a full forward pass and log-likelihood evaluation on random input sequences with lengths ranging from eight events to over half a million events.
+The architectures and mark spaces are the same as in the StackOverflow experiments (see Tables˜6(a) and 7).
+We note: our EasyTPP PyTorch S2P2 is written in pure PyTorch, and hence is not as optimized as other methods (compared to, for instance, IFTPP, which uses a GPU-optimized implementation of the GRU).
+We therefore include the runtimes of a standalone JAX S2P2 implementation, which allows for comparable levels of optimization through JIT compilation.
+We observe the predicted scaling in practice.
+NHP scales linearly across all sequence lengths, and is far outpaced by all other methods.
+The THP scales well before reverting to superlinear scaling, and then runs out of memory.
+The IFTPP is very fast at shorter runtimes, but quickly reverts to linear scaling, due to its simple but highly optimized implementation and inherently sequential operation.
+Both S2P2 implementations scale linearly at long sequence lengths, but have near-constant runtime at shorter sequences.
+At shorter sequence lengths, the more optimized JAX implementation is faster than the unoptimized pure PyTorch implementation.
+While this indicates that there are additional opportunities to accelerate the PyTorch implementation further (e.g., exploiting kernel fusion or writing a lower-level Triton implementation), these results still confirm that our S2P2 can exploit parallel scans to scale to long sequences more effectively than other methods while retaining strong performance.
+Figure 5: Median runtime of various models against increasing sequence lengths when conditioning on a sequence (Algorithm˜1) and for likelihood evaluation (Algorithm˜3) over 20 random seeds (variance negligible).
+S2P2 is faster across a wide range of sequence lengths.
+Crosses indicate where THP runs out of memory or IFTPP throws an error.
 Appendix C Experimental Configurations and Datasets
 C.1 Training Details & Hyperparameter Configurations
+All baseline models used up-to-date PyTorch implementations, provided by the EasyTPP library [Xue et al., 2023] as of May 2025.
 We apply a grid search for all models on all datasets for hyperparameter tuning.
@@
 For models/datasets that require more memory (e.g., large mark space or long sequences), we reduce the batch size and keep them as consistent as possible among all the models on each dataset.
-We use the Adam stochastic gradient optimizer (Kingma & Ba, 2015), with a learning rate of 0.01 and a linear warm-up schedule over the first 1% iterations, followed by a cosine decay.
+We use the Adam stochastic gradient optimizer [Kingma and Ba, 2015], with a learning rate of 0.01 and a linear warm-up schedule over the first 1% iterations, followed by a cosine decay.
 Initial experiments showed this setting generally worked well across different models and datasets leads to convergence within 300 epochs.
@@
 For SAHP, THP, and AttNHP, we searched over all combinations of number of $L$ = {1, 2, 3}, hidden state size = {16, 32, 64, 128}, and number of heads = {1, 2, 4}.
-Finally, for DLHP, we considered combinations for number of layers = {1, 2, 3, 4}, $p$ = {16, 32, 64, 128} and $h$ = {16, 32, 64, 256}.
-We fixed the activation function as GeLU (Hendrycks & Gimpel, 2016) and apply post norm with layer norm (Ba, 2016).
-We fix the dropout as 0.1 for DLHP on the five core benchmark datasets, and add dropout = {0, 0.1} to the grid search for the other three datasets.
+For MHP, we followed their paper to fix $L=4$, then we search over $h$ = {4, 8, 16, 32, 64, 128, 256, 512}.
+Finally, for S2P2, we considered combinations for number of layers = {1, 2, 3, 4}, $p$ = {16, 32, 64, 128} and $h$ = {16, 32, 64, 256}.
+We generally found a range of reasonable hyperparameters to yield similar performance on S2P2, while multiple layers were key for performance (intuitively deep stack allows nonlinear and complex dynamics), but there was not a critical dependence on depth.
+We fixed the activation function as GeLU [Hendrycks and Gimpel, 2016] and apply post norm with layer norm [Ba, 2016].
+We fix the dropout as 0.1 for S2P2 on the five core benchmark datasets, and add dropout = {0, 0.1} to the grid search for the other three datasets.
 Due to the scale of Last.fm and EHRSHOT datasets, we perform a smaller search over architectures that roughly match the parameter counts for all models at three levels: 25k, 50k, 200k, and choose the model with the best validation results.
@@
 We were unable to train any AttNHP on EHRSHOT.
-The final model architectures used are reported in Table 5(a) and Table 5(b).
-These configurations are also included in the supplementary code we include.
-Table 5: Model architectures for the experiments presented in Table 1
-(a) Model architectures for the five EasyTPP benchmark datasets.
+The final model architectures used are reported in Tables˜6(a) and 6(b).
+Table 6: Model architectures for the five EasyTPP benchmark datasets in Table˜2.
+(a) Model architectures for the five EasyTPP benchmark datasets in Table˜2.
 Model Amazon Retweet Taxi Taobao StackOverflow
 RMTPP $h=128$ $h=16$ $h=128$ $h=16$ $h=256$
-NHP $h=128$ $h=64$ $h=128$ $h=128$ $h=64$
 SAHP $h=32,l=2,\mathrm{heads}=2$ $h=32,l=3,\mathrm{heads}=4$ $h=16,l=2,\mathrm{heads}=4$ $h=32,l=1,\mathrm{heads}=1$ $h=64,l=1,\mathrm{heads}=1$
 THP $h=32,l=2,\mathrm{heads}=4$ $h=16,l=3,\mathrm{heads}=4$ $h=128,l=1,\mathrm{heads}=4$ $h=64,l=1,\mathrm{heads}=1$ $h=16,l=2,\mathrm{heads}=4$
+IFTPP $h=64$ $h=64$ $h=32$ $h=64$ $h=64$
+MHP $h=8$ $h=16$ $h=4$ $h=16$ $h=8$
+NHP $h=128$ $h=64$ $h=128$ $h=128$ $h=64$
 AttNHP $h=64,t=16,l=2,\mathrm{heads}=4$ $h=16,t=16,l=2,\mathrm{heads}=4$ $h=16,t=16,l=3,\mathrm{heads}=4$ $h=32,t=16,l=3,\mathrm{heads}=4$ $h=32,t=16,l=2,\mathrm{heads}=4$
-IFTPP $h=64$ $h=64$ $h=32$ $h=64$ $h=64$
-DLHP $h=64,p=128,l=2$ $h=128,p=128,l=2$ $h=128,p=16,l=4$ $h=32,p=16,l=4$ $h=32,p=32,l=3$
-(b) Model architectures for the additional three benchmark datasets.
+S2P2 $h=64,p=128,l=2$ $h=128,p=128,l=2$ $h=128,p=16,l=4$ $h=32,p=16,l=4$ $h=32,p=32,l=3$
+(b) Model architectures for the additional three benchmark datasets in Table˜2.
 Model Last.fm MIMIC-II EHRSHOT
 RMTPP $h=256$ $h=128$ $h=16$
-NHP $h=112$ $h=128$ $h=80$
 SAHP $h=136,l=2,\mathrm{heads}=4$ $h=64,l=2,\mathrm{heads}=4$ $h=8,l=2,\mathrm{heads}=4$
 THP $h=48,l=2,\mathrm{heads}=4$ $h=32,l=3,\mathrm{heads}=4$ $h=32,l=2,\mathrm{heads}=4$
+IFTPP $h=48$ $h=256$ $h=16$
+MHP $h=16$ $h=16$ $h=32$
+NHP $h=112$ $h=128$ $h=80$
 AttNHP $h=28,t=16,l=2,\mathrm{heads}=4$ $h=64,t=16,l=3,\mathrm{heads}=2$ OOM
-IFTPP $h=48$ $h=256$ $h=16$
-DLHP $h=68,p=16,l=2$ $h=256,p=64,l=2$ $h=128,p=32,l=2$
+S2P2 $h=68,p=16,l=2$ $h=64,p=16,l=2$ $h=128,p=32,l=2$
 C.2 Dataset Statistics
-We report the statistics of all eight datasets we used in Table 6.
+We report the statistics of all eight datasets we used in Table˜7.
 We used the HuggingFace version of the five EasyTPP datasets.
 For all datasets, we further ensure the MTPP modeling assumptions are satisfied that no more than two events occur at the same time (i.e., inter-arrival time is strictly positive), and event times do not lie on grid points that are effectively discrete-time events.
-Dataset descriptions and pre-processing details are provided in Section C.3.
+Dataset descriptions and pre-processing details are provided in Section˜C.3.
 Table 7: Statistics of the eight datasets we experiment with.
@@
 For MIMIC-II, we copy Du et al.
-(2016) and keep the 325 test sequences in the test split, and further split the 2,935 training sequences into 2,600 for training and 325 for validation.
+[2016] and keep the 325 test sequences in the test split, and further split the 2,935 training sequences into 2,600 for training and 325 for validation.
 In our pre-processed datasets, Last.fm and EHRSHOT, we randomly partition into subsets containing 70%, 15%, 15% of all sequences for training/validation/test respectively.
 We provide a high-level description of all the datasets we used, followed by our pre-processing procedure of Last.fm and EHRSHOT in more detail.
-Note that for datasets that contain concurrent events or effectively discrete times, we apply a small amount of jittering to ensure no modeling assumptions are violated in the MTPP framework.
-Amazon (Ni et al., 2019) contains user product reviews where product categories are considered as marks.
-Retweet (Zhao et al., 2015) predicts the popularity of a retweet cascade, where the event type is decided by if the retweet comes from users with “small”, “medium”, or “large” influences, measured by number of followers (Mei & Eisner, 2017).
-Taxi data (Whong, 2014; Mei et al., 2019) uses data from the pickups and dropoffs of New York taxi and the marks are defined as the Cartesian product of five discrete locations and two actions (pickup/dropoff).
-Taobao (Xue et al., 2022) describes the viewing patterns of users on an e-commerce site, where item categories are considered as marks.
+Note that for datasets that contain concurrent events or effectively discrete times (e.g., StackOverflow, Retweet), we apply a small amount of jittering to ensure no modeling assumptions are violated in the MTPP framework.
+Amazon [Ni et al., 2019] contains user product reviews where product categories are considered as marks.
+Retweet [Zhao et al., 2015] predicts the popularity of a retweet cascade, where the event type is decided by if the retweet comes from users with “small”, “medium”, or “large” influences, measured by number of followers [Mei and Eisner, 2017].
+Taxi data [Whong, 2014, Mei et al., 2019] uses data from the pickups and dropoffs of New York taxi and the marks are defined as the Cartesian product of five discrete locations and two actions (pickup/dropoff).
+Taobao [Xue et al., 2022] describes the viewing patterns of users on an e-commerce site, where item categories are considered as marks.
 StackOverflow contains the badges (defined as marks) awarded to users on a question-answering website.
-Finally, MIMIC-II (Saeed et al., 2002) records different diseases (used as marks) during hospital visits of patients.
+Finally, MIMIC-II [Saeed et al., 2002] records different diseases (used as marks) during hospital visits of patients.
 We add a small amount of noise to the MIMIC-II event times so that events do not lie on a fixed grid.
 Both StackOverflow and MIMIC-II datasets were first pre-processed by Du et al.
-(2016).
+[2016].
 Last.fm Celma Herrada et al.
-(2009); McFee et al.
-(2012) records 992 users’ music listening habits that has been widely used in MTPP literature (Kumar et al., 2019; Boyd et al., 2020; Bosser & Taieb, 2023).
+[2009], McFee et al.
+[2012] records 992 users’ music listening habits that has been widely used in MTPP literature [Kumar et al., 2019, Boyd et al., 2020, Bosser and Taieb, 2023].
 Mark types are defined as the genres of a song, and each event is a play of a particular genre.
@@
 EHRSHOT Wornow et al.
-(2023) is a newly proposed large dataset of longitudinal de-identified patient medical records, and has rich information such as hospital visits, procedures, and measurements.
+[2023] is a newly proposed large dataset of longitudinal de-identified patient medical records, and has rich information such as hospital visits, procedures, and measurements.
 We introduce an MTPP dataset derived from EHRSHOT, where medical services and procedures are treated as marks, as identified by Current Procedural Terminology (CPT-4) codes.
 Each patient defines an event sequence, and we retain only CPT-4 codes with at least 100 occurrences in the dataset.
-For the $<1\%$ events of events where there are more than 10 codes at a single timestamp, we retain the top 10 codes with the most frequencies and discard the rest.
+For the $1\%$ events of events where there are more than 10 codes at a single timestamp, we retain the top 10 codes with the most frequencies and discard the rest.
 We then add a small amount of random noise to the event time to ensure they are not overlapping.
-This process ensures we still satisfy the MTPP framework, and can reasonably instead compute top-10 accuracy for the next mark prediction.
-Other work has considered extending the MTPP framework to consider simultaneous event occurrence (Chang et al., 2024).
-Then we standardize each sequence to start and end with start and end of a sequence events.
+This process ensures we still satisfy the MTPP framework, and can reasonably instead compute top-10 accuracy for the next mark prediction; other work has considered extending the MTPP framework to consider simultaneous event occurrence [Chang et al., 2024].
+Then we standardize each sequence to start at $t=0$ and pad the start and end of sequences with a specific padded event token.
 Note that we do not score these events.
 Event times are normalized to be in hours.
-We discard sequences that have less than 5 events and a single timestamp.
-This leads to the final version of our dataset to have 668 marks, and the sequence lengths range from 5 to 3955 events, reflecting patient histories that can span multiple years.
+We discard sequences that have less than 5 events or a single timestamp.
+This leads to the final version of our dataset having 668 marks, and the sequence lengths range from 5 to 3955 events, reflecting patient histories that can span multiple years.
+We include code in our forked repository for preparing the EHRSHOT event sequence dataset from the raw EHRSHOT dataset.
+Note that we cannot distribute the raw data (or derivative dataset) under the terms of the original EHRSHOT dataset requiring credentialed access through PhysioNet.
 Appendix D Additional Experimental Results
 D.1 Full Results on Benchmark Datasets
-We provide the full log-likelihood results and corresponding plots in Table 7 and Fig.
-5 respectively, where we decompose the likelihood into time and mark likelihoods.
-The improvement of our DLHP model is mainly driven by better modeling of time, though we also often obtain best- or second-best predictive performance on marks from the next event prediction accuracy results conditioned on true event time in Table 8.
-In aggregate, our model achieves a 1.38 per-event likelihood ratio between itself and the next best method across all datasets (a 38% improvement in likelihood).
+We provide the full log-likelihood results and corresponding plots in Table˜8 and Fig.˜6 respectively, where we decompose the likelihood into time and mark likelihoods.
+The improvement of our S2P2 model is mainly driven by better modeling of time, though we also often obtain best- or second-best predictive performance on marks from the next event prediction accuracy results conditioned on true event time in Table˜2(c).
+In aggregate, our model achieves a 1.33 per-event likelihood ratio between itself and the next best method across all datasets (a 33% improvement in likelihood).
 This is calculated by computing the mean log-likelihood ratio across all datasets and then exponentiating.
 Doing so is equivalent to taking the geometric mean across likelihood ratios.
-We also include root mean square error (RMSE) in Table 9 as one of the commonly used metrics in MTPP literature, with the caveat that it purely describes how well the model captures the expected value of the next-event time.
-In other words, it is not informative when the inter-event times have a multimodal distribution – which we often encounter in practice with real data.
-Intensity-based methods were sampled using the thinning algorithm and averaged over 64 sampled candidate times when memory allows.
-For our method, we include a simple linear probe on the output hidden vector to directly estimate the next event time, which is substantially faster in wallclock runtime than sampling even when accounting for fine-tuning the linear probe parameters.
-For RMSE, IFTPP is the best performer overall; however, across intensity-based methods we find our model to be competitive.
-We posit this is due to the simplified parameterization of IFTPP lending itself well to capturing the expected time of the next event.
+Configuration and training details of all models can be found in Section˜C.1.
+As discussed in Section˜7 and grouped in the results, models with continuous-time hidden states can present a richer class of intensities and often empirically outperform those with discrete hidden states.
+Note for the RMSE results in Table˜2(b), we follow Mei and Eisner [2017] and use the expected next event time as next event time predictions to minimize the Bayes risk.
+Unlike them, however, we estimate these with the trapezoidal rule rather than Monte-Carlo simulation via the thinning algorithm.
+In practice, we have found this to produce an estimator with much lower variance and be faster due to being more readily parallelizable.
+This stands in contrast to the thinning algorithm which has more hyperparameters (e.g., dominating rate, sampling boundary) that can exacerbate bias.
 Table 8: Complete per event log-likelihood results on the held-out test for the eight benchmark datasets we consider, averaged over 5 random seeds.
-In Table 7(a) we show the full log-likelihood.
-We then decompose this log-likelihood into the log-likelihood of the event time in Table 7(b), and the time-conditional log-likelihood of the mark type in Table 7(c).
+In Table˜8(a) we show the full log-likelihood.
+We then decompose this log-likelihood into the log-likelihood of the event time in Table˜8(b), and the time-conditional log-likelihood of the mark type in Table˜8(c).
 OOM indicates out of memory; standard deviation in parentheses.
-We highlight the best-performing model in bold and underline the second-best.
+We box the best-performing model and underline the second-best.
 We also report the average rank of models across datasets as a summary metric (lower is better).
-DLHP is consistently the best or second best-performing model across all datasets.
-(a) Full log-likelihood results (equal to the summation of Table 7(b) and Table 7(c)). Extended version of Table 1.
+S2P2 is consistently the best or second best-performing model across all datasets.
+(a) Full log-likelihood results (equal to the summation of Table˜8(b) and Table˜8(c)). Extended version of Table 2(a).
 Model Per Event Log-Likelihood, $\mathcal{L}_{\text{Total}}$ (nats) ($\uparrow$) Avg. Ranking ($\downarrow$)
 Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
-RMTPP -2.136 (0.003) -7.098 (0.217) 0.346 (0.002) 1.003 (0.004) -2.480 (0.019) -1.780 (0.005) -0.472 (0.026) -8.081 (0.025) 6.3
+RMTPP -2.136 (0.003) -7.098 (0.217) 0.346 (0.002) 1.003 (0.004) -2.480 (0.019) -1.780 (0.005) -0.472 (0.026) -8.081 (0.025) 7.1
+SAHP -2.074 (0.029) -6.708 (0.029) 0.298 (0.057) 1.168 (0.029) -2.341 (0.058) -1.646 (0.083) -0.677 (0.072) -6.804 (0.126) 5.8
+THP -2.096 (0.002) -6.659 (0.007) 0.372 (0.002) 0.790 (0.002) -2.338 (0.014) -1.712 (0.011) -0.577 (0.011) -7.208 (0.096) 6.1
+IFTPP 0.496 (0.002) -10.344 (0.016) 0.453 (0.002) 1.318 (0.017) -2.233 (0.009) -0.492 (0.017) 0.317 (0.052) -6.596 (0.240) 3.0
+MHP -2.091 (0.002) -6.564 (0.015) 0.370 (0.008) 0.636 (0.004) -2.346 (0.012) -1.676 (0.004) -0.351 (0.012) -7.206 (0.407) 5.9
 NHP 0.129 (0.012) -6.348 (0.000) 0.514 (0.004) 1.157 (0.004) -2.241 (0.002) -0.574 (0.011) 0.060 (0.017) -3.966 (0.058) 3.0
-SAHP -2.074 (0.029) -6.708 (0.029) 0.298 (0.057) 1.168 (0.029) -2.341 (0.058) -1.646 (0.083) -0.677 (0.072) -6.804 (0.126) 5.4
-THP -2.096 (0.002) -6.659 (0.007) 0.372 (0.002) 0.790 (0.002) -2.338 (0.014) -1.712 (0.011) -0.577 (0.011) -7.208 (0.096) 5.5
 AttNHP 0.484 (0.077) -6.499 (0.028) 0.493 (0.009) 1.259 (0.022) -2.194 (0.016) -0.592 (0.051) -0.170 (0.077) OOM 3.1
-IFTPP 0.496 (0.002) -10.344 (0.016) 0.453 (0.002) 1.318 (0.017) -2.233 (0.009) -0.492 (0.017) 0.317 (0.052) -6.596 (0.240) 2.9
-DLHP (Ours) 0.781 (0.011) -6.365 (0.003) 0.522 (0.004) 1.304 (0.039) -2.163 (0.009) -0.557 (0.046) 1.243 (0.083) -2.512 (0.369) 1.4
+S2P2 (Ours) 0.781 (0.011) -6.365 (0.003) 0.522 (0.004) 1.304 (0.039) -2.163 (0.009) -0.557 (0.046) 0.919 (0.069) -2.512 (0.369) 1.4
 (b) Per event log-likelihood of the event times (higher is better).
@@
 Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
-RMTPP 0.011 (0.001) -6.191 (0.083) 0.622 (0.002) 2.428 (0.004) -0.797 (0.005) 0.256 (0.007) -0.188 (0.016) -1.913 (0.025) 5.8
+RMTPP 0.011 (0.001) -6.191 (0.083) 0.622 (0.002) 2.428 (0.004) -0.797 (0.005) 0.256 (0.007) -0.188 (0.016) -1.913 (0.025) 6.1
+SAHP 0.115 (0.049) -5.872 (0.062) 0.645 (0.044) 2.604 (0.008) -0.703 (0.031) 0.489 (0.078) -0.244 (0.040) -1.801 (0.049) 4.9
+THP -0.068 (0.002) -5.874 (0.007) 0.621 (0.002) 2.242 (0.002) -0.772 (0.006) 0.220 (0.010) -0.271 (0.004) -1.921 (0.027) 7.0
+IFTPP 2.483 (0.001) -9.500 (0.011) 0.735 (0.002) 2.708 (0.018) -0.662 (0.007) 1.277 (0.016) 0.555 (0.050) -2.640 (0.115) 3.1
+MHP -0.064 (0.002) -5.774 (0.016) 0.620 (0.006) 2.093 (0.004) -0.761 (0.006) 0.230 (0.003) -0.140 (0.008) -2.119 (0.318) 6.4
 NHP 2.116 (0.009) -5.584 (0.001) 0.727 (0.003) 2.578 (0.006) -0.699 (0.002) 1.198 (0.006) 0.225 (0.016) -0.821 (0.045) 3.3
-SAHP 0.115 (0.049) -5.872 (0.062) 0.645 (0.044) 2.604 (0.008) -0.703 (0.031) 0.489 (0.078) -0.244 (0.040) -1.801 (0.049) 4.6
-THP -0.068 (0.002) -5.874 (0.007) 0.621 (0.002) 2.242 (0.002) -0.772 (0.006) 0.220 (0.010) -0.271 (0.004) -1.921 (0.027) 6.4
 AttNHP 2.416 (0.092) -5.726 (0.027) 0.714 (0.010) 2.654 (0.007) -0.684 (0.005) 1.203 (0.015) 0.031 (0.055) OOM 3.3
-IFTPP 2.483 (0.001) -9.500 (0.011) 0.735 (0.002) 2.708 (0.018) -0.662 (0.007) 1.277 (0.016) 0.555 (0.050) -2.640 (0.115) 2.9
-DLHP (Ours) 2.652 (0.009) -5.598 (0.002) 0.733 (0.003) 2.719 (0.038) -0.641 (0.003) 1.257 (0.022) 1.389 (0.053) 0.382 (0.362) 1.4
+S2P2 (Ours) 2.652 (0.009) -5.598 (0.002) 0.733 (0.003) 2.719 (0.038) -0.641 (0.003) 1.257 (0.022) 1.050 (0.065) 0.382 (0.362) 1.4
 (c) Per event log-likelihood of mark type conditioned on the arrival time (higher is better).
@@
 Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
-RMTPP -2.147 (0.003) -0.908 (0.141) -0.276 (0.000) -1.425 (0.002) -1.683 (0.015) -2.035 (0.004) -0.284 (0.014) -6.168 (0.025) 5.9
+RMTPP -2.147 (0.003) -0.908 (0.141) -0.276 (0.000) -1.425 (0.002) -1.683 (0.015) -2.035 (0.004) -0.284 (0.014) -6.168 (0.025) 6.8
+SAHP -2.189 (0.030) -0.836 (0.036) -0.346 (0.024) -1.436 (0.027) -1.638 (0.032) -2.136 (0.070) -0.433 (0.031) -5.003 (0.132) 6.9
+THP -2.028 (0.002) -0.785 (0.001) -0.249 (0.001) -1.451 (0.000) -1.566 (0.008) -1.932 (0.006) -0.306 (0.009) -5.287 (0.107) 5.5
+IFTPP -1.988 (0.001) -0.844 (0.007) -0.282 (0.001) -1.391 (0.005) -1.571 (0.003) -1.769 (0.004) -0.239 (0.002) -3.956 (0.192) 4.1
+MHP -2.027 (0.001) -0.790 (0.003) -0.251 (0.003) -1.456 (0.005) -1.586 (0.006) -1.906 (0.002) -0.210 (0.005) -5.087 (0.296) 5.4
 NHP -1.987 (0.003) -0.764 (0.000) -0.213 (0.002) -1.421 (0.004) -1.542 (0.001) -1.772 (0.006) -0.165 (0.002) -3.144 (0.016) 2.4
-SAHP -2.189 (0.030) -0.836 (0.036) -0.346 (0.024) -1.436 (0.027) -1.638 (0.032) -2.136 (0.070) -0.433 (0.031) -5.003 (0.132) 6.1
-THP -2.028 (0.002) -0.785 (0.001) -0.249 (0.001) -1.451 (0.000) -1.566 (0.008) -1.932 (0.006) -0.306 (0.009) -5.287 (0.107) 5.0
 AttNHP -1.933 (0.024) -0.773 (0.003) -0.221 (0.002) -1.395 (0.016) -1.510 (0.013) -1.795 (0.037) -0.201 (0.025) OOM 2.4
-IFTPP -1.988 (0.001) -0.844 (0.007) -0.282 (0.001) -1.391 (0.005) -1.571 (0.003) -1.769 (0.004) -0.239 (0.002) -3.956 (0.192) 3.8
-DLHP (Ours) -1.871 (0.002) -0.767 (0.000) -0.211 (0.002) -1.415 (0.005) -1.521 (0.008) -1.814 (0.025) -0.145 (0.040) -2.893 (0.009) 1.9
-Figure 5: Visualization of $\mathcal{L}_{\text{Total}}$ decomposed into $\mathcal{L}_{\text{Time}}$ and $\mathcal{L}_{\text{Mark}}$ for all models and all datasets relative to RMTPP, normalized by number of events, as discussed in Section 5.2.
-The improvement of DLHP is mainly driven by better modeling of $\mathcal{L}_{\text{Time}}$.
-Table 8: Next event prediction accuracy (reported as a percentage, higher is better) conditioned on the true event time.
-We report top 1 accuracy for all datasets except for top 10 accuracy for EHRSHOT, due to the pre-processing procedure described in Section C.3.
-We bold the best result per dataset, and underline the runner-up.
-Model Next Mark Accuracy (%) ($\uparrow$) Avg. Ranking ($\downarrow$)
-Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT (Top 10)
-RMTPP 30.89 53.76 91.32 60.78 46.03 52.42 92.32 36.79 6.1
-NHP 39.32 61.43 92.97 61.48 47.21 56.55 94.32 76.41 1.8
-SAHP 31.45 54.97 92.23 60.45 46.54 51.82 87.42 52.24 5.5
-THP 34.63 60.20 91.46 59.99 46.79 53.38 91.09 51.85 5.3
-AttNHP 38.88 60.90 92.74 61.32 48.38 56.39 93.65 OOM 2.6
-IFTPP 35.90 54.92 91.76 60.93 45.66 56.44 93.32 66.38 4.4
-DLHP (Ours) 40.76 61.36 93.27 61.19 47.86 56.28 96.55 79.20 1.9
-Table 9: Comparison of RMSE of the next-event time prediction on the benchmarks we consider in Section 5.2 in the same format as Table 8. Lower RMSE values indicate better performance.
-Model Next event time RMSE ($\downarrow$) Avg. Ranking ($\downarrow$)
-Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
-RMTPP 0.361 16152 0.284 0.127 1.054 15.864 0.754 3445 4.8
-NHP 0.342 15322 0.283 0.127 1.028 15.841 0.739 3430 3.1
-SAHP 0.345 16018 0.285 0.126 1.034 15.830 0.805 3418 3.8
-THP 0.335 15848 0.286 0.126 1.029 15.878 0.781 3504 4.3
-AttNHP 1.214 16220 1.273 0.130 1.311 15.889 0.852 OOM 7.0
-IFTPP 0.332 2568 0.280 0.126 0.975 15.550 0.713 1899 1.0
-DLHP (Ours, sampling) 0.395 15223 0.282 0.127 1.021 15.844 0.764 3421 3.4
-DLHP (Ours, linear probe) 0.337 14642 0.283 0.128 1.038 15.710 0.810 3350 N/A
+S2P2 (Ours) -1.871 (0.002) -0.767 (0.000) -0.211 (0.002) -1.415 (0.005) -1.521 (0.008) -1.814 (0.025) -0.131 (0.014) -2.893 (0.009) 1.9
+Figure 6: Visualization of $\mathcal{L}_{\text{Total}}$ decomposed into $\mathcal{L}_{\text{Time}}$ and $\mathcal{L}_{\text{Mark}}$ for all models and all datasets relative to RMTPP, normalized by number of events, as discussed in Section˜6.
+The improvement of S2P2 is mainly driven by better modeling of $\mathcal{L}_{\text{Time}}$, while it improves both $\mathcal{L}_{\text{Time}}$ and $\mathcal{L}_{\text{Mark}}$.
 D.2 Full Results for Synthetic Poisson Experiments
-We present the full results in Fig.
-6 for all models regarding the synthetic experiments discussed in Section 5.1.
+We present the full results in Fig.˜7 for all models regarding the synthetic Poisson experiments discussed in Section˜5.
 All models are trained until convergence using a set of 5,000 generated sequences, where we use 20 Monte Carlo points per event to estimate the integral of log-likelihood during training to accommodate the sparsity of events.
-We used small models so they do not overfit; model architecture and parameter counts are reported in Table 10.
-We plot the background intensity conditioned on empty sequences using 1,000 equidistant grid points between the start and end points.
+We used small models so they do not overfit; model architecture and parameter counts are reported in Table˜9.
+We plot the estimated intensity conditioned on empty sequences using 1,000 equidistant grid points between the start and end points.
 Our model is the only one that perfectly recovers the underlying ground truth intensity, while also using the fewest parameters.
-Figure 6: Results for all baseline models for the synthetic Poisson experiment introduced in Section 5.1.
+Figure 7: Results for all baseline models for the synthetic Poisson experiment introduced in Section 5.
 The estimated intensity (blue lines) conditioned on an empty sequence are plotted against the ground truth (dotted black lines).
@@
 RMTPP $h=16$ 627
-NHP $h=8$ 1010
 SAHP $h=16,l=2,\mathrm{heads}=4$ 1738
 THP $h=16,l=2,\mathrm{heads}=4$ 1684
+IFTPP $h=16$ 1899
+MHP $h=4$ 2240
+NHP $h=8$ 1010
 AttNHP $h=8,t=2,l=2,\mathrm{heads}=2$ 1178
-IFTPP $h=16$ 1899
-DLHP (Ours) $h=4,p=4,l=2$ 178
+S2P2 (Ours) $h=4,p=4,l=2$ 178
 D.3 Additional Synthetic Results on Multivariate Hawkes Processes
 We evaluate our model and baseline models against the true model on a randomly initiated parametric Hawkes process with three possible marks.
-Following the notation in Section 2.1, we draw all parameters from the following distributions: $\bm{\nu}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.1,0.5]$, $\bm{\alpha}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.5,0.8]$, and $\bm{\beta}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.4,1.2]$ for $i,j\in\{1,2,3\}$.
+Following the notation in Section˜2.1, we draw all parameters from the following distributions: $\bm{\nu}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.1,0.5]$, $\bm{\alpha}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.5,0.8]$, and $\bm{\beta}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Unif}[0.4,1.2]$ for $i,j\in\{1,2,3\}$.
 All models are trained until convergence using a set of 50,000 generated sequences, where we use 20 Monte Carlo points per event to estimate the integral of log-likelihood during training.
-Model architecture and parameter counts are reported in Table 11.
-We plot three example sequences drawn for an additional test set for each model in Figs.
-7 and 8, using 1,000 equidistant grid points for any inter-event interval.
+Model architecture and parameter counts are reported in Table˜10.
+We plot three example sequences drawn for an additional test set for each model in Fig.˜8, using 1,000 equidistant grid points for any inter-event interval.
 Dotted lines refer to the intensities under the true underlying parametric model; solid lines are different model estimates from trained models.
-As we see in inhomogeneous Poisson processes, our model can recover the ground truth intensities with the fewest parameters.
-Both neural Hawkes processes and our DLHP show almost perfect recovery of parametric Hawkes processes, especially before seeing any event happening, and at event times.
-It is also worth noting that our model is 7-9$\times$ quicker than NHP and AttNHP regarding wall-clock runtime on a single A5000 GPU.
-Our results on synthetic experiments validate the model’s ability to recover the ground truth intensities.
-We further evaluate all models quantitatively using 1,000 test sequences generated from the same multivariate Hawkes process and evaluated both log-likelihood and RMSE for the immediate next event.
-We see our method competitive again on both metrics.
-Table 11: Model architectures and parameter counts for parametric Hawkes processes experiments.
+[…]
+Table 10: Model architectures and parameter counts for multivariate Hawkes processes experiments.
 Model Architecture # Parameters
 RMTPP $h=16$ 697
-NHP $h=8$ 1046
 SAHP $h=16,l=2,\mathrm{heads}=4$ 1902
 THP $h=16,l=2,\mathrm{heads}=4$ 1756
+IFTPP $h=16$ 1965
+MHP $h=4$ 2264
+NHP $h=8$ 1046
 AttNHP $h=8,t=2,l=2,\mathrm{heads}=2$ 1230
-IFTPP $h=16$ 1965
-DLHP (Ours) $h=8,p=4,l=2$ 358
-Figure 7: Our proposed DLHP model trained with 50k training sequences drawn from a randomly instantiated multivariate Hawkes process. Three example test sequences are plotted for each model.
-Figure 8: Baseline models trained with 50k training sequences drawn from a randomly instantiated multivariate Hawkes process. Three example test sequences are plotted for each model.
-Table 12: Performance comparison of models on the synthetic Hawkes process experiment presented above.
-Higher per event log-likelihood indicates better performance, whereas lower root mean squared error (RMSE) indicates better performance.
+S2P2 (Ours) $h=8,p=4,l=2$ 358
+Table 11: Performance comparison of models on the multivariate Hawkes processes experiment presented above.
+Higher per-event log-likelihood indicates better performance, whereas lower root mean squared error (RMSE) indicates better performance.
 Model Total Log-Likelihood $\mathcal{L}_{\mathrm{Total}}$ ($\uparrow$) Next-Event Time RMSE ($\downarrow$)
-RMTPP -0.550 0.570
-NHP -0.530 0.565
-SAHP -0.537 0.569
-THP -0.543 0.570
-AttNHP -0.533 0.655
-IFTPP -0.534 0.540
-DLHP (Ours) -0.527 0.566
-D.4 Ablation for Different DLHP Variants
-We perform an ablation study of different model variants that we proposed on all datasets and summarize the results in Table 13.
+RMTPP -0.550 0.648
+SAHP -0.537 0.647
+THP -0.543 0.648
+IFTPP -0.534 0.647
+MHP -0.551 0.648
+NHP -0.530 0.647
+AttNHP -0.533 0.652
+S2P2 (Ours) -0.527 0.647
+Figure 8: Our proposed S2P2 model and baseline models trained with 50k training sequences drawn from a randomly instantiated multivariate Hawkes process where $K=3$.
+For each model, three example test sequences are plotted; the locations of colored bars $|||$ indicate the true event times.
+D.4 Full Results for Hawkes and Self-Correcting Process
+We train our approach on synthetic data generated from known, classical temporal point processes, namely the self-correcting and Hawkes processes.
+These are characterized by intensity functions $\lambda_{t}^{\text{SC}}=\exp\left(t-0.5N_{t}\right)$ and $\lambda_{t}^{\text{H}}=0.5+\sum_{i=1}^{N_{t}}0.5\exp(t_{i}-t)$, respectively.
+Models were fit using data drawn from these processes, 6,000 sequences for training and 2,000 for validation.
+The learned intensity functions, evaluated on a held-out test sequence, can be seen for the self-correcting process data in Fig.˜9 and for the Hawkes process data in Fig.˜10 for all methods.
+The hyperparameters for the models were chosen by a grid search (see Table˜12).
+We see many of the models, including ours, do well at capturing the ground truth intensity.
+Figure 9: Synthetic self-correcting process experiment visualization of predicted intensities compared to the ground truth intensity for a given held-out sequence.
+The vertical lines present for IFTPP are due to the conversion from density to intensity being unstable near $\Delta t=0$.
+Figure 10: Synthetic Hawkes process experiment visualization of predicted intensities compared to the ground truth intensity for a given held-out sequence.
+The vertical lines present for IFTPP are due to the conversion from density to intensity being unstable near $\Delta t=0$.
+Table 12: Model architectures and parameter counts for synthetic Hawkes and self-correcting process experiments.
+Model Architecture # Parameters
+RMTPP $h=16$ 627
+SAHP $h=16,l=3,\mathrm{heads}=4$ 2554
+THP $h=16,l=3,\mathrm{heads}=4$ 2500
+IFTPP $h=32$ 6859
+MHP $h=16$ 13556
+NHP $h=32$ 14658
+AttNHP $h=16,t=16,l=2,\mathrm{heads}=4$ 13298
+S2P2 (Ours) $h=16,p=16,l=4$ 7202
+D.5 Full Results for Long-Range Dependency Experiment
+To measure the ability to capture long-range dependencies by neural MTPPs, we constructed a generative process with long-range dependencies.
+For this, we generate sequences over the time window of $[0,100]$, with three possible marks.
+The first mark is a “distractor” mark, meaning it has no influence over other events.
+These events are drawn from a homogeneous Poisson process with rate 1.
+The second mark is a “trigger” mark, which are directly tied to the third “target” mark.
+Triggers are also drawn from a homogeneous Poisson process with rate 0.1.
+For every trigger event $(t_{i},k_{i}=2)$ drawn, a corresponding target event $(t_{j},k_{j}=3)$ is generated conditionally independent of all other events according to $t_{j}|t_{i}\sim\mathcal{N}(t_{i}+40,0.1)$.
+All models were trained with the same hyperparameters as in Table˜12.
+The predicted intensity functions for a single sequence can be seen in Fig.˜11, and the likelihood ratio between the trained models and the ground truth process on held-out test sequences can be found in Table˜13.
+We can see that both S2P2 and AttNHP do very well, both qualitatively and quantitatively.
+This is expected as both architectures are well suited for long-range dependencies while still being continuous-time models, allowing for expressive intensity functions.
+Table 13: Likelihood ratios between models and ground truth process for held-out data on long-range experiment.
+Model Ground Truth RMTPP SAHP THP IFTPP MHP NHP AttNHP S2P2
+Lik. Ratio 100% 80.4% 81.0% 94.0% 88.0% 87.3% 87.9% 99.7% 97.8%
+Figure 11: Synthetic long-range experiment visualization of predicted intensities compared to the ground truth intensity for a given held-out sequence.
+The vertical lines present for IFTPP are due to the conversion from density to intensity being unstable near $\Delta t=0$.
+D.6 Ablation for Different S2P2 Variants
+We perform an ablation study of different model variants that we proposed on all datasets and summarize the results in Table˜14.
 We train EHRSHOT using 10% of its training data because larger dataset scale requires more training time (but use the original validation and test sets for model selection and reporting results).
@@
 We select backward discretization with input dependence for the results in the main paper.
-Table 13: Ablation for different model variants log-likelihood (LL).
-ID stands for input-dependent, see Section 3.4.
-Backward and Forward respectively refer to using $\mathbf{u}_{t_{i-1}}$ and $\mathbf{u}_{t_{i}-}$ (i.e., the previous right limit or current left limit), see Section B.4.
-Dataset Model variant LL Arrival time LL Mark LL conditioned on time
+Table 14: Ablation for different model variants log-likelihood ($L$).
+ID stands for input-dependent, see Section˜$3.$ Backward and Forward respectively refer to using $\mathbf{u}_{t_{i-1}}$ and $\mathbf{u}_{t_{i}-}$ (i.e., the previous right limit or current left limit), see Section˜B.4.
+Dataset Model variant $L$ Arrival time LL Mark LL conditioned on time
 Amazon Forward 0.705 2.617 -1.912
@@
 Backward + ID -4.684 -0.641 -4.043
-D.5 Model Calibration
-To further probe the models, we evaluate the calibration metrics of MTPPs that are proposed in literature (Bosser & Taieb, 2023), which has a different focus than log-likelihood-based evaluation.
-On a high level, calibration describes how well the uncertainty in the model is reflected in the observed data.
-However, a model can achieve perfect calibration by predicting the marginal distribution, so better calibration does not necessarily transform into better predictive performance.
-We therefore present these metrics as a secondary metric (secondary to log-likelihood per Daley & Vere-Jones (2003)) for investigating the performance of different models.
-We provide summarized statistics for both probabilistic calibration error (PCE) for time calibration and expected calibration error (ECE) for mark calibration in Table 14, and visualize the calibration curves in Figs.
-9 and 10.
-From our results, all MTPP models are well-calibrated on most of the datasets, especially on mark predictions.
+D.7 Model Calibration
+To further probe the models, we evaluate the calibration of MTPPs, as proposed by Bosser and Taieb [2023].
+Calibration has a different focus than log-likelihood-based or accuracy-based evaluation.
+Calibration instead describes how well the uncertainty in the model is reflective of the observed data.
+It is important to note, however, a model can achieve perfect calibration simply by predicting the marginal distribution.
+Better calibration therefore does not necessarily indicate better predictive performance — only better calibrated errors — and so should be taken in context with the performance under other metrics.
+We provide summarized statistics for both probabilistic calibration error (PCE) for time calibration and expected calibration error (ECE) for mark calibration in Table˜15, and visualize the calibration curves in Figs.˜12 and 13.
+We see that, on the whole, MTPP models produce fairly well-calibrated predictions.
+IFTPP is the best calibrated of the models, this may be as a result of having parametric distributions for inter-arrival time (although IFTPP does fail on some datasets such as Retweet).
+The S2P2 is particularly well calibrated in time (PCE) among intensity-based methods, suggesting again that our S2P2 is capturing time dependencies better than other models.
+It is also the second-best calibrated on mark prediction (ECE) on average.
 Table 15: Calibration results for the models and datasets tests.
 (a) Probabilistic calibration error (PCE) for time calibration in percentage.
-Model Probabilistic Calibration Error (PCE) ($\downarrow$)
+Model Probabilistic Calibration Error (PCE) ($\downarrow$) Avg. Ranking ($\downarrow$)
 Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
-RMTPP 13.67 7.71 3.51 9.97 1.78 11.54 4.23 13.07
-NHP 8.16 0.14 1.27 6.66 1.42 4.60 5.87 7.42
-SAHP 10.79 6.23 1.75 3.11 1.38 10.44 2.00 20.02
-THP 12.38 5.60 3.33 16.32 1.92 10.97 1.05 13.73
-AttNHP 6.50 1.08 0.53 2.85 1.46 1.36 4.84 OOM
-IFTPP 1.54 23.35 0.47 0.61 0.68 0.30 1.89 15.14
-DLHP (Ours) 5.81 0.91 0.15 1.91 0.82 0.73 10.35 12.16
+RMTPP 13.67 (0.03) 7.93 (0.62) 3.50 (0.03) 0.22 (0.16) 1.94 (0.10) 1.56 (0.01) 3.63 (0.37) 12.60 (0.37) 6.1
+SAHP 12.04 (1.02) 8.51 (1.86) 2.52 (0.99) 3.18 (0.21) 1.50 (0.57) 2.53 (1.86) 2.28 (0.44) 20.20 (1.09) 5.3
+THP 12.38 (0.05) 5.68 (0.08) 3.34 (0.02) 6.36 (0.04) 2.06 (0.11) 1.02 (0.08) 1.10 (0.06) 13.46 (0.45) 5.4
+IFTPP 1.59 (0.09) 23.85 (0.26) 0.40 (0.10) 1.61 (0.74) 0.84 (0.34) 0.46 (0.44) 1.75 (0.33) 16.58 (3.34) 2.6
+MHP 12.22 (0.04) 4.89 (0.16) 3.43 (0.05) 8.77 (0.40) 1.58 (0.13) 1.25 (0.05) 6.21 (0.18) 15.24 (0.92) 6.0
+NHP 8.45 (0.28) 0.20 (0.19) 0.87 (0.50) 7.40 (0.68) 1.51 (0.11) 4.70 (0.13) 5.92 (0.14) 7.70 (0.49) 3.6
+AttNHP 6.36 (0.63) 2.09 (0.85) 0.84 (0.27) 3.08 (0.16) 1.65 (0.24) 1.43 (0.14) 4.70 (0.33) OOM 3.7
+S2P2 (Ours) 5.88 (0.17) 0.44 (0.27) 0.55 (0.33) 2.07 (0.32) 1.03 (0.15) 1.38 (0.52) 11.70 (0.68) 12.06 (0.54) 2.8
 (b) Expected calibration error (ECE) for mark calibration in percentage.
-Model Expected Calibration Error (ECE) ($\downarrow$)
+Model Expected Calibration Error (ECE) ($\downarrow$) Avg. Ranking ($\downarrow$)
 Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
-RMTPP 6.42 2.06 2.24 1.52 2.09 2.49 2.87 8.22
-NHP 8.02 0.41 0.77 5.28 1.15 2.73 1.92 4.21
-SAHP 8.93 8.66 6.96 2.87 1.48 7.79 5.58 4.80
-THP 2.14 1.27 1.65 6.51 0.94 4.24 2.42 9.95
-AttNHP 2.35 0.33 0.57 2.56 1.31 0.51 1.97 OOM
-IFTPP 0.60 2.93 0.87 1.48 1.41 0.59 1.66 1.92
-DLHP (Ours) 1.30 0.45 0.52 1.48 1.74 0.31 1.31 2.47
+RMTPP 6.58 (0.15) 3.99 (4.28) 2.42 (0.16) 1.89 (0.24) 2.10 (0.27) 2.47 (0.45) 2.79 (0.43) 8.47 (0.31) 5.8
+SAHP 8.17 (2.00) 6.27 (2.23) 6.77 (0.21) 2.68 (0.35) 1.71 (0.77) 6.26 (4.30) 5.41 (0.26) 5.85 (1.95) 6.8
+THP 2.06 (0.17) 1.26 (0.11) 1.76 (0.07) 6.51 (0.03) 0.81 (0.14) 3.42 (0.70) 2.16 (0.39) 8.95 (0.91) 5.0
+IFTPP 0.46 (0.10) 0.95 (1.12) 0.55 (0.19) 1.20 (0.20) 1.28 (0.54) 0.66 (0.05) 1.39 (0.23) 1.99 (0.61) 1.8
+MHP 1.65 (0.16) 1.18 (0.12) 1.91 (0.11) 4.15 (0.36) 0.82 (0.18) 2.83 (0.50) 2.22 (0.24) 10.00 (1.71) 4.8
+NHP 8.30 (0.21) 0.35 (0.06) 0.79 (0.10) 5.59 (0.69) 1.31 (0.16) 3.41 (0.41) 2.24 (0.32) 4.18 (0.69) 4.9
+AttNHP 3.13 (0.61) 0.52 (0.16) 0.56 (0.10) 2.47 (0.12) 1.37 (0.42) 0.61 (0.16) 2.23 (0.50) OOM 3.4
+S2P2 (Ours) 0.88 (0.34) 0.52 (0.13) 0.58 (0.12) 1.96 (0.67) 1.98 (0.19) 1.01 (0.63) 1.62 (0.24) 2.51 (0.44) 3.0
 Figure 12: Reliability diagram for predicted inter-arrival time for each model on all datasets. Diagonal dashed lines refer to perfect calibration.
@@
 The grey dashed lines indicate the overall prediction accuracy of the model for the next event conditioned on true event time.
-Finally, in Figs.
-11 and 12 we plot the log-likelihood of time and mark respectively, versus their corresponding calibration results, to provide an overall view of the performances of different models.
-Our DLHP model consistently achieves higher log-likelihood while maintaining good calibration on both time and mark components on most datasets.
+Finally, in Figs.˜14 and 15 we plot the log-likelihood of time and mark respectively, versus their corresponding calibration results, to provide an overall view of the performances of different models.
+Our S2P2 model consistently achieves higher log-likelihood while maintaining good calibration on both time and mark components on most datasets.
 Figure 14: Log-likelihood of time vs. PCE for all models grouped by datasets. Higher log-likelihood and lower PCE are better (i.e., top left corner).